### v1

# Revisiting Graph Neural Networks on Graph-level Tasks: Comprehensive Experiments, Analysis, and Improvements

Abstract.
Graphs are essential data structures for modeling complex interactions in domains such as social networks, molecular structures, and biological systems. Graph-level tasks, which predict properties or classes for the entire graph, are critical for applications, such as molecular property prediction and subgraph counting. Graph Neural Networks (GNNs) have shown promise in these tasks, but their evaluations are often limited to narrow datasets, tasks, and inconsistent experimental setups, restricting their generalizability. To address these limitations, we propose a unified evaluation framework for graph-level GNNs. This framework provides a standardized setting to evaluate GNNs across diverse datasets, various graph tasks (e.g., graph classification and regression), and challenging scenarios, including noisy, imbalanced, and few-shot graphs. Additionally, we propose a novel GNN model with enhanced expressivity and generalization capabilities. Specifically, we enhance the expressivity of GNNs through a $k$-path rooted subgraph approach, enabling the model to effectively count subgraphs (e.g., paths and cycles). Moreover, we introduce a unified graph contrastive learning algorithm for graphs across diverse domains, which adaptively removes unimportant edges to augment graphs, thereby significantly improving generalization performance. Extensive experiments demonstrate that our model achieves superior performance against fourteen effective baselines across twenty-seven graph datasets, establishing it as a robust and generalizable model for graph-level tasks.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2020.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at %leave␣empty␣if␣no␣availability␣url␣should␣be␣setURL_TO_YOUR_ARTIFACTS.
1. Introduction
Graphs represent objects as nodes and their relationships as edges, which are key data structures across various domains to model complex interactions (Bonifati et al., 2024; Fan, 2022; Fang et al., 2022), such as social networks, molecular structures, biological systems, etc. Graph-level tasks involve predicting properties or labels of entire graphs, rather than individual nodes or edges. These tasks are crucial in various domains, such as subgraph counting in database management systems (Fichtenberger and Peng, 2022; Li and Yu, 2024), molecular property prediction in chemistry (Zhong and Mottin, 2023; Hu et al., 2020b), and protein classification in biology (Morris et al., 2020; Li et al., 2024a). Recently, graph neural networks (Veličković et al., 2017; Hamilton et al., 2017; Liao et al., 2022; Cui et al., 2021; Li et al., 2024c) have emerged as powerful tools for modeling and analyzing graph-structured data. They learn node representations by iteratively aggregating information from their neighbors and obtain graph representations based on node representations for graph-level tasks.
Depending on the technique, current GNNs designed for graph-level tasks can be categorized into five types: node-based, hierarchical pooling (HP)-based, subgraph-based, graph learning (GL)-based, and self-supervised learning (SSL)-based GNNs. Firstly, node-based GNNs (Kipf and Welling, 2016; Hamilton et al., 2017; Xu et al., 2019; Veličković et al., 2017) learn representations for individual nodes, which are then used to form a graph-level representation by a Readout function, such as averaging the representations of all nodes. Secondly, HP-based GNNs (Mesquita et al., 2020; Dhillon et al., 2007; Li et al., 2022b; Bianchi et al., 2020) employ hierarchical pooling techniques to progressively reduce the graph size while preserving its essential structural information. HP-based GNNs aim to capture multi-level graph structures to learn a graph representation.
Thirdly, subgraph-based GNNs (Ding et al., 2024; Yan et al., 2024; Bevilacqua et al., 2024; Papp and Wattenhofer, 2022) separate a graph into several subgraphs. They learn the representations for each subgraph and then obtain the final graph representation based on these subgraph representations. Fourthly, GL-based GNNs (Fatemi et al., 2023; Zhiyao et al., 2024; Li et al., 2024d; Li et al., 2024b) assume that graphs are noisy and incomplete. These methods aim to reconstruct the graph structure and node features to enhance graph quality, enabling the learning of more accurate and reliable graph representations. Finally, SSL-based GNNs (Sun et al., 2024; Wang et al., 2022b; Inae et al., 2023; Li et al., 2024a; Thakoor et al., 2021) aim to pretrain GNNs on unlabeled graphs to learn generalizable node representations, which can then be fine-tuned with labeled data. Specifically, they pretrain GNNs by either predicting existing graph properties (e.g., node degrees and attributes) (Hou et al., 2022; Inae et al., 2023; Zhang et al., 2021; Zang et al., 2023; Wang et al., 2022b) or maximizing the similarity between graph representations obtained from different augmentations (e.g., edge dropping) of the same graph (Lee et al., 2022; Hassani and Khasahmadi, 2020; Li et al., 2024a; Sun et al., 2024).
Although various GNNs being designed for graph-level tasks, their evaluations are often restricted to a narrow range of domain-specific datasets and insufficient baseline comparisons (Zhiyao et al., 2024; Errica et al., 2019; Wang et al., 2024a). This limited scope fails to accurately reflect the advancements and capabilities of existing GNNs in handling graph-level tasks. To address the need for fair and comprehensive empirical evaluations, we first identify three key issues in existing evaluation frameworks, thereby enhancing our evaluation capabilities.
  • Issue 1. Insufficient Coverage of Data Domains. The evaluations of existing approaches often concentrate on a limited range of datasets, typically in specific domains such as chemistry or biology. This narrow focus can cause models to overfit to the unique characteristics of these datasets, limiting the generalizability and applicability of GNNs to other domain datasets, such as social networks or geometric graphs.
  • Issue 2. Insufficient Graph-level Tasks and Scenarios. Current evaluation frameworks only focus on one type of graph-level task, such as graph classification for molecular graphs. The insufficient coverage of graph tasks neglects the diversity of potential graph-level applications, such as cycle counting and path counting in graphs. Additionally, current methods only assess GNNs under the assumption of enough labeled graphs. However, there are many other realistic scenarios. For example, the graphs labels are few and imbalanced, and the graph datasets are noisy.
  • Issue 3. Lacking One Unified Evaluation Pipeline. The absence of a standardized evaluation pipeline leads to inconsistent reporting and comparison of results across studies. Different works may use varying metrics, datasets, and experimental setups, complicating the assessment of progress and the identification of truly effective models. A unified evaluation framework would facilitate more transparent and reliable comparisons
To address the above three issues, we propose a comprehensive framework to ensure a more comprehensive and fair assessment of GNNs for graph-level tasks. Firstly, we categorize existing GNNs for graph-level tasks into five types and analyze each type in depth. Secondly, we propose a unified and fair evaluation pipeline with standardized data splitting and experimental settings. We use graph datasets from diverse domains. including biology, chemistry, social networks, and geometric graphs. Thirdly, we evaluate existing graph-level GNNs on both graph classification and graph regression (e.g., cycle counting). We also consider diverse scenarios beyond those with ample labeled graphs, including noisy graphs, imbalanced graphs, and the few-shot setting.
Additionally, alongside the comprehensive evaluation framework, we propose an improved GNN model for graph-level tasks, which achieves both high expressivity and strong generalization capabilities. Firstly, to enhance the expressivity of GNNs, we introduce a general $k$-path rooted subgraph GNN, which samples subgraphs for each $k$-length path starting from every node. Our model learns representations for each subgraph and aggregates these into a comprehensive graph representation. We theoretically prove that k-path rooted subgraph GNNs can, at a minimum, count the number of $k+2$-Path and $k+3$-Cycle. Secondly, to improve generalization across diverse domains and scenarios, we propose an adaptive graph contrastive learning approach to augment graphs. By analyzing edge importance across different domains, we propose a unified metric based on node degree to remove unimportant edges, ensuring the augmented graphs retain their original structural patterns. Our model is then optimized by maximizing the similarity between augmented graphs and predicting labels consistently with the original graph. This approach makes our model resistant to noise, thereby improving its generalization ability.
We summarize the contributions of this paper as follows.
  • We systematically revisit GNNs for graph-level tasks and categorize them into five types, i.e., node-based, hierarchical pooling-based, subgraph-based, graph learning-based, and self-supervised learning-based. We provide an in-depth analysis of their distinct techniques for learning graph representations, offering a clear understanding of their strengths and limitations.
  • We propose a unified evaluation framework to address limited domain diversity, insufficient task coverage, and the lack of standardization. It includes diverse datasets, supports multiple tasks, and ensures fair comparisons across challenging scenarios like noise, imbalance, and few-shot learning.
  • We propose an expressive and generalizable GNN model, which enhances expressivity with a $k$-path rooted subgraph approach for counting paths and cycles, and improves generalization by adaptive graph contrastive learning that removes unimportant edges to handle noise and diverse domains.
  • Extensive experiments on 13 graph classification datasets, 14 graph regression datasets, and various scenarios (e.g., few-shot and noisy graphs) demonstrate the superior performance of the proposed method over state-of-the-art GNNs.
2. Preliminary and Related Work
In this section, we first introduce the graph-level tasks and then summarize existing GNNs used for these tasks. Important notations are summarized in Table 1.
2.1. Graph-level Tasks
Graphs are a fundamental data structure to model relations between nodes in real-world datasets, such as social networks (Cohen, 2016; Linghu et al., 2020), chemistry graphs (Chen et al., 2024; Cohen, 2016), etc. Formally, we denote a graph as $G_{i}(V_{i},\mathbf{A}_{i},\mathbf{X}_{i})$, where $V_{i}$, $\mathbf{A}_{i}\in\{0,1\}^{|V_{i}|\times|V_{i}|}$, $\mathbf{X}_{i}\in\{0,1\}^{|V_{i}|\times d_{x}}$, denote nodes, adjacency matrix, and node features, respectively. For each node $v\in V_{i}$, $N_{i}(v)=\{u:\mathbf{A}_{i}[u][{v}]=1\}$ is the $1$-hop neighbors of ${v}$ and $N_{i}^{l}(v)$ is the neighbors of $v$ within $l$ hops. In general, there are two types of graph-level tasks, graph classification and graph regression.
These two type of tasks involve predicting either categorical labels (classification) or a continuous value (regression) for an entire graph. Formally, given a labeled graph dataset $\mathcal{LG}=\{(G_{i},y_{i})\}_{i=1}^{n}$, each graph $G_{i}$ is labeled with discrete label vector $y_{i}\in\{0,1\}^{\mathcal{|Y|}}$ on labels $\mathcal{Y}$ for classification, or a continuous value $y_{i}\in\mathbb{R}$ for regression. The goal is to learn a mapping function $f_{\theta}$ that generalizes to unseen graphs as follows:
(1) $f_{\theta}:G_{j}\to y_{j},\ \text{where}\ y_{j}\in\{0,1\}^{\mathcal{|Y|}}\ \text{or}\quad y_{j}\in\mathbb{R}.$
Graph classification and graph regression are critical and popular across various domains, such as databases, chemistry, and biology. Specifically, in database and data management fields, graph classification tasks include query plan strategy classification (Zhao et al., 2022a; Zhou et al., 2020) and community classification (Fang et al., 2020; Su et al., 2022), which help optimize database queries and understand data relationships. Graph regression tasks involve subgraph counting (Zhao et al., 2021; Li and Yu, 2024) and cardinality estimation of graph queries (Schwabe and Aco, 2024; Wang et al., 2022a; Teng et al., 2024). Also, biology benefits from graph classification for protein structure analysis and motif identification (Zang et al., 2023), and graph regression helps predict gene expression levels (Otal et al., 2024).
2.2. Graph Neural Networks
Recently, GNNs (Demirci et al., 2022; Zou et al., 2023; Shao et al., 2022; Guliyev et al., 2024; Li and Chen, 2023; Huang et al., 2024; Wang et al., 2024b; Gao et al., 2024; Zhang et al., 2023) have emerged as powerful tools to learn node representations, due to their ability to capture complex relationships within graph structures. Existing GNNs for graph-level tasks can generally be categorized into five types: node-based, pooling-based, subgraph-based, graph learning-based, and self-supervised learning-based GNNs.
2.2.1. Node-based Graph Neural Networks
As shown in Figure 1 (a), node-based GNNs first learn representations for individual nodes, which then are used to form a graph-level representation. Generally, each GNN $f_{\theta}$ consists of two basic operations, i.e., AGG($\cdot$) and COM($\cdot$(Hamilton et al., 2017; Li and Chen, 2021), which can be formulated as two trainable parameter matrices $\mathbf{W}_{agg}$ and $\mathbf{W}_{com}$. Formally, given a graph $G_{i}(V_{i},\mathbf{A}_{i},\mathbf{X}_{i})$ and a node $v\in V_{i}$, in the $l$-th layer, $\mathsf{AGG}^{(l)}(\cdot)$ in Equation (2) first aggregates the hidden representation $\mathbf{h}_{i}^{(l-1)}(u)$ and edge representation $\mathbf{e}_{uv}$ of each neighbor $u$ of node $v$ and obtain the aggregated result $\mathbf{m}^{(l)}_{i}(v)$. Then, as shown in Equation (3), $\mathsf{COM}^{(l)}(\cdot)$ combines the aggregated neighbor information $\mathbf{m}_{i}^{(l)}(v)$ and the latest information $\mathbf{h}_{i}^{(l-1)}(v)$ of node $v$ to obtain the $l$-th layer representation $\mathbf{h}_{i}^{(l)}(v)$.
(2) $\mathbf{m}_{i}^{(l)}(v)$ $=\mathsf{AGG}^{(l)}\left(\{(\mathbf{h}_{i}^{(l-1)}(u),\mathbf{e}_{uv}):u\in\mathcal{N}_{i}(v)\}\right)$
(3) $\mathbf{h}_{i}^{(l)}(v)$ $=\sigma\left(\mathsf{COM}^{(l)}\left(\mathbf{h}_{i}^{(l-1)}(v),\mathbf{m}_{i}^{(l)}(v)\right)\right),$
where $\sigma$ denotes a non-linear function (e.g., ReLU (Li and Yuan, 2017)). $\mathbf{h}_{i}^{(0)}(v)$ is initialized as $\mathbf{X}_{i}[v]$. After $L$ layers aggregation, we can obtain node representations $\mathbf{H}_{i}(V_{i})\in\mathbb{R}^{|V_{i}|\times d_{H}}$. Then, the graph representation $\mathbf{h}_{i}$ of $G_{i}$ can be summarized from the node representation $\mathbf{H}_{i}(V_{i})$ by a $\mathsf{READOUT}$ function (Li et al., 2024a) as follows:
(4) $\mathbf{h}_{i}=\mathsf{READOUT}(\mathbf{H}_{i}(V_{i})),$
where READOUT is a permutation-invariant function, such as AVERAGE, SUM, or MAX functions (Li et al., 2024a). The $\mathsf{SUM}$ function as an example is $\mathbf{h}_{i}=\sum_{v\in V_{i}}\mathbf{H}_{i}(V_{i})[v]$.
Model Optimization. Then, we can use a decoder (e.g., 2-layer MLPs (Zhu et al., 2021)) to predict discrete class or continuous property value for each graph $G_{i}$ based $\mathbf{h}_{i}$. In summary, given the labeled training dataset $\mathcal{LG}=\{(G_{i},y_{i})\}_{i=1}^{n}$ and the prediction $y^{*}_{i}$ for each graph $G_{i}$ by the GNN $f_{\theta}$, the GNN $f_{\theta}$ can be optimized by minimizing the task loss as follows:
(5) $\theta^{*}=\arg\min_{\theta}\frac{1}{|\mathcal{LG}|}\sum_{G_{i}\in\mathcal{LG}}{\mathcal{L}_{task}(f_{\theta},G_{i},y_{i})}.$
(6) $s.t.\quad$ $\mathcal{L}_{task}(y^{*}_{i},y_{i})=\begin{cases}-\sum_{y\in{y}_{i}}y\log y^{*}_{i}[y],&{y}_{i}\in\{0,1\}^{|\mathcal{Y}|}\\ \|y^{*}_{i}-y_{i}\|_{2},&y_{i}\in\mathbb{R}\end{cases}$
2.2.2. Hierarchical Pooling-based Graph Neural Networks
As shown in Figure 1 (b), hierarchical pooling (HP)-based GNNs are designed to capture hierarchical structures in the graph, which can reduce the graph size progressively while preserving its essential structural information. This is achieved by clustering nodes into groups and summarizing their features, allowing the model to focus on condensed graph structures and finally obtain the representation of each graph.
As mentioned in Equations (2) and (3), at the $l$-th layer, node-based GNNs learn node representations using the adjacency matrix $\mathbf{A}_{i}$ and the node hidden representations $\mathbf{H}_{i}^{(l-1)}(V_{i})$. In contrast, at the $l$-th layer, HP-based GNNs first generate a new cluster assignment matrix $\mathbf{S}_{i}^{(l)}\in\{0,1\}^{n_{l-1}\times n_{l}}$, where $n_{l}<n_{l-1}$. This matrix maps the nodes $V_{i}^{(l-1)}$ from the $(l-1)$-th layer to a cluster index in $\{1,\ldots,n_{l}\}$. The new adjacency matrix $\mathbf{A}_{i}^{(l)}\in\mathbb{R}^{n_{l}\times n_{l}}$ is then reconstructed based on the cluster assignment matrix $\mathbf{S}_{i}^{(l)}$ as follows:
(7) $\mathbf{A}_{i}^{(l)}=\mathbf{S}_{i}^{(l)\top}\mathbf{A}_{i}^{(l-1)}\mathbf{S}_{i}^{(l)}.$
Subsequently, $n_{l}$ new nodes $V_{i}^{(l)}$ are created to represent the $n_{l}$ clusters. Then, the node hidden representations $\mathbf{H}^{(l-1)}_{i}(V^{(l)}_{i})$ can obtained by a READOUT operation on nodes each cluster $k\in\{1,\ldots,n_{l}\}$ as follows:
(8) $\mathbf{H}^{(l-1)}_{i}(V^{(l)}_{i})[{k}]=\textsf{READOUT}(\ \mathbf{H}^{(l-1)}_{i}(V^{(l-1)}_{i})[V_{k}]),$
where $V_{k}=\{u\mid\mathbf{S}_{i}^{(l)}[u][k]=1\}$ are nodes in the cluster $k$.
Depending on the technique of generating the cluster matrix $\mathbf{S}_{i}^{(l)}$, existing pooling-based GNNs can be classified into three types.
Table 1. Summary on important notations.
Symbols Meanings
$G_{i}(V_{i},\mathbf{A}_{i},\mathbf{X}_{i})$ The graph $G_{i}$
$y_{i}$ Discrete label or continuous property value of $G_{i}$
$N^{l}_{i}{(v)},N_{i}(v)$ The $l$-hop and 1-hop neighbors of node $v$ in $G_{i}$
$D_{i}(v),\bar{D}_{i}$ Degree of node $v$ and average degree in $G_{i}$
$f_{\theta}$ GNN model
$l,L$ GNN layer index and the total layer number
$\mathbf{h}^{(l)}_{i}(v)$ The $l$-th layer node representation of $v$ in $G_{i}$
$\mathbf{h}_{i}(v)$ Final node representation of $v$ in $G_{i}$
$\mathbf{H}^{(l)}_{i}(V_{j})$ The $l$-th layer representation of nodes $V_{j}$ in $G_{i}$
$\mathbf{H}_{i}(V_{j})$ Final representation of nodes $V_{j}$ in $G_{i}$
$\mathbf{h}_{i}$ Graph representation of $G_{i}$
${y}^{*}_{i}$ Prediction of GNN $f_{\theta}$ for $G_{i}$
$\mathbf{S}_{i}^{(l)}$ Cluster assignment matrix at $l$-th layer
$\hat{G}^{*}_{i}(V_{i},\hat{\mathbf{A}}^{*}_{i},\hat{\mathbf{X}}^{*}_{i})$ The reconstructed graph for $G_{i}$
$\tilde{G}_{i},\hat{G}_{i}$ Augmented positive views for $G_{i}$
$s({\mathbf{h}}_{i},{\mathbf{h}_{j}})$ Similarity score between $G_{i}$ and $G_{j}$
$imp_{v,u}$ Edge importance score between $v$ and $u$
$\vec{t}_{v_{i}}^{k}\in\mathcal{T}_{v_{i}}^{k}$ $k$-tuple set of each node $v_{i}\in V$
$G^{s}_{\vec{t}_{v_{i}}^{k}}$ $k$-path rooted subgraph based on $\vec{t}_{v_{i}}^{k}$
$G^{s,a}_{\vec{t}_{v_{i}}^{k}}$ Feature augmented subgraph for $G^{s}_{\vec{t}_{v_{i}}^{k}}$
$\mathcal{L}_{task}(\cdot)$ Task loss (Equation (5))
$\mathcal{L}_{cy}(\cdot)$ Consistency-aware loss (Equation (23))
Figure 1. The five types of current GNNs for graph-level tasks.
  • Similarity-based. Similarity-based approaches (Mesquita et al., 2020; Dhillon et al., 2007; Li et al., 2022b) cluster nodes based on the similarity of nodes features or structures. These methods rely on predefined similarity metrics or graph partitioning algorithms to generate the cluster assignment matrix. Graclus (Mesquita et al., 2020; Dhillon et al., 2007) and CC-GNN (Li et al., 2022b) assign nodes to different clusters based on feature similarity and graph structure.
  • Node Dropping-based. Node dropping-based methods (Cangea et al., 2018; Gao and Ji, 2019; Lee et al., 2019) focus on reducing the number of nodes in the graph by learning an importance score for each node. Nodes with the highest scores are retained, while others are dropped, effectively shrinking the graph size in a controlled manner. TopKPool (Cangea et al., 2018; Gao and Ji, 2019) and SAGPool (Lee et al., 2019) aim to learn an importance score for each node and retain only the top-$n_{l}$ nodes with the highest scores at layer $l$. In other words, they only assign one node for each cluster and directly drop the other nodes.
  • Learning-based. Learning-based approaches (Ying et al., 2018; Bianchi et al., 2020; Vaswani, 2017; Baek et al., 2021; Diehl, 2019) use neural network architectures to learn the cluster assignment matrix $\mathbf{S}_{i}^{(l)}$ based on the graph’s node features and structure, which adaptively learn how to cluster nodes in a way that preserves the graph’s essential information. Specifically, DiffPool (Ying et al., 2018) and MinCutPool (Bianchi et al., 2020) use non-linear neural networks and and GMT (Baek et al., 2021) use the multi-head attention mechanism (Vaswani, 2017), to learn a cluster matrix $\mathbf{S}_{i}^{(l)}$ based on the learned node features $\mathbf{H}_{i}^{(l-1)}$ and adjacent matrix $\mathbf{A}_{i}^{(l-1)}$. EdgePool (Diehl, 2019) proposes to learn the edge score between each pair of connected nodes and then generate the cluster matrix accordingly.
2.2.3. Subgraph-based Graph Neural Networks
Recent studies propose subgraph-based GNNs with enhanced expressive power, enabling them to more effectively capture substructure information. As shown in Figure 1 (c), the core idea is to decompose the input graph into a set of overlapping or non-overlapping subgraphs and then learn representations for each subgraph to improve the final graph representation. Formally, given a graph $G_{i}(V_{i},\mathbf{A}_{i},\mathbf{X}_{i})$, existing models first generate $n_{s}$ subgraphs ${G_{i,j}(V_{i,j},\mathbf{A}_{i,j},\mathbf{X}_{i,j})}_{j=1}^{n_{s}}$. Then, for each subgraph $G_{i,j}$, the node representation of $v\in V_{i,j}$ is learned by Equations (2) and (3). In this approach, each node $v\in V_{i}$ can have multiple representations because it may appear in multiple sampled subgraphs. To derive the final node representation $\mathbf{h}_{i}(v)$, these representations are merged. A common approach is to use a READOUT function to combine them, as follows:
(9) $\mathbf{h}_{i}(v)=\textsf{READOUT}(\mathbf{h}_{i,j}(v)|v\in V_{i,j}).$
Depending on the techniques of generating the subgraphs, current subgraph-based approaches can be classified into three types.
  • Graph Element Deletion-based. These approaches (Cotta et al., 2021; Ding et al., 2024; Papp et al., 2021; Bevilacqua et al., 2022) propose deleting specific nodes or edges to create subgraphs, enabling GNNs to focus on the most informative parts of the graph by removing noise or less relevant elements. For example, DropGNN (Papp et al., 2021) and ESAN (Bevilacqua et al., 2022) generate subgraphs through random edge deletion to enhance model expressiveness. Conversely, SGOOD (Ding et al., 2024) abstracts a superstructure from the original graphs and applies sampling and edge deletion on this superstructure, thereby creating a more diverse graphs.
  • Rooted Subgraph-based. These approaches (Bevilacqua et al., 2024; You et al., 2021; Yan et al., 2024; Frasca et al., 2022; Huang et al., 2022; Papp and Wattenhofer, 2022; Qian et al., 2022; Yang et al., 2023; Zhang and Li, 2021; Zhao et al., 2022b) focus on generating subgraphs centered around specific nodes, referred to as root nodes, with the aim of capturing the structural role and local topology of the root node within the subgraph. The primary motivation is to enhance the representational power of GNNs by encoding the root node’s position within the subgraph, as well as its relationships to other nodes within the same subgraph. Specifically, I2GNN (Huang et al., 2022), ECS (Yan et al., 2024), and ID-GNN (You et al., 2021) propose to append side information to the root nodes or edges to capture the position information, such as an ID identifier for root nodes (e.g., the root $v$ is 1 and other nodes are 0) (Huang et al., 2022; You et al., 2021), and node degree and shortest distance information (Yan et al., 2024). Also, NestGNN (Zhang and Li, 2021) and GNN-AK (Zhao et al., 2022b) use rooted subgraphs with different hops to capture the hierarchical relationships within graphs for root nodes.
  • $k$-hop Subgraph-based. Similar to rooted subgraph-based approaches (Nikolentzos et al., 2020; Feng et al., 2022; Abu-El-Haija et al., 2019; Yao et al., 2023; Sandfelder et al., 2021; Wang et al., 2021), $k$-hop subgraph-based approaches construct subgraphs based on the $k$-hop neighborhood of each node $v$, focusing on capturing the local structure around each node. Unlike rooted subgraph-based approaches, which aggregate information from 1-hop neighbors, $k$-hop subgraph-based approaches aggregate information not only from 1-hop neighbors but also directly from nodes up to $k$ hops away. Specifically, MixHop (Abu-El-Haija et al., 2019) utilizes a graph diffusion kernel to gather multi-hop neighbors and computes the final representation. SEK-GNN (Yao et al., 2023), KP-GNN (Feng et al., 2022), EGO-GNN (Sandfelder et al., 2021), and $k-$hop GNN (Nikolentzos et al., 2020) progressively updates node representations by aggregating information from neighbors within $k$-hops. MAGNA (Wang et al., 2021) proposes to learn more accurate weights between each pair of nodes based on all the paths between them within $k$-hops.
2.2.4. Graph Learning-based Graph Neural Networks
Due to uncertainty and complexity in data collection, graph structures may be redundant, biased, or noisy. Consequently, GNNs cannot learn reliable and accurate graph representations, leading to incorrect predictions. Therefore, as shown in Figure 1 (d), current researchers (Fatemi et al., 2023; Zhiyao et al., 2024; Li et al., 2024d) propose reconstructing the graph structure and node features to improve the quality of the learned graph representations. Given a labeled graph set $\mathcal{LG}=\{(G_{i},y_{i})\}_{i=1}^{n}$, the graph learning-based approaches can be formulated as bi-level optimization problem.
(10) $\theta^{*}=\min_{\theta\in\Theta}{\frac{1}{|\mathcal{LG}|}\sum_{G_{i}\in\mathcal{LG}}\mathcal{L}_{task}({f}_{\theta},\hat{G}^{*}_{i}(V_{i},\hat{\mathbf{A}}^{*}_{i},\hat{\mathbf{X}}^{*}_{i}),y_{i})}.$
(11) $s.t.\quad$ $\hat{\mathbf{A}}^{*}_{i},\hat{\mathbf{X}}^{*}_{i}=\arg\min_{\hat{\mathbf{A}}_{i},\hat{\mathbf{X}}_{i}}\mathcal{L}_{gl}(f_{\theta^{*}},\hat{G}_{i}(V_{i},\hat{\mathbf{A}}_{i},\hat{\mathbf{X}}_{i}),y_{i}),\forall G_{i}\in\mathcal{G}$
At the low level in Equation (11), current approaches propose different graph learning objectives $\mathcal{L}_{gl}(\cdot)$ to reconstruct graph structure $\mathbf{A}_{i}$ and node features $\mathbf{X}_{i}$. Then, in Equation (10), $\hat{G}^{*}_{i}(V_{i},\hat{\mathbf{A}}^{*}_{i},\hat{\mathbf{X}}^{*}_{i})$ will be used to optimize the GNNs by the loss function in Equation (5).
Depending on the techniques of reconstructing graphs, current GL-based GNNs can be categorized into three types.
  • Preprocessing-based. Preprocessing-based approaches (Wu et al., 2019; Li et al., 2022a; Entezari et al., 2020) reconstruct graphs and then use them to optimize GNNs directly. The basic idea is to first reveal the common graph patterns and modify the graph structure and node features to recover these patterns. For example, GNN-Jaccard (Wu et al., 2019) and GNAT (Li et al., 2022a) observe that similar nodes tend to be connected by edges. Therefore, they remove edges between dissimilar nodes and add edges between similar ones. Also, GNN-SVD (Entezari et al., 2020) observes that noisy edges and node features tend to increase the rank of the adjacency matrix. Therefore, GNN-SVD reconstructs graphs by decreasing the rank of the adjacency matrix.
  • Jointly Training-based. In contrast to static preprocessing-based methods, jointly training-based approaches aim to optimize the GNN while iteratively reconstructing the graph structure and node features alongside the GNN model parameters. This iterative bi-level optimization enables the GNN to dynamically adapt the graph structure and features based on task-specific objectives. For instance, several approaches (Jin et al., 2021, 2020b; Li et al., 2024b; Franceschi et al., 2019; Sun et al., 2022; Luo et al., 2021), such as ADGNN (Li et al., 2024b), ProGNN (Jin et al., 2020b), and SimPGCN (Jin et al., 2021), propose reconstructing the edges between each pair of nodes by minimizing the GNN loss on the reconstructed structure while also reducing the rank of the adjacency matrix. Alternatively, several approaches (Zhang et al., 2019; Zhou et al., 2024; Wang et al., 2023), such as MOSGSL (Zhou et al., 2024) and HGP-SL (Zhang et al., 2019), focus on first partitioning each graph into subgraphs based on node similarities and predefined motifs, and then reconstructing the edges between these subgraphs rather than at the individual node level.
2.2.5. Self Supervised Learning-based Graph Neural Networks
Self-supervised learning (SSL) has become a powerful paradigm to pretrain GNNs without the need for labeled data, which can capture the node patterns and graph patterns. As shown in Figure 1 (e), the key idea of SSL approaches is to create supervised signals directly from the structure and node features of the unlabeled graph itself, leveraging the graph’s inherent properties to guide the learning process. Formally, given a set of unlabeled graphs $\mathcal{UG}=\{G_{i}(V_{i},\mathbf{A}_{i},\mathbf{X}_{i})\}_{i=1}^{|\mathcal{UG}|}$, the GNN $f_{\theta}$ is pretrained as follows:
(12) $\theta^{\prime}=\arg\min_{\theta}\frac{1}{|\mathcal{UG}|}\sum_{G_{i}\in\mathcal{UG}}\mathcal{L}_{ssl}(f_{\theta},G_{i},Signal_{i}),$
where $Signal_{i}$ is the supervised signals from the unlabeled graph $G_{i}$ and $\theta^{\prime}$ is the optimized GNN parameters. Then, the pretrained $f_{\theta^{\prime}}$ can be used to predict graph labels or properties. Formally, given the set of labeled graphs $\mathcal{LG}=\{G_{j}(V_{j}\mathbf{A}_{j},\mathbf{X}_{j}),y_{j}\}_{j=1}^{|\mathcal{LG}|}$, the GNN $f_{\theta^{\prime}}$ is optimized as follows:
(13) $\theta^{*}=\arg\min_{\theta^{\prime}}\frac{1}{|\mathcal{LG}|}\sum_{G_{j}\in\mathcal{LG}}\mathcal{L}_{task}(f_{\theta^{\prime}},G_{j},y_{j}),$
where the task loss $\mathcal{L}_{task}(\cdot)$ is defined in Equation (5).
Depending on the technique of generating supervised signals from unlabeled graphs, SSL-based GNN approaches can be broadly categorized into two main types.
  • Pretext Task-based. Pretext task-based approaches (Hou et al., 2022; Inae et al., 2023; Zhang et al., 2021; Zang et al., 2023; Hu et al., 2020a; Jin et al., 2020a; Wang et al., 2022b) design auxiliary tasks to help the model learn useful representations from the graph structure and features without requiring external labels. Current approaches design diverse node-level and graph-level tasks to capture intrinsic graph patterns (Wang et al., 2022b), such as predicting node attribute, node degree, and the node number. For example, HMGNN (Zang et al., 2023) pretrains GNNs by predicting the links between nodes and the number of nodes in each graph. MGSSL (Zhang et al., 2021) first masks the edges among motifs and then predicts the missing edges. MoAMa (Inae et al., 2023) first masks node features and then uses GNNs to reconstruct the node features based on the masked graphs. GraphMAE (Hou et al., 2022) and GPTGNN (Hu et al., 2020a) predict both node attributes and edges.
  • Graph Contrastive Learning-based. Graph contrastive learning (GCL)-based GNNs (Wang et al., 2022c; Perozzi et al., 2014; Lee et al., 2022; Hassani and Khasahmadi, 2020; Yuan et al., 2021; Sun et al., 2024) aim to learn representations by contrasting positive and negative examples. The basic idea is to maximize the similarity between different views or augmentations of the same node or graph (positive pairs) while minimizing the similarity between different nodes or graphs (negative pairs). In such a way, GCL-based approaches can distinguish the similarity between nodes and graphs. In general, the SSL loss $\mathcal{L}_{ssl}(\cdot)$ in Equation (12) can be formulated as follows.
    (14) $\theta^{\prime}=\arg\min_{\theta}{\frac{1}{|\mathcal{UG}|}\sum_{G_{i}\in\mathcal{UG}}\mathcal{L}_{cl}(f_{\theta},\hat{G}_{i},\tilde{G}_{i},Neg_{i})}.$
    (15) $s.t.$ $\tilde{G}_{i},\hat{G}_{i}=\arg\min_{\tilde{G}_{i},\hat{G}_{i}}{\mathcal{L}_{positive}(G_{i},\mathcal{T})},\forall G_{i}\in\mathcal{UG},$
    where $\mathcal{L}_{cl}(\cdot)$ is the contrastive loss. Additionally, $\mathcal{L}_{positive}(\cdot)$ is the view generation loss used to generate two positive views (i.e., $\tilde{G}_{i}$ and $\hat{G}_{i}$ for each graph $G_{i}$) where $\mathcal{T}$ is a set of augmentation operations for generating the views. Negative views ($Neg_{i}$) are a set of graphs, which are typically randomly sampled from all graphs or from graphs with labels different from $G_{i}$. One typical $\mathcal{L}_{cl}(f_{\theta},\hat{G}_{i},\tilde{G}_{i},Neg_{i})$ in Equation (14) can be defined based on InfoNCE loss (Zhu et al., 2021; Yeh et al., 2022) as follows:
    (16) $\mathcal{L}_{cl}(\cdot)=-\log\frac{\text{s}(\mathbf{\hat{h}}_{i},\mathbf{\tilde{h}}_{i})}{\sum_{G^{\prime}_{i}\in\{\hat{G}_{i},\tilde{G}_{i}\}}\sum_{G_{j}\in A(G_{i})\}}s(\mathbf{\hat{h}}^{\prime}_{i},\mathbf{\tilde{h}}_{j})},$
    where $A(G_{i})=\hat{G}_{i}\cup\tilde{G}_{i}\cup{Neg}_{i}$ includes positive views and negative samples of $G_{i}$, and $\hat{\mathbf{h}}_{i}$ and $\tilde{\mathbf{h}}_{i}$ are the representations of graph $\hat{G}_{i}$ and $\tilde{G}_{i}$, respectively. Also, $s({\mathbf{h}}_{i},{\mathbf{h}_{j}})=\exp(cosine({\mathbf{h}}_{i},{\mathbf{h}}_{j})/\tau)$ is the similarity score between ${\mathbf{h}}_{i}$ and ${\mathbf{h}}_{j}$ based the parameter $\tau$.
    The techniques in Equation (15) are employed to generate positive views for each node. Specifically, similarity-based methods (Wang et al., 2022c; Perozzi et al., 2014; Lee et al., 2022) utilize both node features and graph structures to identify nodes in the original graph, which are similar to a given target node, as the positive nodes. Also, diffusion-based approaches (Hassani and Khasahmadi, 2020; Yuan et al., 2021; Sun et al., 2024) modify the graph’s structure (e.g., by deleting or adding edges) based on global topological information, such as personalized PageRank (PPR) (Haveliwala, 1999) and motif-preserving (Sun et al., 2024). Perturbation-based methods (Thakoor et al., 2021; Zhu et al., 2020; Zhu et al., 2021; Li et al., 2024a), on the other hand, introduce changes to both the graph’s structure and node features.
3. A Generalizable and Expressive GNN
We propose a simple yet effective contrastive learning-based GNN with high generalization and expressivity capabilities on graph-level tasks. As shown in Figure 1, our proposed GNN model consists of three components, i.e., adaptive graph augmentation, $k$-Path rooted subgraph encoder, and consistency-aware model training.
Figure 2. Framework overview.
3.1. Adaptive Graph Augmentation
To enhance the generalization ability of the GNN model, we propose using contrastive learning to optimize GNNs for graph-level tasks. As described in Section 2, the key step is to generate positive graph pairs for each graph. Here, we propose a unified contrastive learning approach for graphs in diverse domains, such as geometric graphs and biology. The basic idea is to first measure the importance of each edge concerning the whole graph and then drop the unimportant edges based on weighted sampling to generate positive pairs for $G$.
Specifically, given a graph $G_{i}(V_{i},\mathbf{A}_{i},\mathbf{X}_{i})$ with edges $E$, the edge importance of each edge $e_{v,u}\in E_{i}$ is calculated based on node degree of $v$ and $u$ (Li et al., 2024a; Zhu et al., 2020), i.e., $imp_{v,u}=\log\left(\frac{D_{i}(v)+D_{i}(u)}{2}+1\right)$, where edge connected with higher-degree nodes tends to be important because nodes with higher degrees in a graph are bridges and core nodes in a functional motif as follows:
  • Geometric Graphs. High-degree nodes act as crucial bridges, and removing their edges significantly impacts the number of paths and cycles within this graph.
  • Social Networks. Higher-degree nodes are crucial as they often represent key influencers or hubs in the network. These nodes connect different parts of the network, facilitating communication and information flow. Their removal can fragment the network and disrupt connectivity significantly.
  • Chemistry Graphs. High-degree atoms mainly determine the properties of a motif. For example, phosphorus (P) in phosphate groups ($\text{PO}_{4}$) is bonded to four oxygen atoms (O). While oxygen atoms contribute to properties like electronegativity and polarity, phosphorus is often more critical in determining the overall functionality and behavior of the phosphate group.
Then, the dropping probability $p_{v,u}$ of each edge $e_{v,u}$ can be normalized using min-max normalization (Zhu et al., 2020; Patro, 2015; Li et al., 2024a) as follows:
(17) $p_{{v,u}}=\mu\cdot\frac{imp_{max}-imp_{v,u}}{imp_{max}-imp_{min}},$
where $\mu\in[0,1]$ is a hyperparameter, and $imp_{max}=\max_{e_{i,j}\in E_{i}}imp_{i,j}$ and $imp_{min}=\min_{e_{i,j}\in E_{i}}imp_{i,j}$ are the largest and smallest edge score, respectively. Then, for each graph $G_{i}$, we can generate two positive graphs $\hat{G}_{i}$ and $\tilde{G}_{i}$ by randomly dropping unimportant edges based on Equation (17).
3.2. $k$-Path Rooted Subgraph Encoder
A more expressive GNN (Xu et al., 2019; Huang et al., 2022). can better encode substructures like paths, cycles, and other motifs, which are essential for solving downstream tasks. To enhance the expressivity ability, we propose a $k$-Path rooted subgraph encoder following (Huang et al., 2022; You et al., 2021).
3.2.1. $k$-Path Rooted Subgraph and Feature Augmentation
Given a graph $G(V,\mathbf{A},\mathbf{X})$, we define a $k$-tuple set of each node $v_{i}\in V$ as $\mathcal{T}_{v_{i}}^{k}=\{\vec{t}_{v_{i}}^{k}=(v_{i},v_{i+1},v_{i+2},\ldots,v_{i+k-1})\}$, where there is an edge between each consecutive pair $(v_{i},v_{i+1})$ and $v_{i+j}$ is actually a $j$-hop neighbor of node $v_{i}$. Particularly, for each node $v_{i}\in V$, the number of $k$-tuples starting from $v_{i}$ is $O(\overline{n}_{nei}^{k})$, where $\overline{n}_{nei}$ is the average neighbor number of nodes in $G$. Then, for each tuple in $\vec{t}_{v_{i}}^{k}\in\mathcal{T}_{v_{i}}^{k}$, we can extract a subgraph based on nodes in $\vec{t}_{v_{i}}^{k}$. Formally, given a $k$-tuple nodes $\vec{t}_{v_{i}}^{k}=(v_{i},v_{i+1},v_{i+2},\ldots,v_{i+k-1})$ and the graph $G$, we denote the rooted subgraph of $\vec{t}_{v_{i}}^{k}$ as $G^{s}_{\vec{t}_{v_{i}}^{k}}(V^{s}_{\vec{t}_{v_{i}}^{k}},\mathbf{A}^{s}_{\vec{t}_{v_{i}}^{k}},\mathbf{X}^{s}_{\vec{t}_{v_{i}}^{k}})$, which extracts $L$-hop neighbors of each node in $\vec{t}_{v_{i}}^{k}$, i.e., $V^{s}_{\vec{t}_{v_{i}}^{k}}=\cup_{v_{j}\in\vec{t}_{v_{i}}^{k}}{\cup_{l=0}^{L}\mathcal{N}^{l}(v_{j})}$ and keep the edges among nodes $V^{s}_{\vec{t}_{v_{i}}^{k}}$ in $G$.
Then, given each rooted subgraph $G^{s}_{\vec{t}_{v_{i}}^{k}}(V^{s}_{\vec{t}_{v_{i}}^{k}},\mathbf{A}^{s}_{\vec{t}_{v_{i}}^{k}},\mathbf{X}^{s}_{\vec{t}_{v_{i}}^{k}})$ of nodes $\vec{t}_{v_{i}}^{k}$, we add identifiers for each node in $\vec{t}_{v_{i}}^{k}$ based on the consecutive order. Specifically, for each node $v_{i+j}\in\vec{t}_{v_{i}}^{k}$, its node feature $\mathbf{x}(v_{i+j})\in\mathbb{R}^{d_{x}}$ is augmented as $\mathbf{x}(v_{i+j})=\mathbf{x}(v_{i+j})\oplus\mathbf{e}_{j,k}\in\mathbb{R}^{(d_{x}+k)}$, where $\mathbf{e}_{j,k}\in\{0,1\}^{k}$ is a one hot vector, where the $j$-th element is 1 and others are 0. For nodes $v_{m}\in V^{s}_{\vec{t}_{v_{i}}^{k}}\setminus\vec{t}_{v_{i}}^{k}$, its node feature $\mathbf{x}(v_{m})\in\mathbb{R}^{d_{x}}$ of is augmented as $\mathbf{x}(v_{m})=\mathbf{x}(v_{m})\oplus\mathbf{0}_{k}\in\mathbb{R}^{(d_{x}+k)}$, where $\mathbf{0}_{k}$ is a vector of length $k$ and all its components are 0. Then, we denote the feature augmented subgraph $G^{s}_{\vec{t}_{v_{i}}^{k}}$ as $G^{s,a}_{\vec{t}_{v_{i}}^{k}}(V^{s}_{\vec{t}_{v_{i}}^{k}},\mathbf{A}^{s}_{\vec{t}_{v_{i}}^{k}},\mathbf{X}^{s,a}_{\vec{t}_{v_{i}}^{k}})$.
3.2.2. Graph Representation Learning
Given each augmented rooted subgraph $G^{s,a}_{\vec{t}_{v_{i}}^{k}}$, the node representations of nodes $V^{s}_{\vec{t}_{v_{i}}^{k}}$ can be learned by a GNN encoder as $\mathbf{H}({V^{s}_{\vec{t}_{v_{i}}^{k}}})$, such as GraphSAGE (Hamilton et al., 2017) or GatedGNN (Li et al., 2016). Then, the representations of each node $v_{i}$ can be learned based on all subgraphs of each $\vec{t}_{v_{i}}^{k}\in\mathcal{T}_{v_{i}}^{k}$ as follows:
(18) $\mathbf{h}({v_{i}})=\mathsf{READOUT}_{1}(\{\mathbf{H}(V^{s}_{\vec{t}_{v_{i}}^{k}})[v_{i}]|\vec{t}_{v_{i}}^{k}\in\mathcal{T}_{v_{i}}^{k}\}),$
Where $\mathsf{READOUT}_{1}(\cdot)$ is a readout function, such as MaxPooling. Therefore, the final graph representation of $G(V,\mathbf{A},\mathbf{X})$ can be learned as follows:
(19) $\mathbf{h}=\mathsf{READOUT}_{2}(\{\mathbf{h}({v_{i}})|v_{i}\in V\}),$
The graph representation $\mathbf{h}$ can be used to predict graph labels or graph properties, as mentioned in Equation (1).
3.2.3. Theoretical Analysis
Theorem 3.1.
Given the graph $G(V,\mathbf{A},\mathbf{X})$, the $k$-Path rooted subgraph of the $k$-tuple node set $\mathcal{T}_{v_{i}}^{k}$ can count the number of $k+2$-Path, starting from the node $v_{i}$.
Proof.
We firstly give Lemma 3.2 and then prove Theorem 3.1.
Lemma 3.2.
Given a graph denoted as $G(V,\mathbf{A},\mathbf{X})$ and a $k$-tuple $\vec{t}_{v_{i}}^{k}=(v_{i},v_{i+1},v_{i+2},\ldots,v_{i+k-1})$ starting from node $v_{i}$, our model can count the number of $(k+2)$-Path between $v_{i}$ and any node $v_{j}\notin\vec{t}^{k}_{v_{i}}$ that pass through the path $(v_{i},v_{i+1},v_{i+2},\ldots,v_{i+k-1})$.
Proof.
Note that since we extract the subgraphs based on the $k$ tuple $\vec{t}_{v_{i}}^{k}$, we can know whether a node belong to this $\vec{t}_{v_{i}}^{k}$. Given a node $v_{j}\in V$ and $v\notin\vec{t}^{k}_{v_{i}}$, we can count the $k+2$-Path between $v_{i}$ and $v_{j}$ by aggregating information from neighbor. Firstly, we can identify whether a node $v_{j}$ is the one-hop neighbor of the last element of $\vec{t}_{v_{i}}^{k}$ (i.e., $v_{i+k-1}$) as follows:
(20) $\mathbf{h}^{(1)}_{\vec{t}_{v_{i}}^{k}}(v_{j})=\mathbb{I}(v_{j}\notin\vec{t}^{k}_{v_{i}}[1:k-1])\sum_{v_{m}\in N(v_{j})}{\mathbb{I}(v_{m}=v_{i+k-1})}$
where $\vec{t}_{v_{i}}^{k}[1:k-1]=(v_{i},v_{i+1},\cdots,v_{i+k-2})$ is the first $k-1$ elements of $\vec{t}_{v_{i}}^{k}$, the indicator function $\mathbb{I}(condition)=1$ if the $condition$ is true. Specifically, in Equation (20), if $v_{j}$ is the one-hop neighbor $v_{i+k-1}$, the hidden representation $\mathbf{h}^{(1)}_{\vec{t}_{v_{i}}^{k}}(v_{j})=1$ otherwise $\mathbf{h}^{(1)}_{\vec{t}_{v_{i}}^{k}}(v_{j})=0$.
Then, in Equation (21), for each node $v_{j}\notin\vec{t}^{k}_{v_{i}}$, $\mathbf{h}^{(2)}_{\vec{t}_{v_{i}}^{k}}(v_{j})$ indicate the number of 2-Path from node $v_{i+k-1}$ to the node $v_{j}$.
(21) $\mathbf{h}^{(2)}_{\vec{t}_{v_{i}}^{k}}(v_{j})=\mathbb{I}(v_{j}\notin\vec{t}^{k}_{v_{i}})\sum_{v_{m}\in N(v_{j})}{\mathbf{h}^{(1)}_{\vec{t}_{v_{i}}^{k}}(v_{m})}$
Finally, in Equation (22), for each node $v_{j}\notin\vec{t}^{k}_{v_{i}}$, $\mathbf{h}^{(2)}_{\vec{t}_{v_{i}}^{k}}(v_{j})$ indicate the number of 3-Path from node $v_{i+k-1}$ to the node $v_{j}$.
(22) $\mathbf{h}^{(3)}_{\vec{t}_{v_{i}}^{k}}(v_{j})=\mathbb{I}(v_{j}\notin\vec{t}^{k}_{v_{i}})\sum_{v_{m}\in N(v_{j})}{(\mathbf{h}^{(2)}_{\vec{t}_{v_{i}}^{k}}(v_{m})-\mathbf{h}^{(1)}_{\vec{t}_{v_{i}}^{k}}(v_{j}))},,$
If $v_{j}$ is a one-hop neighbor of $v_{i+k-1}$, there exists a recursive 3-Path of the form $(v_{i+k-1},v_{j},v_{m},v_{j})$. To eliminate these recursive 3-Path for $v_{j}$, we use $\mathbf{h}^{(2)}_{\vec{t}_{v_{i}}^{k}}(v_{m})-\mathbf{h}^{(1)}_{\vec{t}_{v_{i}}^{k}}(v_{j})$.
Since $\vec{t}^{k}_{v_{i}}$ is a $k-1$ path from node $v_{i}$ to node $v_{i+k-1}$ and $\mathbf{h}^{(3)}_{\vec{t}_{v_{i}}^{k}}(v_{j})$ is the number of 3-Path from $v_{i+k-1}$ to $v_{j}$, $\mathbf{h}^{(3)}_{\vec{t}_{v_{i}}^{k}}(v_{j})$ is the number of $k+2$-Path from $v_{i}$ to $v_{j}$ that pass through the path $\vec{t}^{k}_{v_{i}}$.
Based on Lemma 3.2, given the graph $G(V,\mathbf{A},\mathbf{X})$, we can compute the total number of $k+2$ path in the graph $G$ that starts from node $v_{i}$ as follows: $n^{k+2}_{v_{i}}=\sum_{\vec{t}^{k}_{v_{i}}\in\mathcal{T}^{k}_{v_{i}}}\sum_{v_{j}\in V}{\mathbf{h}^{(3)}_{\vec{t}_{v_{i}}^{k}}(v_{j})}.$
Theorem 3.3.
Given the graph $G(V,\mathbf{A},\mathbf{X})$, the $k$-Path rooted subgraph of the $k$-tuple node set $\mathcal{T}_{v_{i}}^{k}$ can count the number of $k+3$-Cycle of the node $v_{i}$.
Proof.
Based on the Theorem 3.1, we can counting the number of $k+2$-Path from each node $v_{i}$ that passes through the path $\vec{t}^{k}_{v_{i}}\in\mathcal{T}^{k}_{v_{i}}$, then we can cunt the number of $k+3$-Cycle of node $v_{i}$ based on its neighbor $N(v_{i})$ as $\sum_{v_{j}\in N(v_{i})}\sum_{\vec{t}^{k}_{v_{i}}\in\mathcal{T}^{k}_{v_{i}}}\mathbf{h}^{(3)}_{\vec{t}^{k}_{v_{i}}}$. ∎
3.3. Consistency-aware Model Training
We propose a contrastive loss and a task-aware consistency loss to optimize GNNs, enhancing the generalizability of learned representations and effectively capturing graph patterns. Given a batch of graphs $\mathcal{G}_{B}=\{(G_{i},y_{i})\}_{i=1}^{B}$ with size $B$, we generate two positive views $\hat{G}_{i}$ and $\tilde{G}_{i}$ for each $G_{i}\in\mathcal{G}_{B}$ by dropping edges from $G_{i}$ based on Equation (17). The contrastive loss $\mathcal{L}_{cl}(f_{\theta},G_{i})$ for each graph $G_{i}$ is defined as follows:
$-\log\frac{\sum_{\{G_{j}|y_{j}=y_{i}\}}\sum_{\mathbf{h}^{\prime}_{i}\in\{\hat{\mathbf{h}}_{i},\tilde{\mathbf{h}}_{i}\}}\sum_{\mathbf{h}^{\prime}_{j}\in\{\hat{\mathbf{h}}_{j},\tilde{\mathbf{h}}_{j}\}}{s({\mathbf{h}}^{\prime}_{i},{\mathbf{h}}^{\prime}_{j})+s(\hat{\mathbf{h}}_{i},\mathbf{\tilde{h}}_{i})}}{\sum_{G_{i},G_{j}\in\mathcal{G}_{B}}\sum_{\mathbf{h}^{\prime}_{i}\in\{\hat{\mathbf{h}}_{i},\tilde{\mathbf{h}}_{i}\}}\sum_{\mathbf{h}^{\prime}_{i}\in\{\hat{\mathbf{h}}_{j},\tilde{\mathbf{h}}_{j}\}}{s(\hat{\mathbf{h}}_{i},\hat{\mathbf{h}}_{j})}},$
where $\{G_{j}|y_{j}=y_{i}\}$ denotes the graphs in $\mathcal{G}_{B}$ that has the same label of $G_{i}$, and $s({\mathbf{h}}_{i},{\mathbf{h}_{j}})=\exp(cosine({\mathbf{h}}_{i},{\mathbf{h}}_{j})/\tau)$ is the similarity score between ${\mathbf{h}}_{i}$ and ${\mathbf{h}}_{j}$ based the temperature parameter $\tau$.
Also, we define a task-aware consistency loss. Note that two positive pairs $\hat{G}_{i}$ and $\tilde{G}_{i}$ are created by dropping unimportant edges from $G_{i}$. For graph classification tasks, the graph label of $\hat{G}_{i}$ and $\tilde{G}_{i}$ should remain similar to the original graph $G_{i}$. For graph regression tasks, such as counting tasks, the cycle (resp. path ) number in $\hat{G}_{i}$ and $\tilde{G}_{i}$ should be less than that in $G_{i}$, since dropping edges decreases the cycle and path numbers in $G_{i}$. Based on this analysis, we define the task-aware consistency loss $\mathcal{L}_{cy}(f_{\theta},\hat{G}_{i},\tilde{G}_{i},y_{i})$ for each graph $G_{i}$ as follows:
(23) $\mathcal{L}_{cy}(\cdot)=\begin{cases}-\sum_{y\in{y}_{i}}(\log{\hat{y}^{*}}_{i}[y]+\log{\tilde{y}^{*}}_{i}[y]),&{y}_{i}\in\{0,1\}^{|\mathcal{Y}|}\\ \mathsf{ReLU}(\hat{y}^{*}_{i}-y_{i})+\mathsf{ReLU}(\tilde{y}^{*}_{i}-y_{i}),&y_{i}\in\mathbb{R}\end{cases}$
In summary, we pretrain the GNNs by the combination of contrastive learning loss and task-aware consistency loss.
$\theta^{\prime}=\arg\min_{\theta}$ $\frac{1}{B}\sum_{G_{i}\in\mathcal{G}_{B}}\mathcal{L}_{cl}(f_{\theta},\hat{G}_{i},\tilde{G}_{i},Neg_{i})$
(24) $+\mathcal{L}_{cy}(f_{\theta},\hat{G}_{i},\tilde{G}_{i},y_{i})$
where $\mathcal{L}_{cl}(\cdot)$ is the contrastive loss in Equation (16), and $\mathcal{L}_{cy}(\cdot)$ is the task-aware consistency loss in Equation (23). Then, we can fine-tune the GNNs with labeled data $\mathcal{LG}$ based on the task-specific loss in Equation (5).
4. Experiments
We conduct experiments to analyze baselines and our model on both graph classification and regression tasks. Specifically, we introduce datasets and baselines in Section 4.1 and Section 4.2. We present the hyperparameter setting and the evaluation metric in Section 4.3 and Section 4.4, respectively. Effectiveness and efficiency are discussed in Section 4.5 and Section 4.6. Additionally, we evaluate baselines on more scenarios in Section 4.7.
4.1. Datasets
We use graphs from four domains, including social networks (SN), biology (BIO), chemistry (CHE), and geometric (GE) graphs. In particular, for dataset splitting, if a standard split is publicly available and commonly used in existing works, we adopt the standard split. If there is no standard split, we use a 10-fold cross-validation approach. The detailed data statistics and split ratio are in Table 2.
4.1.1. Social Networks
  • IMDB-B (Morris et al., 2020) and IMDB-M (Morris et al., 2020). Two movie datasets from the IMDB (imd, 2024), where nodes are actors, edges denote co-appearances, and labels indicate movie genres. IMDB-B and IMDB-M are binary and multiple-class classification datasets, respectively.
  • REDDIT-B (Morris et al., 2020). A binary social network dataset from Reddit (red, 2024), where nodes represent users, edges indicate interactions, and graphs are labeled as QA-based or discussion-based.
  • COLLAB-M (Morris et al., 2020). A multi-class scientific collaboration network, where nodes are researchers and edges indicate co-authorship. Each graph is labeled with different research fields.
4.1.2. Biology
  • PROTEIN-B (Morris et al., 2020). A binary protein dataset where nodes represent secondary structure elements, edges denote proximity, and graphs indicate if the molecule is an enzyme.
  • DD-B (Morris et al., 2020). Derived from a molecular protein dataset, where nodes represent amino acids, and edges indicate adjacency. Each graph is labeled to classify different types of proteins.
  • ENZYMES (Morris et al., 2020). Derived from the BRENDA enzyme database, it represents nodes as secondary structural elements, edges as adjacency, and labels each graph with one of six EC classes.
  • MolHIV (Hu et al., 2020b) and MolTox21 (Hu et al., 2020b). They are molecular datasets, where nodes represent atoms, and edges are chemical bonds. Labels in MolHIV indicate whether the molecule has HIV inhibition, and labels in MolTox21 indicate the toxicity type.
4.1.3. Chemistry
MUTAG, NCI1, MolBACE, and MolPCBA are molecular graph datasets, with nodes represent atoms, node labels indicating the atomic species (e.g., N, C, O), and edges represent chemical bonds between atoms.
  • MUTAG (Morris et al., 2020). The labels are whether a molecule has mutagenic activity, referring to its potential to cause genetic mutations.
  • NCI1 (Morris et al., 2020). The labels specify whether a molecule shows a positive response in assays for specific anticancer activity.
  • MolBACE (Hu et al., 2020b). The labels indicate whether a molecule inhibits the BACE enzyme (Beta-secretase 1).
  • MolPCBA (Hu et al., 2020b). The labels represent the activity of a molecule across 128 different bioassays, indicating whether the compound is active in a given experimental condition.
Table 2. Data statistics. CV denotes Cross-Validation.
Type Data Graphs $\overline{\mbox{Nodes}}$ $\overline{\mbox{Edges}}$ Feat Classes Split Ratio
SN IMDB-B 1,000 19.77 96.53 N/A 2 10-fold CV
IMDB-M 1,500 13.00 65.94 N/A 3 10-fold CV
REDDIT-B 2,000 429.63 497.75 N/A 2 10-fold CV
COLLAB 5,000 74.49 2,457.78 N/A 3 10-fold CV
BIO PROTEINS 1,113 39.06 72.82 3 2 10-fold CV
DD 1,178 284.32 715.66 89 2 10-fold CV
ENZYMES 600 32.63 62.14 3 6 10-fold CV
MolHIV 41,127 25.50 54.93 9 2 8/1/1
MolTox21 7,831 18.90 39.22 9 12 8/1/1
CHE MUTAG 188 17.93 19.79 7 2 10-fold CV
NCI1 4,110 29.87 32.30 37 2 10-fold CV
MolBACE 1,513 34.13 36.91 204 2 8/1/1
MolPCBA 437,929 26.00 28.10 128 128 8/1/1
GE $\{\mathbf{3,4,5}\}$-Cycle 5,000 18.80 31.34 1 1 3/2/5
$\mathbf{\{6,7,8\}}$-Cycle 5,000 18.80 31.34 1 1 3/2/5
$\mathbf{\{4,5,6\}}$-Path 5,000 18.80 31.34 1 1 3/2/5
4-Clique 5,000 18.80 31.34 1 1 3/2/5
Tailed Tri. 5,000 18.80 31.34 1 1 3/2/5
Chordal Cyc. 5,000 18.80 31.34 1 1 3/2/5
Triangle Rect. 5,000 18.80 31.34 1 1 3/2/5
4.1.4. Geometric Graphs
The counting task aims to predict the number of specific substructures in each graph. We collect 14 datasets (Zhao et al., 2022b). Labels include the counts of various substructures, including 3-Cycle, 4-Cycle, 5-Cycle, 6-Cycle, 7-Cycle, 8-Cycle, 4-Path, 5-Path, 6-Path, 4-Clique, Tailed-Triangle (Tailed Tri.), Chordal-Cycle, and Triangle-Rectangle.
4.2. Baselines
We compare and comprehensively investigate our proposed model with 14 representative and effective baselines in five types.
4.2.1. Node-based GNNs
  • GCN (Kipf and Welling, 2016). It utilizes graph convolutional layers to aggregate information from neighbors, serving as a fundamental GNN.
  • GIN (Xu et al., 2019). The GIN baseline employs sum pooling and utilizes Graph Isomorphism Network as the message passing layer.
  • GraphSAGE (SAGE) (Hamilton et al., 2017). It aggregates information from sampled neighbors using different functions, such as mean, LSTM, or pooling, to update node embeddings in a scalable manner.
4.2.2. Pooling-based GNNs
  • TopK (Cangea et al., 2018; Gao and Ji, 2019). It performs node drop by selecting the top-ranked nodes based on a learnable scoring function.
  • GMT (Baek et al., 2021). It condenses nodes into important ones based on grouping-matrix and create edges via attention mechanism.
  • EPool (Diehl, 2019). It progressively merges nodes by contracting high-scoring edges between adjacent nodes through edge clustering.
4.2.3. Subgraph-based GNNs
  • ECS (Yan et al., 2024). It integrates a base GNN with precomputed structural embeddings to efficiently count substructures using distance information and message passing.
  • AK+ (Zhao et al., 2022c). It extends base GNNs by replacing the local aggregation mechanism with a GNN that encodes general subgraphs instead of star-shaped patterns.
  • I2GNN (Huang et al., 2022). It enhances Subgraph MPNNs by employing multiple node identifiers, assigning distinct identifiers to the root node and its neighbors within each subgraph.
4.2.4. GL-based GNNs
  • VIBGSL (VIB) (Sun et al., 2022). It uses the Information Bottleneck principle to mask irrelevant features and learn a task-relevant graph structure for improved GNN performance.
  • HGPSL (HGP) (Zhang et al., 2019). It selects key nodes to form an induced subgraph and refines the graph structure using sparse attention-based structure learning.
  • MOSGSL (MO) (Zhou et al., 2024). It reconstructs key subgraphs using a motif-driven guidance module that captures and aligns discriminative structural patterns
4.2.5. SSL-based GNNs
  • RGC (Sun et al., 2024). It employs a diverse-curvature GCN and motif-aware contrastive learning on a stable kernel layer to model motif regularity and learn node representations.
  • MVGRL(MVG) (Hassani and Khasahmadi, 2020). It uses diffusion algorithms, such as Personalized PageRank, to generate two positive views of the graph and maximize their similarity.
  • Ours. It enhances expressivity with a $k$-path rooted subgraph approach and improves generalization through adaptive graph contrastive learning, ensuring robust performance across diverse tasks and scenarios.
4.3. Hyperparameter and Hardware Setting
For classification datasets, we set the batch size as 32 for four larger datasets (REDDIT-B, COLLAB, DD, and MolPCBA) and 128 for the other datasets. For regression datasets, we set the batch size as 256 The baselines and our model use different hyperparameters, which are fine-tuned by grid search based on validation data. For example, the hidden dimension is selected from $\{64,128,256,512\}$, the learning rate from $\{1e-2,1e-3,1e-4,1e-5\}$, and the dropout rate from $\{0,0.1,0.2,0.3,0.4,0.5\}$. Our model sets $k=1$ for graph classification, and sets $k=3$ for regression with 3-path sampling uniformly. All models are trained for a maximum of 2000 epochs, with early stopping applied if no improvement is observed on the validation set within 50 epochs.
All codes are executed on a CentOS 7 machine equipped with a 20-core Intel® Xeon® Silver 4210 CPU @ 2.20GHz, 8 NVIDIA GeForce RTX 2080 Ti GPUs, and 256GB of RAM.
4.4. Evaluation Metric
We define the effectiveness and efficiency metrics as follows.
4.4.1. Effectiveness Metric
Given a graph set $\mathcal{LG}=\{(G_{i},y_{i})\}_{i=1}^{|\mathcal{LG}|}$, we denote the prediction label of each graph $G_{i}$ as $\hat{y}_{i}$. For graph classification tasks, we use the Strict Accuracy (Acc), Micro-F1 (Mi-F1), and Macro-F1 (Ma-F1). Particularly, if each graph only has one label, the Micro-F1 is same as Accuracy.
  • Strict Accuracy (Acc). Strict accuracy is defined as $Acc=\frac{1}{|\mathcal{LG}|}\sum_{G_{i}\in\mathcal{LG}}{\mathbb{I}(y_{i}=\hat{y}_{i})}$, where $\mathbb{I}(y_{i}=\hat{y}_{i})=1$ if only $y_{i}=\hat{y}_{i}$.
  • Micro-F1 (Mi-F1). The Micro-precision is defined as $Mi\text{-}P=\frac{\sum_{G_{i}\in\mathcal{LG}}{|y_{i}\cap\hat{y}_{i}|}}{\sum_{G_{j}\in\mathcal{LG}}{|\hat{y}_{j}|}}$ and Micro-recall is $Mi\text{-}R=\frac{\sum_{G_{i}\in\mathcal{LG}}{|y_{i}\cap\hat{y}_{i}|}}{\sum_{G_{j}\in\mathcal{LG}}{|{y}_{j}|}}$. Then, the Micro-F1 is defined as $Mi\textit{-}F1=\frac{2\times Mi\text{-}P\times Mi\text{-}R}{Mi\text{-}P+Mi\text{-}R}$.
  • Macro-F1 (Ma-F1). The Macro-precision is defined as $Ma\text{-}P=\frac{1}{|\mathcal{LG}|}\sum_{G_{i}\in\mathcal{LG}}\frac{|y_{i}\cap\hat{y}_{i}|}{|\hat{y}_{i}|}$ and Macro-recall is defined as $Ma\text{-}R=\frac{1}{|\mathcal{LG}|}\sum_{G_{i}\in\mathcal{LG}}\frac{|y_{i}\cap\hat{y}_{i}|}{|y_{i}|}$, Macro-F1 $Ma\text{-}F1=\frac{2\times Ma\text{-}P\times Ma\text{-}R}{Ma\text{-}P+Ma\text{-}R}$.
For graph regression tasks, we use the Mean Absolute Error (MAE) and R2 as follows.
  • Mean Absolute Error (MAE). Mean Absolute Error is defined as $MAE=\frac{1}{|\mathcal{LG}|}\sum_{G_{i}\in\mathcal{LG}}{|y_{i}-\hat{y}_{i}|}$.
  • R2. The R2 is defined as $R2=1-\frac{\sum_{G_{i}\in\mathcal{LG}}(y_{i}-\hat{y}_{i})^{2}}{\sum_{G_{i}\in\mathcal{LG}}(y_{i}-\bar{y})^{2}}\in[0,1]$, where $\bar{y}=\frac{1}{|\mathcal{LG}|}\sum_{G_{i}\in\mathcal{LG}}y_{i}$ is the mean of the true values. Higher $R2$ indicates better prediction performance.
4.4.2. Efficiency Metric
We evaluate the efficiency of models on both graph classification and regression tasks based on the training time (s), inference time (s), memory usage (MB) in training and inference phrases.
Table 3. Evaluation on graph classification. The best and the second performance are highlighted in bold and in underline.
Type Dataset Metric Node-based Pooling-based Subgraph-based GL-based SSL-based
GCN GIN SAGE TopK GMT EPool ECS AK+ I2GNN VIB HGP MO RGC MVG Ours
SN IMDB-B Acc 0.6840 0.7100 0.6860 0.6860 0.7440 0.6980 0.7150 0.7290 0.7070 0.7250 0.7030 0.7310 0.6230 0.6970 0.7550
F1 0.6963 0.7117 0.6991 0.7105 0.7625 0.7146 0.7158 0.7391 0.7014 0.7185 0.7078 0.7422 0.6321 0.6940 0.7693
IMDB-M Acc 0.4620 0.4820 0.4650 0.4670 0.5080 0.4750 0.4727 0.4967 OOM 0.4720 0.4650 0.5067 0.4130 0.4867 0.5233
F1 0.4393 0.4706 0.4512 0.4435 0.4850 0.4563 0.4466 0.4839 OOM 0.4443 0.4376 0.4874 0.3867 0.4788 0.5087
REDDIT-B Acc 0.9305 0.8965 0.9094 0.9280 0.9195 0.9260 OOM OOM OOM 0.8276 OOM 0.8625 OOM OOM 0.9325
F1 0.9322 0.9026 0.9132 0.9284 0.9263 0.9303 OOM OOM OOM 0.8298 OOM 0.8722 OOM OOM 0.9330
COLLAB Acc 0.7644 0.7328 0.7406 0.7556 0.8164 0.7696 OOM OOM OOM 0.7528 0.6988 0.8278 OOM 0.7688 0.7862
F1 0.7454 0.7099 0.7169 0.7395 0.7973 0.7520 OOM OOM OOM 0.7114 0.6702 0.8046 OOM 0.7320 0.7432
BIO PROTEINS Acc 0.7286 0.7250 0.7367 0.7277 0.7440 0.7286 0.7061 0.7475 0.7071 0.7322 0.7304 0.7232 0.7015 0.7367 0.7457
F1 0.6596 0.6432 0.6617 0.6712 0.7058 0.6598 0.5838 0.6978 0.6026 0.6641 0.6653 0.6442 0.6386 0.6993 0.6998
DD Acc 0.7310 0.7233 0.7542 0.7156 0.7802 0.7343 OOM 0.7776 0.7335 0.7632 0.7598 0.7632 OOM OOM 0.7632
F1 0.6676 0.6555 0.6707 0.6384 0.7252 0.6687 OOM 0.7267 0.6402 0.6845 0.6841 0.6844 OOM OOM 0.6886
ENZYMES Acc 0.4650 0.4833 0.5183 0.4800 0.4950 0.4783 0.4617 0.5267 0.4650 0.4417 0.4667 0.5250 0.4917 0.5333 0.5533
F1 0.4775 0.4781 0.5098 0.4794 0.4237 0.4737 0.4592 0.5279 0.4628 0.4337 0.4695 0.5309 0.4902 0.5314 0.5671
MolHIV Acc 0.9696 0.9701 0.9695 0.9687 0.9691 0.9696 0.9689 0.9728 0.9703 0.9686 0.9689 0.9701 0.9686 0.9705 0.9742
F1 0.3192 0.3466 0.2991 0.2221 0.2861 0.3213 0.2537 0.3682 0.3413 0.2203 0.2518 0.2502 0.2225 0.3628 0.3693
MolTOX21 Acc 0.5548 0.5553 0.5529 0.5523 0.5412 0.5514 0.5463 0.5561 0.5580 0.5283 0.5350 0.5384 0.5112 0.5565 0.5727
Mi-F1 0.9120 0.9114 0.9096 0.9083 0.9109 0.9101 0.9121 0.9138 0.9137 0.8993 0.9004 0.9048 0.8986 0.9126 0.9174
Ma-F1 0.3628 0.3601 0.3418 0.2240 0.3452 0.3765 0.3858 0.3850 0.3815 0.2020 0.2125 0.2180 0.1981 0.3847 0.3882
CHE MUTAG Acc 0.8041 0.8570 0.8199 0.8094 0.8254 0.8091 0.8091 0.8409 0.8062 0.7643 0.7706 0.7790 0.7067 0.8607 0.9102
F1 0.8583 0.8907 0.8644 0.8590 0.8642 0.8506 0.8558 0.8770 0.8512 0.8255 0.8296 0.8447 0.8012 0.8912 0.9310
NCI1 Acc 0.8158 0.8154 0.8146 0.8169 0.7662 0.8160 0.7866 0.8187 0.7603 0.7816 0.7825 0.7852 0.6995 0.7560 0.8253
F1 0.8160 0.8153 0.8149 0.8171 0.7708 0.8168 0.7914 0.8190 0.7673 0.7827 0.7831 0.7855 0.6997 0.7557 0.8249
MoIBACE Acc 0.6711 0.6645 0.6477 0.6756 0.6513 0.6749 0.7125 0.6803 0.6078 0.6019 0.6689 0.6294 0.6089 0.7099 0.7895
F1 0.7154 0.7351 0.7099 0.7264 0.7366 0.7436 0.7400 0.7382 0.7055 0.7001 0.6545 0.7077 0.7047 0.7366 0.8242
MolPCBA Acc 0.5456 0.5466 0.5451 0.5461 0.5447 ¿3 days OOM 0.5523 0.5414 OOM OOM OOM OOM OOM 0.5531
Mi-F1 0.9850 0.9853 0.6945 0.9849 0.9836 ¿3 days OOM 0.9860 0.9850 OOM OOM OOM OOM OOM 0.9858
Ma-F1 0.1200 0.1500 0.1317 0.1333 0.0167 ¿3 days OOM 0.2140 0.2064 OOM OOM OOM OOM OOM 0.2152
Table 4. Evaluation on graph regression. The best and the second performance are highlighted in bold and in underline.
Type Dataset Metric Node-based Pooling-based Subgraph-based GL-based SSL-based
GCN GIN SAGE TopK GMT EPool ECS AK+ I2GNN VIB HGP MO RGC MVG Ours
GE 3-Cycle MAE 0.4396 0.3962 0.5119 0.4290 0.4251 0.4238 0.0192 0.0022 0.0009 0.8776 0.4369 0.5410 0.4744 0.4231 0.0006
R2 0.6970 0.7546 0.8081 0.7044 0.7107 0.5679 0.9996 0.9998 0.9999 0.0096 0.6850 0.5003 0.6062 0.7174 0.9999
4-Cycle MAE 0.2806 0.2536 0.5407 0.2772 0.2720 0.2696 0.0148 0.0223 0.0062 0.7851 0.2745 0.5442 0.5398 0.2734 0.0039
R2 0.8229 0.8864 0.4011 0.8331 0.8477 0.8406 0.9997 0.9990 0.9998 0.0157 0.8350 0.4032 0.4436 0.8606 0.9999
5-Cycle MAE 0.2782 0.1858 0.4608 0.2399 0.2338 0.2659 0.0716 0.0340 0.0115 0.9021 0.2763 0.4532 0.4570 0.2180 0.0078
R2 0.8334 0.9364 0.4815 0.8614 0.8936 0.8553 0.9862 0.9972 0.9996 0.0019 0.8190 0.6081 0.4939 0.9060 0.9998
6-Cycle MAE 0.3014 0.1895 0.4690 0.2302 0.1788 0.2926 0.0862 0.0584 0.0368 0.8876 0.3035 0.4557 0.5027 0.1773 0.0275
R2 0.7926 0.9162 0.6189 0.8803 0.9279 0.8073 0.9569 0.9956 0.9969 0.0021 0.7891 0.6219 0.6077 0.9301 0.9982
7-Cycle MAE 0.4011 0.2111 0.5898 0.2852 0.1569 0.3942 0.1560 0.0563 0.0488 0.8273 0.4220 0.5713 0.5836 0.1500 0.0414
R2 0.5887 0.8635 0.4594 0.7921 0.9354 0.6091 0.9460 0.9937 0.9946 0.0064 0.5487 0.4883 0.4692 0.9404 0.9957
8-Cycle MAE 0.4757 0.2632 0.5285 0.2910 0.1377 0.4701 0.1153 0.0493 0.0399 0.7430 0.4790 0.5340 0.4808 0.1286 0.0337
R2 0.3570 0.7220 0.2212 0.7553 0.9428 0.3840 0.9703 0.9949 0.9963 0.0330 0.3520 0.2051 0.3319 0.9466 0.9974
4-Path MAE 0.7154 0.4273 0.7340 0.5273 0.1606 0.6361 0.0243 0.0150 0.0083 0.7784 0.5019 0.7316 0.7270 0.1512 0.0066
R2 0.1574 0.6349 0.0898 0.5270 0.9459 0.3188 0.9921 0.9996 0.9998 0.0279 0.5390 0.0909 0.1310 0.9534 0.9999
5-Path MAE 0.6848 0.3949 0.6669 0.5189 0.1559 0.6133 0.0134 0.0153 0.0092 0.7510 0.4846 0.6933 0.7053 0.1407 0.0081
R2 0.1050 0.6358 0.1675 0.5015 0.9473 0.3254 0.9995 0.9995 0.9998 0.0283 0.5204 0.0823 0.0650 0.9562 0.9999
6-Path MAE 0.6164 0.3913 0.6594 0.5629 0.1394 0.6930 0.0138 0.0133 0.0089 0.7571 0.4552 0.6627 0.6589 0.1299 0.0079
R2 0.2470 0.5958 0.1286 0.4133 0.9539 0.1408 0.9995 0.9996 0.9998 0.0250 0.5459 0.1282 0.1280 0.9607 0.9998
4-Clique MAE 0.3432 0.3449 0.3498 0.3430 0.3427 0.3578 0.0091 0.0092 0.0005 0.3866 0.2355 0.3422 0.3424 0.1797 0.0004
R2 0.1612 0.1348 0.1086 0.1213 0.1022 0.1057 0.9662 0.9964 0.9999 0.0890 0.8896 0.1604 0.1610 0.9002 0.9999
Tailed Triangle MAE 0.3506 0.2886 0.3812 0.3465 0.3665 0.3405 0.0189 0.0146 0.0034 0.8934 0.3791 0.3891 0.4095 0.3277 0.0018
R2 0.7884 0.8606 0.7303 0.7813 0.7673 0.8053 0.9989 0.9995 0.9999 0.0176 0.7336 0.7008 0.6245 0.8177 0.9999
Chordal Cycle MAE 0.4377 0.3683 0.4390 0.4216 0.3690 0.4246 0.0295 0.0336 0.0044 0.8720 0.4441 0.4413 0.4728 0.3671 0.0028
R2 0.6346 0.7109 0.6002 0.6414 0.6848 0.6421 0.9970 0.9951 0.9999 0.0333 0.5746 0.5841 0.5684 0.7293 0.9999
Triangle Rectangle MAE 0.4657 0.4034 0.4794 0.4736 0.4068 0.4890 0.3619 0.3464 0.3271 0.8263 0.4286 0.4755 0.5409 0.4301 0.3123
R2 0.5211 0.6291 0.5335 0.4921 0.6235 0.4702 0.6895 0.7063 0.7241 0.0338 0.5828 0.5037 0.3852 0.5691 0.7501
4.5. Effectiveness Evaluation
In this subsection, we evaluate the effectiveness of GNNs on graph tasks from two aspects. Firstly, in Section 4.5.1, we assess whether existing GNNs can capture the core semantic patterns of graphs, where the core semantic patterns mainly determine graph labels. Secondly, in Section 4.5.2, we evaluate the expressivity ability of GNNs based, i.e., whether GNNs can distinguish graph geometric patterns, such as the number of cycles and paths.
Table 5. Efficiency evaluation. We report the total training time (Train (s)), and total inference time (Infer (s)), and GPU memory during training (T-Mem) and inference (I-Mem).
Type Dataset Metric Node-based Pooling-based Subgraph-based GL-based SSL-based
GCN GIN SAGE TopK GMT EPool ECS AK+ I2GNN VIB HGP MO RGC MVG Ours
SN REDDIT-B Train(s) 192.44 182.98 179.55 218.03 758.22 8365.60 OOM OOM OOM OOM OOM 371.86 OOM OOM 411.93
Infer(s) 28.42 26.21 26.04 32.95 30.45 1953.91 OOM OOM OOM OOM OOM 79.35 OOM OOM 34.50
T-Mem(MB) 287 288 310 287 2120 360 OOM OOM OOM OOM OOM 604 OOM OOM 4287
I-Mem(MB) 190 122 135 190 244 194 OOM OOM OOM OOM OOM 289 OOM OOM 190
BIO ENZYMES Train(s) 26.05 26.34 25.23 31.95 38.85 375.25 312.46 350.25 470.05 120.48 67.5 75.29 495.23 57.53 55.45
Infer(s) 5.19 5.33 5.04 6.84 7.17 76.30 68.74 68.93 94.92 36.12 19.20 28.27 85.11 7.76 5.80
T-Mem(MB) 83 78 77 83 228 103 582 2369 1656 323 186 506 282 324 246
I-Mem(MB) 42 36 36 42 46 45 270 879 790 152 95 194 54 58 42
CHE MUTAG Train(s) 10.78 11.03 10.05 13.50 15.96 60.13 52.05 88.70 67.33 44.37 35.32 39.99 80.49 25.15 18.72
Infer(s) 3.86 3.89 3.51 5.27 4.01 17.74 18.91 21.18 25.10 15.80 11.63 15.18 29.78 7.64 4.22
T-Mem(MB) 42 44 44 42 82 46 422 775 1306 56 48 431 134 140 102
I-Mem(MB) 19 20 21 19 24 19 245 210 722 26 27 74 24 25 19
GE 4-Cycle Train(s) 166.50 160.95 140.88 180.57 1224.39 482.95 1180.51 645.97 1487.66 395.71 284.34 394.60 1770.94 300.21 1503.16
Infer(s) 256.17 241.02 235.40 279.43 2058.62 1033.68 1776.91 920.13 2219.63 549.60 441.95 583.55 2675.22 461.87 2255.94
T-Mem(MB) 88 85 85 88 380 104 1214 1734 3080 262 269 470 238 237 4327
I-Mem(MB) 64 62 62 64 260 68 342 608 726 104 152 270 103 96 737
4.5.1. Graph Classification Tasks
We evaluate baselines and our model on the graph classification tasks in Table 3. Firstly, node-based GNNs, such as GCN and GIN, cannot achieve satisfactory performance. These models learn node representations and use global pooling, which overlooks local graph structures, resulting in suboptimal outcomes compared to other GNN types. Secondly, pooling-based approaches propose to progressively reduce the graph size hierarchically and progressively learn the graph representations, which cannot achieve the best performance. This is because they simply keep the top-k most important nodes and concentrate similar nodes into a community to learn graph representation, ignoring local graph structure information. Thirdly, subgraph-based approaches effectively capture intricate local structures within the graphs, making them well-suited for bioinformatics applications where understanding detailed interactions is essential. However, their computational intensity and potential overfitting to subgraph patterns present notable challenges. Fourthly, graph learning-based approaches reconstruct graph structure and node features to remove noisy information, thereby enhancing performance on noisy datasets, such as social networks. Lastly, SSL-based approaches propose to learn generalizable representations based on pretraining on unlabeled graphs and fine-tuning on labeled graphs. Particularly, our proposed model achieves the best or comparable performance on all datasets. This is because we propose adaptive contrastive learning and sample subgraphs to learn local structure information, thereby learning a generalizable and informative graph representation.
4.5.2. Graph Regression Tasks
For graph regression, we mainly evaluate the ability of GNNs to capture the semantics and core patterns of graphs, based on cycle counting and path counting for each graph. Lower MAE and higher R2 values indicate better performance. As shown in Table 4, node-based (except GIN), pooling-based, GL-based, and SSL-based approaches (except our model) do not achieve satisfactory performance. This is because these approaches are not designed to enhance the expressivity of GNNs, i.e., they cannot distinguish whether two graphs are isomorphic or have the same cycles. In contrast, GIN, subgraph-based GNNs, and our model aim to theoretically increase the expressivity of GNNs, thereby achieving satisfying performance. Particularly, our model performs the best on all datasets. This is because our model is guaranteed to have better expressivity, such as counting 8-Cycle, while other approaches have low expressivity. Additionally, we propose a unified contrastive learning approach to augment graphs and achieve more generalizable graph representations.
4.6. Efficiency Evaluation
We evaluate the efficiency of GNNs in terms of time and GPU peak memory consumption during both the training and inference phases. As shown in Table 5, node-based GNNs are the most efficient because they directly learn node representations from the original graphs without any additional processing. Secondly, the pooling-based approach, TopK also achieves good efficiency and low memory usage by retaining only the most important nodes However, GMT and EPool are slightly less efficient as it computes use nodes features to compute the cluster representation and pairwise similarities between connected nodes to form clusters, respectively. Thirdly, subgraph-based and GL-based approaches are less efficient due to their reliance on additional graph processing. Subgraph-based GNNs sample subgraphs for each root node or generate multiple subgraphs and learn representations for each one, which increases time and memory costs. Similarly, GL-based approaches iteratively optimize the graph structure and node features while training the GNNs, further adding to resource consumption. Lastly, SSL-based approaches are computationally intensive in training because they augment graphs into positive pairs and optimize GNNs by maximizing the similarity between these pairs, which increases time and memory usage. However, during inference, they do not require graph augmentation and operate directly on the original graphs, resulting in efficiency comparable to node-based GNNs.
(a) IMDB-M
(b) ENZYMES
(c) NCI1
(d) MUTAG
Figure 3. Robustness evaluation.
(a) IMDB-M
(b) ENZYMES
(c) NCI1
(d) 4-Cycle
Figure 4. Imbalance data evaluation.
(a) IMDB-M
(b) ENZYMES
(c) NCI1
(d) 4-Cycle
Figure 5. Few-shot evaluation.
4.7. More Scenarios
We evaluate GNNs in more realistic scenarios, including noisy graphs, imbalanced graphs, and few-shot graphs. To ensure clarity, we select one representative graph dataset from each domain for evaluation, i.e., IMDB-M, ENZYMES, NCI1, and 4-Cycle. Additionally, we compare the most effective GNN model from each GNN category, including GCN, GMT, with AK+ and MO for classification tasks, I2GNN and HGPSL for regression tasks, with our model.
4.7.1. Robustness Evaluation
We evaluate the robustness of existing GNNs on noisy and incomplete graphs. Specifically, for each graph $G(V,\mathbf{A},\mathbf{X})$, we randomly delete $\alpha||\mathbf{A}||_{0}$ existing edges by varying a ratio $\alpha\in\{0.1,0.2,0.3,0.4,0.5\}$. As shown in Figure 3, as $\alpha$ increases, the performance of GNNs decreases and our proposed model outperform other baselines under the same ratio $\alpha$. Specifically, node-based GCN shows the sharply decline due to their reliance on the original graph structure and global pooling, which fail to capture local patterns effectively. Also, Pooling-based approach GMT also performs unsatisfying as it separates nodes into different clusters based on node similarity and connectivity, while edge deletions significantly affect the cluster results. Subgraph-based approach AK+ performs better by capturing intricate local patterns, but their reliance on subgraph sampling makes them sensitive to edge deletions. GL-based method MO shows resilience by dynamically learning graph structures. Particularly, our proposed model outperforms all baselines, maintaining strong performance on all $\alpha$ values. By leveraging adaptive contrastive learning and subgraph sampling, it effectively captures local structures and mitigates the impact of noise, resulting in superior robustness.
4.7.2. Imbalance Data Evaluation
In the graph classification task, given a dataset $\mathcal{G}={(G_{i},y_{i})}$, we simulate class imbalance by setting the proportions of training samples per class as $\{1,\frac{1}{2^{\beta}},\frac{1}{3^{\beta}},\dots,\frac{1}{|\mathcal{Y}|^{\beta}}\}$, where $\beta\in\{0,0.5,1,1.5,2\}$ controls the imbalance ratio. The number of samples in the first class is fixed under all $\beta$ values. For graph regression tasks, where $y_{i}\in[y_{\text{min}},y_{\text{max}}]$, we partition the label range into three equal buckets and vary the proportions of the groups as $\{1,\frac{1}{2^{\beta}},\frac{1}{3^{\beta}}\}$ to simulate imbalance. A higher $\beta$ indicates a greater level of imbalance. As shown in Figure 4, GCN and GMT struggle with both classification and regression tasks due to their reliance on global graph aggregation, which dilutes the features of minority class nodes and edges in imbalanced datasets. Similarly, AK+ (and I2GNN) and MO (and HGP), despite incorporating subgraph-level information or reconstructing graph structures, fail to explicitly address class imbalance, limiting their ability to preserve minority class features under skewed distributions. In contrast, our model demonstrates superior robustness across all tasks. By effectively capturing task-relevant subgraphs and leveraging adaptive contrastive learning, it mitigates the dominance of majority class features and balances feature representations, preserving the features of minority classes even under severe imbalance.
4.7.3. Few-shot Evaluation
Specifically, For graph classification tasks, given a training graph dataset $\mathcal{G}=\{(G_{i},y_{i})\}$, we set the number of training graphs per class as $\gamma\in\{5,10,15,20,25\}$. For graph regression tasks, we partition the range $[y_{\text{min}},y_{\text{max}}]$ into five equal buckets to create discrete groups and sample $\gamma$ training instances for each bucket. As shown in Figure 5, GCN and GMT struggle in few-shot settings due to their reliance on global graph aggregation, which oversmooths features and fails to preserve task-relevant substructures. Similarly, AK+, I2GNN, MO, and HGP, while incorporating subgraph-level information, lack mechanisms to prioritize task-relevant features or adapt to low-data conditions, leading to suboptimal performance. In contrast, our model demonstrates superior performance across all tasks by explicitly capturing task-relevant subgraphs and leveraging adaptive contrastive learning, achieving robust generalization even with very limited training data.
5. Conclusion
In this paper, we systematically revisit existing GNNs for graph-level tasks, categorizing them into five types: node-based, hierarchical pooling-based, subgraph-based, graph learning-based, and self-supervised learning-based approaches. To address limitations in current evaluation practices, we proposed a unified evaluation framework that ensures comprehensive and fair comparisons by incorporating diverse datasets, multiple tasks, and challenging scenarios such as noisy, imbalanced, and few-shot graphs. Furthermore, we propose a novel GNN model that enhances expressivity through a $k$-path rooted subgraph approach and improves generalization via adaptive graph contrastive learning. Extensive experiments demonstrated the superior performance of our model across a wide range of graph-level tasks and twenty-seven datasets.

### v2

@@
-Revisiting Graph Neural Networks on Graph-level Tasks: Comprehensive Experiments, Analysis, and Improvements
+Revisiting Graph Neural Networks for Graph-level Tasks: Taxonomy, Empirical Study, and Future Directions
 Abstract.
-Graphs are essential data structures for modeling complex interactions in domains such as social networks, molecular structures, and biological systems.
-Graph-level tasks, which predict properties or classes for the entire graph, are critical for applications, such as molecular property prediction and subgraph counting.
-Graph Neural Networks (GNNs) have shown promise in these tasks, but their evaluations are often limited to narrow datasets, tasks, and inconsistent experimental setups, restricting their generalizability.
-To address these limitations, we propose a unified evaluation framework for graph-level GNNs.
-This framework provides a standardized setting to evaluate GNNs across diverse datasets, various graph tasks (e.g., graph classification and regression), and challenging scenarios, including noisy, imbalanced, and few-shot graphs.
-Additionally, we propose a novel GNN model with enhanced expressivity and generalization capabilities.
-Specifically, we enhance the expressivity of GNNs through a $k$-path rooted subgraph approach, enabling the model to effectively count subgraphs (e.g., paths and cycles).
-Moreover, we introduce a unified graph contrastive learning algorithm for graphs across diverse domains, which adaptively removes unimportant edges to augment graphs, thereby significantly improving generalization performance.
-Extensive experiments demonstrate that our model achieves superior performance against fourteen effective baselines across twenty-seven graph datasets, establishing it as a robust and generalizable model for graph-level tasks.
-PVLDB Reference Format:
-PVLDB, 14(1): XXX-XXX, 2020.
-doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License.
-Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license.
-For any use beyond those covered by this license, obtain permission by emailing info@vldb.org.
-Copyright is held by the owner/author(s).
-Publication rights licensed to the VLDB Endowment.
-Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
-doi:XX.XX/XXX.XX
-PVLDB Artifact Availability:
-The source code, data, and/or other artifacts have been made available at %leave␣empty␣if␣no␣availability␣url␣should␣be␣setURL_TO_YOUR_ARTIFACTS.
+Graphs are fundamental data structures for modeling complex interactions in domains such as social networks, molecular structures, and biological systems.
+Graph-level tasks, which involve predicting properties or labels for entire graphs, are crucial for applications like molecular property prediction and subgraph counting.
+While Graph Neural Networks (GNNs) have shown significant promise for these tasks, their evaluations are often limited by narrow datasets, task coverage, and inconsistent experimental setups, hindering their generalizability.
+In this paper, we present a comprehensive experimental study of GNNs on graph-level tasks, systematically categorizing them into five types: node-based, hierarchical pooling-based, subgraph-based, graph learning-based, and self-supervised learning-based GNNs.
+To address these challenges, we propose a unified evaluation framework OpenGLT for graph-level GNNs.
+OpenGLT standardizes the evaluation process across diverse datasets, multiple graph tasks (e.g., classification and regression), and real-world scenarios, including noisy, imbalanced, and few-shot graphs.
+Extensive experiments are conducted on 16 baseline models across five categories, evaluated on 13 graph classification and 13 graph regression datasets.
+These experiments provide comprehensive insights into the strengths and weaknesses of existing GNN architectures.
+Graph Neural Networks, Graph-level tasks, Unified evaluation
+copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 9–13, 2026; Jeju, Koreabooktitle: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’26), August 9–13, 2026, Jeju, Koreaisbn: 978-1-4503-XXXX-X/26/08copyright: none
 1. Introduction
 Graphs represent objects as nodes and their relationships as edges, which are key data structures across various domains to model complex interactions (Bonifati et al., 2024; Fan, 2022; Fang et al., 2022), such as social networks, molecular structures, biological systems, etc.
-Graph-level tasks involve predicting properties or labels of entire graphs, rather than individual nodes or edges.
-These tasks are crucial in various domains, such as subgraph counting in database management systems (Fichtenberger and Peng, 2022; Li and Yu, 2024), molecular property prediction in chemistry (Zhong and Mottin, 2023; Hu et al., 2020b), and protein classification in biology (Morris et al., 2020; Li et al., 2024a).
-Recently, graph neural networks (Veličković et al., 2017; Hamilton et al., 2017; Liao et al., 2022; Cui et al., 2021; Li et al., 2024c) have emerged as powerful tools for modeling and analyzing graph-structured data.
-They learn node representations by iteratively aggregating information from their neighbors and obtain graph representations based on node representations for graph-level tasks.
-Depending on the technique, current GNNs designed for graph-level tasks can be categorized into five types: node-based, hierarchical pooling (HP)-based, subgraph-based, graph learning (GL)-based, and self-supervised learning (SSL)-based GNNs.
-Firstly, node-based GNNs (Kipf and Welling, 2016; Hamilton et al., 2017; Xu et al., 2019; Veličković et al., 2017) learn representations for individual nodes, which are then used to form a graph-level representation by a Readout function, such as averaging the representations of all nodes.
-Secondly, HP-based GNNs (Mesquita et al., 2020; Dhillon et al., 2007; Li et al., 2022b; Bianchi et al., 2020) employ hierarchical pooling techniques to progressively reduce the graph size while preserving its essential structural information.
-HP-based GNNs aim to capture multi-level graph structures to learn a graph representation.
-Thirdly, subgraph-based GNNs (Ding et al., 2024; Yan et al., 2024; Bevilacqua et al., 2024; Papp and Wattenhofer, 2022) separate a graph into several subgraphs.
-They learn the representations for each subgraph and then obtain the final graph representation based on these subgraph representations.
-Fourthly, GL-based GNNs (Fatemi et al., 2023; Zhiyao et al., 2024; Li et al., 2024d; Li et al., 2024b) assume that graphs are noisy and incomplete.
-These methods aim to reconstruct the graph structure and node features to enhance graph quality, enabling the learning of more accurate and reliable graph representations.
-Finally, SSL-based GNNs (Sun et al., 2024; Wang et al., 2022b; Inae et al., 2023; Li et al., 2024a; Thakoor et al., 2021) aim to pretrain GNNs on unlabeled graphs to learn generalizable node representations, which can then be fine-tuned with labeled data.
-Specifically, they pretrain GNNs by either predicting existing graph properties (e.g., node degrees and attributes) (Hou et al., 2022; Inae et al., 2023; Zhang et al., 2021; Zang et al., 2023; Wang et al., 2022b) or maximizing the similarity between graph representations obtained from different augmentations (e.g., edge dropping) of the same graph (Lee et al., 2022; Hassani and Khasahmadi, 2020; Li et al., 2024a; Sun et al., 2024).
-Although various GNNs being designed for graph-level tasks, their evaluations are often restricted to a narrow range of domain-specific datasets and insufficient baseline comparisons (Zhiyao et al., 2024; Errica et al., 2019; Wang et al., 2024a).
-This limited scope fails to accurately reflect the advancements and capabilities of existing GNNs in handling graph-level tasks.
-To address the need for fair and comprehensive empirical evaluations, we first identify three key issues in existing evaluation frameworks, thereby enhancing our evaluation capabilities.
+Early graph representation learning methods (Perozzi and others, 2014; Grover and Leskovec, 2016; Wang et al., 2016; Dong et al., 2017) laid the foundation of graph learning.
+Graph-level tasks aim at predicting properties of an entire graph, rather than individual nodes or edges.
+These tasks are crucial in domains including subgraph counting in database management (Fichtenberger and Peng, 2022; Li and Yu, 2024), molecular property prediction in chemistry (Zhong and Mottin, 2023; Hu et al., 2020a), and protein classification in bioinformatics (Morris et al., 2020; Li et al., 2024a).
+Recently, graph neural networks (GNNs) (Wang et al., 2023b; Xiang et al., 2025; Liu et al., 2025; Song et al., 2023; Liao et al., 2022; Cui et al., 2021; Li et al., 2024e) have emerged as powerful tools for graph-structured and anomaly detection tasks (Ma et al., 2023).
+They learn node representations by iteratively aggregating neighbor information to obtain graph representations.
+Depending on the approach, we categorize GNNs for graph-level tasks into five categories: node-based, hierarchical pooling (HP)-based, subgraph-based, graph learning (GL)-based, and self-supervised learning (SSL)-based methods.
+Node-based GNNs (Kipf and Welling, 2016; Hamilton et al., 2017; Xu et al., 2019; Veličković et al., 2017) compute node representations through message passing and aggregate them via a permutation-invariant readout function, such as averaging, to form the final graph embedding.
+HP-based GNNs (Mesquita et al., 2020; Dhillon et al., 2007; Li et al., 2022b; Bianchi et al., 2020) apply pooling operations to reduce the graph size and capture hierarchical structure, yielding multi-level graph representations.
+Subgraph-based GNNs (Ding et al., 2024; Yan et al., 2024; Bevilacqua et al., 2024; Papp and Wattenhofer, 2022) divide the graph into subgraphs, learn a representation for each, and then aggregate these to represent the whole graph.
+GL-based GNNs (Fatemi et al., 2023; Zhiyao et al., 2024; Li et al., 2024d, b) enhance graph quality by reconstructing structure and features.
+SSL-based GNNs (Sun et al., 2024; Wang et al., 2022b; Inae et al., 2023; Li et al., 2024a; Thakoor et al., 2021) pretrain on unlabeled data, either by predicting graph properties or maximizing agreement between augmented views of the same graph.
+Although various GNNs have been designed for graph-level tasks, their evaluations are often restricted to a narrow range of domain-specific datasets and insufficient baseline comparisons (Zhiyao et al., 2024; Errica et al., 2020; Li et al., 2023; Wang et al., 2024b).
+To ensure fair comparison, we identify and address three key shortcomings in current evaluation frameworks.
   • Issue 1.
-Insufficient Coverage of Data Domains.
-The evaluations of existing approaches often concentrate on a limited range of datasets, typically in specific domains such as chemistry or biology.
-This narrow focus can cause models to overfit to the unique characteristics of these datasets, limiting the generalizability and applicability of GNNs to other domain datasets, such as social networks or geometric graphs.
+No clear taxonomy for GNNs on graph-level Tasks.
+Graph-level tasks require different approaches than node-level tasks, yet a clear taxonomy for GNNs on graph-level tasks is lacking.
+This gap hinders holistic understanding and systematic comparison of models.
   • Issue 2.
-Insufficient Graph-level Tasks and Scenarios.
-Current evaluation frameworks only focus on one type of graph-level task, such as graph classification for molecular graphs.
-The insufficient coverage of graph tasks neglects the diversity of potential graph-level applications, such as cycle counting and path counting in graphs.
-Additionally, current methods only assess GNNs under the assumption of enough labeled graphs.
-However, there are many other realistic scenarios.
-For example, the graphs labels are few and imbalanced, and the graph datasets are noisy.
+Restricted Coverage of GNN Architectures.
+Most evaluations focus on a limited set of architectures, such as node-based GNNs, while often ignoring more expressive models like subgraph-based GNNs.
+This narrow coverage limits performance comparisons and overlooks the strengths of diverse approaches.
   • Issue 3.
-Lacking One Unified Evaluation Pipeline.
-The absence of a standardized evaluation pipeline leads to inconsistent reporting and comparison of results across studies.
-Different works may use varying metrics, datasets, and experimental setups, complicating the assessment of progress and the identification of truly effective models.
-A unified evaluation framework would facilitate more transparent and reliable comparisons
-To address the above three issues, we propose a comprehensive framework to ensure a more comprehensive and fair assessment of GNNs for graph-level tasks.
-Firstly, we categorize existing GNNs for graph-level tasks into five types and analyze each type in depth.
-Secondly, we propose a unified and fair evaluation pipeline with standardized data splitting and experimental settings.
-We use graph datasets from diverse domains.
-including biology, chemistry, social networks, and geometric graphs.
-Thirdly, we evaluate existing graph-level GNNs on both graph classification and graph regression (e.g., cycle counting).
-We also consider diverse scenarios beyond those with ample labeled graphs, including noisy graphs, imbalanced graphs, and the few-shot setting.
-Additionally, alongside the comprehensive evaluation framework, we propose an improved GNN model for graph-level tasks, which achieves both high expressivity and strong generalization capabilities.
-Firstly, to enhance the expressivity of GNNs, we introduce a general $k$-path rooted subgraph GNN, which samples subgraphs for each $k$-length path starting from every node.
-Our model learns representations for each subgraph and aggregates these into a comprehensive graph representation.
-We theoretically prove that k-path rooted subgraph GNNs can, at a minimum, count the number of $k+2$-Path and $k+3$-Cycle.
-Secondly, to improve generalization across diverse domains and scenarios, we propose an adaptive graph contrastive learning approach to augment graphs.
-By analyzing edge importance across different domains, we propose a unified metric based on node degree to remove unimportant edges, ensuring the augmented graphs retain their original structural patterns.
-Our model is then optimized by maximizing the similarity between augmented graphs and predicting labels consistently with the original graph.
-This approach makes our model resistant to noise, thereby improving its generalization ability.
-We summarize the contributions of this paper as follows.
-  • We systematically revisit GNNs for graph-level tasks and categorize them into five types, i.e., node-based, hierarchical pooling-based, subgraph-based, graph learning-based, and self-supervised learning-based.
-We provide an in-depth analysis of their distinct techniques for learning graph representations, offering a clear understanding of their strengths and limitations.
-  • We propose a unified evaluation framework to address limited domain diversity, insufficient task coverage, and the lack of standardization.
-It includes diverse datasets, supports multiple tasks, and ensures fair comparisons across challenging scenarios like noise, imbalance, and few-shot learning.
-  • We propose an expressive and generalizable GNN model, which enhances expressivity with a $k$-path rooted subgraph approach for counting paths and cycles, and improves generalization by adaptive graph contrastive learning that removes unimportant edges to handle noise and diverse domains.
-  • Extensive experiments on 13 graph classification datasets, 14 graph regression datasets, and various scenarios (e.g., few-shot and noisy graphs) demonstrate the superior performance of the proposed method over state-of-the-art GNNs.
-2. Preliminary and Related Work
-In this section, we first introduce the graph-level tasks and then summarize existing GNNs used for these tasks. Important notations are summarized in Table 1.
+Insufficient Data Diversity.
+Current evaluations typically use datasets from a narrow range of domains, such as chemistry or biology.
+This limited diversity can lead to overfitting and restricts the generalizability of GNNs to other domains like social networks or different types of graphs.
+  • Issue 4.
+Narrow Task and Scenario Scope.
+Current evaluation frameworks typically focus on a single type of graph-level task, such as molecular graph classification, and overlook diverse applications like cycle or path counting.
+They also assume access to ample clean labeled data, neglecting real-world challenges such as noise, class imbalance, or limited labeled graphs.
+  • Issue 5.
+Inconsistent Evaluation Pipelines.
+The lack of a standardized evaluation pipeline results in inconsistent comparisons.
+Different works often use varying data splits, tuning protocols, and evaluation metrics, hindering fair assessment of model performance.
+A unified evaluation framework is needed for transparent and reliable comparisons.
+To address these five issues, we introduce OpenGLT, a comprehensive framework designed to provide a fair and thorough assessment of GNNs for graph-level tasks.
+We begin by systematically categorizing existing GNNs for graph-level tasks into five distinct types and conduct an in-depth analysis of each type to understand their unique strengths and limitations.
+Next, we introduce a unified evaluation pipeline with standardized data splitting and experimental settings.
+This pipeline incorporates graph datasets from diverse domains, including biology, chemistry, social networks, and motif graphs, ensuring broad and representative evaluations.
+Furthermore, we comprehensively evaluate the effectiveness and efficiency of existing graph-level GNNs on both graph classification and graph regression tasks, such as cycle counting.
+Beyond standard tasks, we also evaluate GNNs in real-world scenarios, including noisy graphs, imbalanced datasets, and few-shot learning settings.
+This unified framework ensures a holistic and equitable evaluation of GNNs, promoting fair comparisons and advancing their applicability across various graph-level tasks.
+We summarize the contributions of this paper on graph-level benchmarking as follows.
+  • We systematically revisit GNNs for graph-level tasks and categorize them into five types, with in-depth analysis in Section 3.
+  • We propose a unified open-source evaluation framework, OpenGLT, which covers diverse tasks, datasets, and scenarios.
+  • We conduct extensive experiments on 13 graph classification and 13 graph regression datasets, systematically comparing the effectiveness and efficiency of state-of-the-art GNNs, providing insights into their relative strengths, limitations, and directions.
+2. Preliminaries and Related Work
+In this section, we first introduce graph-level tasks and then review existing experimental studies related to these graph-level tasks. Important notations are summarized in Table 1.
 2.1. Graph-level Tasks
-Graphs are a fundamental data structure to model relations between nodes in real-world datasets, such as social networks (Cohen, 2016; Linghu et al., 2020), chemistry graphs (Chen et al., 2024; Cohen, 2016), etc.
-Formally, we denote a graph as $G_{i}(V_{i},\mathbf{A}_{i},\mathbf{X}_{i})$, where $V_{i}$, $\mathbf{A}_{i}\in\{0,1\}^{|V_{i}|\times|V_{i}|}$, $\mathbf{X}_{i}\in\{0,1\}^{|V_{i}|\times d_{x}}$, denote nodes, adjacency matrix, and node features, respectively.
-For each node $v\in V_{i}$, $N_{i}(v)=\{u:\mathbf{A}_{i}[u][{v}]=1\}$ is the $1$-hop neighbors of ${v}$ and $N_{i}^{l}(v)$ is the neighbors of $v$ within $l$ hops.
-In general, there are two types of graph-level tasks, graph classification and graph regression.
-These two type of tasks involve predicting either categorical labels (classification) or a continuous value (regression) for an entire graph.
-Formally, given a labeled graph dataset $\mathcal{LG}=\{(G_{i},y_{i})\}_{i=1}^{n}$, each graph $G_{i}$ is labeled with discrete label vector $y_{i}\in\{0,1\}^{\mathcal{|Y|}}$ on labels $\mathcal{Y}$ for classification, or a continuous value $y_{i}\in\mathbb{R}$ for regression.
-The goal is to learn a mapping function $f_{\theta}$ that generalizes to unseen graphs as follows:
-(1) $f_{\theta}:G_{j}\to y_{j},\ \text{where}\ y_{j}\in\{0,1\}^{\mathcal{|Y|}}\ \text{or}\quad y_{j}\in\mathbb{R}.$
-Graph classification and graph regression are critical and popular across various domains, such as databases, chemistry, and biology.
-Specifically, in database and data management fields, graph classification tasks include query plan strategy classification (Zhao et al., 2022a; Zhou et al., 2020) and community classification (Fang et al., 2020; Su et al., 2022), which help optimize database queries and understand data relationships.
-Graph regression tasks involve subgraph counting (Zhao et al., 2021; Li and Yu, 2024) and cardinality estimation of graph queries (Schwabe and Aco, 2024; Wang et al., 2022a; Teng et al., 2024).
-Also, biology benefits from graph classification for protein structure analysis and motif identification (Zang et al., 2023), and graph regression helps predict gene expression levels (Otal et al., 2024).
-2.2. Graph Neural Networks
-Recently, GNNs (Demirci et al., 2022; Zou et al., 2023; Shao et al., 2022; Guliyev et al., 2024; Li and Chen, 2023; Huang et al., 2024; Wang et al., 2024b; Gao et al., 2024; Zhang et al., 2023) have emerged as powerful tools to learn node representations, due to their ability to capture complex relationships within graph structures.
+Graphs are a fundamental data structure to model relations between nodes in real-world datasets.
+Consequently, graph-level tasks, including classification and regression, are widely applied across domains, such as recommendation systems in social networks (Cohen, 2016; Linghu et al., 2020), molecular property and activity prediction in chemistry (Chen et al., 2024; Cohen, 2016; Mandal et al., 2022; Niazi and Mariam, 2023; Gilmer et al., 2017), and protein analysis, motif identification, or gene expression prediction in biology (Zang et al., 2023; Otal et al., 2024).
+Formally, we denote a graph as $G_{i}(V_{i},\mathbf{A}_{i},\mathbf{X}_{i})$, where $V_{i}$, $\mathbf{A}_{i}\in\{0,1\}^{|V_{i}|\times|V_{i}|}$, $\mathbf{X}_{i}\in\mathbb{R}^{|V_{i}|\times d_{x}}$, denote nodes, adjacency matrix, and node features, respectively.
+In general, graph-level tasks are to learn a mapping from $G_{i}$ to a label (classification) or a continuous value (regression).
+2.2. Related Works of Experimental Studies
+As summarized in Table 2, existing benchmarks and surveys (Zhiyao et al., 2024; Li et al., 2024c; Hu et al., 2020a; Errica et al., 2020; Wang et al., 2024b) often face four key limitations (Zhiyao et al., 2024; Li et al., 2024c; Hu et al., 2020a; Errica et al., 2020; Wang et al., 2024b): (i) a lack of systematic taxonomy for graph-level tasks; (ii) a predominant focus on node-based GNNs (Dwivedi et al., 2023; Hu et al., 2020a; Morris et al., 2020), leaving diverse architectures partially absent with poor categorization; (iii) the neglect of realistic scenarios such as noise, imbalance, and few-shot settings, which hinders robustness assessment; and (iv) a lack of comprehensive efficiency metrics for usability evaluation.
+Table 1. Summary on important notations.
+Symbols Meanings
+$G_{i}(V_{i},\mathbf{A}_{i},\mathbf{X}_{i})$ The graph $G_{i}$
+$y_{i}$ Discrete label or continuous value of $G_{i}$
+$N^{l}_{i}{(v)},N_{i}(v)$ The $l$-hop and 1-hop neighbors of $v$ in $G_{i}$
+$f_{\theta}$ GNN model
+$l,L$ GNN layer index and the total layer number
+$\mathbf{h}^{(l)}_{i}(v)$ The $l$-th layer representation of $v$ in $G_{i}$
+$m^{(l)}_{i}(v)$ The $l$-th layer message aggregated intermediate for node $v$
+$e_{uv}$ Edge feature of edge $(u,v)$ in $G_{i}$
+$\mathbf{h}_{i}(v)$ Final node representation of $v$ in $G_{i}$
+$\mathbf{H}^{(l)}_{i}(V_{j})$ $l$-th layer representation of nodes $V_{j}$ in $G_{i}$
+$\mathbf{H}_{i}(V_{j})$ Final representation of nodes $V_{j}$ in $G_{i}$
+$\mathrm{READOUT}(\cdot)$ Graph-level aggregation function producing $\mathbf{h}_{i}$ from node/cluster embeddings
+$\mathbf{h}_{i}$ Graph representation of $G_{i}$
+${y}^{*}_{i}$ Prediction of GNN $f_{\theta}$ for $G_{i}$
+$\mathbf{S}_{i}^{(l)}$ Cluster assignment matrix at $l$-th layer
+$\hat{G}^{*}_{i}(V_{i},\hat{\mathbf{A}}^{*}_{i},\hat{\mathbf{X}}^{*}_{i})$ The reconstructed graph for $G_{i}$
+$\tilde{G}_{i},\hat{G}_{i}$ Augmented positive views for $G_{i}$
+$s({\mathbf{h}}_{i},{\mathbf{h}_{j}})$ Similarity score between $G_{i}$ and $G_{j}$
+$\mathcal{L}_{task}(\cdot)$ Task loss (Equation (4))
+$\mathcal{L}_{ssl}(\cdot)$ Self-supervised learning loss (Equation (11))
+$\mathcal{L}_{cl}(\cdot)$ Contrastive learning loss (Equation (15))
+Table 2.
+Summary of existing surveys and experimental studies on GNNs for graph-level tasks.
+Sur.
+and Exp.
+denote Survey and Experiments, respectively.
+Taxo.
+denotes Taxonomy, Subg.
+denotes Subgraph.
+Additionally, FewS., Imba., Effect., and Effic.
+denote Few-shot, Imbalanced, Effectiveness, and Efficiency, respectively.
+Paper Paper Type Taxo. GNN Type Data Type Scenarios Eval Metric
+Sur. Exp. Node Pool Subg. GL SSL SN BIO CHE MC Clean Noise FewS. Imba. Effect. Effic.
+GNNS (Wu et al., 2020)
+TUD (Morris et al., 2020)
+OGB (Hu et al., 2020a)
+GNNB (Dwivedi et al., 2023)
+ReGCB (Li et al., 2024c)
+FGNNB (Errica et al., 2020)
+GPB (Wang et al., 2024b)
+OpenGLT
+3. Graph Neural Networks for Graph-level Tasks
+Recently, GNNs (Demirci et al., 2022; Zou et al., 2023; Shao et al., 2022; Guliyev et al., 2024; Li and Chen, 2023; Huang et al., 2024; Wang et al., 2024a; Gao et al., 2024; Zhang et al., 2023) have emerged as powerful tools to learn node representations, due to their ability to capture complex relationships within graph structures.
 Existing GNNs for graph-level tasks can generally be categorized into five types: node-based, pooling-based, subgraph-based, graph learning-based, and self-supervised learning-based GNNs.
-2.2.1. Node-based Graph Neural Networks
-As shown in Figure 1 (a), node-based GNNs first learn representations for individual nodes, which then are used to form a graph-level representation.
-Generally, each GNN $f_{\theta}$ consists of two basic operations, i.e., AGG($\cdot$) and COM($\cdot$(Hamilton et al., 2017; Li and Chen, 2021), which can be formulated as two trainable parameter matrices $\mathbf{W}_{agg}$ and $\mathbf{W}_{com}$.
+3.1. Node-based GNNs
+As shown in Figure 1 (a), node-based GNNs first learn latent representations for individual nodes, which then are used to form a graph-level representation.
+Generally, each GNN $f_{\theta}$ consists of two fundamental operations, i.e., AGG($\cdot$) and COM($\cdot$(Hamilton et al., 2017; Li and Chen, 2021), which can be parameterized by learnable matrices $\mathbf{W}_{agg}$ and $\mathbf{W}_{com}$, respectively.
 Formally, given a graph $G_{i}(V_{i},\mathbf{A}_{i},\mathbf{X}_{i})$ and a node $v\in V_{i}$, in the $l$-th layer, $\mathsf{AGG}^{(l)}(\cdot)$ in Equation (1) first aggregates the hidden representation $\mathbf{h}_{i}^{(l-1)}(u)$ and edge representation $\mathbf{e}_{uv}$ of each neighbor $u$ of node $v$ and obtain the aggregated result $\mathbf{m}^{(l)}_{i}(v)$.
-Then, as shown in Equation (3), $\mathsf{COM}^{(l)}(\cdot)$ combines the aggregated neighbor information $\mathbf{m}_{i}^{(l)}(v)$ and the latest information $\mathbf{h}_{i}^{(l-1)}(v)$ of node $v$ to obtain the $l$-th layer representation $\mathbf{h}_{i}^{(l)}(v)$.
+Subsequently, as shown in Equation (2), $\mathsf{COM}^{(l)}(\cdot)$ combines the aggregated neighbor information $\mathbf{m}_{i}^{(l)}(v)$ and the latest information $\mathbf{h}_{i}^{(l-1)}(v)$ of node $v$ to obtain the $l$-th layer graph-level representation $\mathbf{h}_{i}^{(l)}(v)$.
 (1) $\mathbf{m}_{i}^{(l)}(v)$ $=\mathsf{AGG}^{(l)}\left(\{(\mathbf{h}_{i}^{(l-1)}(u),\mathbf{e}_{uv}):u\in\mathcal{N}_{i}(v)\}\right)$
@@
 Model Optimization.
-Then, we can use a decoder (e.g., 2-layer MLPs (Zhu et al., 2021)) to predict discrete class or continuous property value for each graph $G_{i}$ based $\mathbf{h}_{i}$.
+Then, we can use a decoder (e.g., 2-layer Multi-Layer Perceptrons (MLPs) (Zhu and others, 2021)) to predict discrete class or continuous property value for each graph $G_{i}$ based $\mathbf{h}_{i}$.
 In summary, given the labeled training dataset $\mathcal{LG}=\{(G_{i},y_{i})\}_{i=1}^{n}$ and the prediction $y^{*}_{i}$ for each graph $G_{i}$ by the GNN $f_{\theta}$, the GNN $f_{\theta}$ can be optimized by minimizing the task loss as follows:
@@
 (5) $s.t.\quad$ $\mathcal{L}_{task}(y^{*}_{i},y_{i})=\begin{cases}-\sum_{y\in{y}_{i}}y\log y^{*}_{i}[y],&{y}_{i}\in\{0,1\}^{|\mathcal{Y}|}\\ \|y^{*}_{i}-y_{i}\|_{2},&y_{i}\in\mathbb{R}\end{cases}$
-2.2.2. Hierarchical Pooling-based Graph Neural Networks
+3.2. Hierarchical Pooling-based GNNs
 As shown in Figure 1 (b), hierarchical pooling (HP)-based GNNs are designed to capture hierarchical structures in the graph, which can reduce the graph size progressively while preserving its essential structural information.
 This is achieved by clustering nodes into groups and summarizing their features, allowing the model to focus on condensed graph structures and finally obtain the representation of each graph.
+Figure 1. The five types of current GNNs for graph-level tasks.
 As mentioned in Equations (1) and (2), at the $l$-th layer, node-based GNNs learn node representations using the adjacency matrix $\mathbf{A}_{i}$ and the node hidden representations $\mathbf{H}_{i}^{(l-1)}(V_{i})$.
 In contrast, at the $l$-th layer, HP-based GNNs first generate a new cluster assignment matrix $\mathbf{S}_{i}^{(l)}\in\{0,1\}^{n_{l-1}\times n_{l}}$, where $n_{l}<n_{l-1}$.
-This matrix maps the nodes $V_{i}^{(l-1)}$ from the $(l-1)$-th layer to a cluster index in $\{1,\ldots,n_{l}\}$.
+This matrix maps the nodes $V_{i}^{(l-1)}$ from the (l1)(l-1)-th layer to a cluster index in $\{1,\ldots,n_{l}\}$.
 The new adjacency matrix $\mathbf{A}_{i}^{(l)}\in\mathbb{R}^{n_{l}\times n_{l}}$ is then reconstructed based on the cluster assignment matrix $\mathbf{S}_{i}^{(l)}$ as follows:
@@
 Depending on the technique of generating the cluster matrix $\mathbf{S}_{i}^{(l)}$, existing pooling-based GNNs can be classified into three types.
-Table 1. Summary on important notations.
-Symbols Meanings
-$G_{i}(V_{i},\mathbf{A}_{i},\mathbf{X}_{i})$ The graph $G_{i}$
-$y_{i}$ Discrete label or continuous property value of $G_{i}$
-$N^{l}_{i}{(v)},N_{i}(v)$ The $l$-hop and 1-hop neighbors of node $v$ in $G_{i}$
-$D_{i}(v),\bar{D}_{i}$ Degree of node $v$ and average degree in $G_{i}$
-$f_{\theta}$ GNN model
-$l,L$ GNN layer index and the total layer number
-$\mathbf{h}^{(l)}_{i}(v)$ The $l$-th layer node representation of $v$ in $G_{i}$
-$\mathbf{h}_{i}(v)$ Final node representation of $v$ in $G_{i}$
-$\mathbf{H}^{(l)}_{i}(V_{j})$ The $l$-th layer representation of nodes $V_{j}$ in $G_{i}$
-$\mathbf{H}_{i}(V_{j})$ Final representation of nodes $V_{j}$ in $G_{i}$
-$\mathbf{h}_{i}$ Graph representation of $G_{i}$
-${y}^{*}_{i}$ Prediction of GNN $f_{\theta}$ for $G_{i}$
-$\mathbf{S}_{i}^{(l)}$ Cluster assignment matrix at $l$-th layer
-$\hat{G}^{*}_{i}(V_{i},\hat{\mathbf{A}}^{*}_{i},\hat{\mathbf{X}}^{*}_{i})$ The reconstructed graph for $G_{i}$
-$\tilde{G}_{i},\hat{G}_{i}$ Augmented positive views for $G_{i}$
-$s({\mathbf{h}}_{i},{\mathbf{h}_{j}})$ Similarity score between $G_{i}$ and $G_{j}$
-$imp_{v,u}$ Edge importance score between $v$ and $u$
-$\vec{t}_{v_{i}}^{k}\in\mathcal{T}_{v_{i}}^{k}$ $k$-tuple set of each node $v_{i}\in V$
-$G^{s}_{\vec{t}_{v_{i}}^{k}}$ $k$-path rooted subgraph based on $\vec{t}_{v_{i}}^{k}$
-$G^{s,a}_{\vec{t}_{v_{i}}^{k}}$ Feature augmented subgraph for $G^{s}_{\vec{t}_{v_{i}}^{k}}$
-$\mathcal{L}_{task}(\cdot)$ Task loss (Equation (5))
-$\mathcal{L}_{cy}(\cdot)$ Consistency-aware loss (Equation (23))
-Figure 1. The five types of current GNNs for graph-level tasks.
   • Similarity-based.
 Similarity-based approaches (Mesquita et al., 2020; Dhillon et al., 2007; Li et al., 2022b) cluster nodes based on the similarity of nodes features or structures.
-These methods rely on predefined similarity metrics or graph partitioning algorithms to generate the cluster assignment matrix.
+These methods rely on predefined similarity metrics such as cosine similarity of node features, or graph partitioning algorithms to generate the cluster assignment matrix.
 Graclus (Mesquita et al., 2020; Dhillon et al., 2007) and CC-GNN (Li et al., 2022b) assign nodes to different clusters based on feature similarity and graph structure.
@@
 EdgePool (Diehl, 2019) proposes to learn the edge score between each pair of connected nodes and then generate the cluster matrix accordingly.
-2.2.3. Subgraph-based Graph Neural Networks
+3.3. Subgraph-based GNNs
 Recent studies propose subgraph-based GNNs with enhanced expressive power, enabling them to more effectively capture substructure information.
@@
 MAGNA (Wang et al., 2021) proposes to learn more accurate weights between each pair of nodes based on all the paths between them within $k$-hops.
-2.2.4. Graph Learning-based Graph Neural Networks
-Due to uncertainty and complexity in data collection, graph structures may be redundant, biased, or noisy.
-Consequently, GNNs cannot learn reliable and accurate graph representations, leading to incorrect predictions.
-Therefore, as shown in Figure 1 (d), current researchers (Fatemi et al., 2023; Zhiyao et al., 2024; Li et al., 2024d) propose reconstructing the graph structure and node features to improve the quality of the learned graph representations.
+3.4. Graph Learning-based GNNs
+Due to uncertainty and complexity in data collection, real-world graphs often contain redundant, biased, or noisy edges and features.
+When operating on such imperfect structures, vanilla GNNs may learn spurious correlations and thus fail to produce reliable graph representations, ultimately leading to incorrect predictions.
+To mitigate this issue, as shown in Figure 1 (d), recent work (Fatemi et al., 2023; Zhiyao et al., 2024; Li et al., 2024d) propose to learn from purified or reconstructed graph structure and enhanced node features that better reflect the underlying signal to improve the quality of the learned graph representations.
 Given a labeled graph set $\mathcal{LG}=\{(G_{i},y_{i})\}_{i=1}^{n}$, the graph learning-based approaches can be formulated as bi-level optimization problem.
 (9) $\theta^{*}=\min_{\theta\in\Theta}{\frac{1}{|\mathcal{LG}|}\sum_{G_{i}\in\mathcal{LG}}\mathcal{L}_{task}({f}_{\theta},\hat{G}^{*}_{i}(V_{i},\hat{\mathbf{A}}^{*}_{i},\hat{\mathbf{X}}^{*}_{i}),y_{i})}.$
-(11) $s.t.\quad$ $\hat{\mathbf{A}}^{*}_{i},\hat{\mathbf{X}}^{*}_{i}=\arg\min_{\hat{\mathbf{A}}_{i},\hat{\mathbf{X}}_{i}}\mathcal{L}_{gl}(f_{\theta^{*}},\hat{G}_{i}(V_{i},\hat{\mathbf{A}}_{i},\hat{\mathbf{X}}_{i}),y_{i}),\forall G_{i}\in\mathcal{G}$
+$s.t.\quad$ $\hat{\mathbf{A}}^{*}_{i},\hat{\mathbf{X}}^{*}_{i}=\arg\min_{\hat{\mathbf{A}}_{i},\hat{\mathbf{X}}_{i}}\mathcal{L}_{gl}\bigl(f_{\theta^{*}},\hat{G}_{i}(V_{i},\hat{\mathbf{A}}_{i},\hat{\mathbf{X}}_{i}),y_{i}\bigr),$
+(10) $\hfill\forall G_{i}\in\mathcal{G}$
 At the low level in Equation (10), current approaches propose different graph learning objectives $\mathcal{L}_{gl}(\cdot)$ to reconstruct graph structure $\mathbf{A}_{i}$ and node features $\mathbf{X}_{i}$.
@@
 Depending on the techniques of reconstructing graphs, current GL-based GNNs can be categorized into three types.
-  • Preprocessing-based.
-Preprocessing-based approaches (Wu et al., 2019; Li et al., 2022a; Entezari et al., 2020) reconstruct graphs and then use them to optimize GNNs directly.
+  • Preprocessing-based. Preprocessing-based approa-
+    ches (Wu et al., 2019; Li et al., 2022a; Entezari et al., 2020) reconstruct graphs and then use them to optimize GNNs directly.
 The basic idea is to first reveal the common graph patterns and modify the graph structure and node features to recover these patterns.
@@
 For instance, several approaches (Jin et al., 2021, 2020b; Li et al., 2024b; Franceschi et al., 2019; Sun et al., 2022; Luo et al., 2021), such as ADGNN (Li et al., 2024b), ProGNN (Jin et al., 2020b), and SimPGCN (Jin et al., 2021), propose reconstructing the edges between each pair of nodes by minimizing the GNN loss on the reconstructed structure while also reducing the rank of the adjacency matrix.
-Alternatively, several approaches (Zhang et al., 2019; Zhou et al., 2024; Wang et al., 2023), such as MOSGSL (Zhou et al., 2024) and HGP-SL (Zhang et al., 2019), focus on first partitioning each graph into subgraphs based on node similarities and predefined motifs, and then reconstructing the edges between these subgraphs rather than at the individual node level.
-2.2.5. Self Supervised Learning-based Graph Neural Networks
+Alternatively, several approaches (Zhang et al., 2019; Zhou et al., 2024; Wang et al., 2023a), such as MOSGSL (Zhou et al., 2024) and HGP-SL (Zhang et al., 2019), focus on first partitioning each graph into subgraphs based on node similarities and predefined motifs, and then reconstructing the edges between these subgraphs rather than at the individual node level.
+3.5. Self Supervised Learning-based GNNs
 Self-supervised learning (SSL) has become a powerful paradigm to pretrain GNNs without the need for labeled data, which can capture the node patterns and graph patterns.
@@
 where the task loss $\mathcal{L}_{task}(\cdot)$ is defined in Equation (4).
-Depending on the technique of generating supervised signals from unlabeled graphs, SSL-based GNN approaches can be broadly categorized into two main types.
+Depending on the specific technique used to auto-generate supervised signals from unlabeled graphs, SSL-based GNN approaches can be broadly categorized into two main paradigms.
   • Pretext Task-based.
-Pretext task-based approaches (Hou et al., 2022; Inae et al., 2023; Zhang et al., 2021; Zang et al., 2023; Hu et al., 2020a; Jin et al., 2020a; Wang et al., 2022b) design auxiliary tasks to help the model learn useful representations from the graph structure and features without requiring external labels.
+Pretext task-based approaches  (Hou et al., 2022; Inae et al., 2023; Zhang et al., 2021; Zang et al., 2023; Hu et al., 2020b; Jin et al., 2020a; Wang et al., 2022b) design auxiliary tasks to help the model learn useful representations from the graph structure and features without requiring external labels.
 Current approaches design diverse node-level and graph-level tasks to capture intrinsic graph patterns (Wang et al., 2022b), such as predicting node attribute, node degree, and the node number.
@@
 MoAMa (Inae et al., 2023) first masks node features and then uses GNNs to reconstruct the node features based on the masked graphs.
-GraphMAE (Hou et al., 2022) and GPTGNN (Hu et al., 2020a) predict both node attributes and edges.
+GraphMAE (Hou et al., 2022) and GPTGNN (Hu et al., 2020b) predict both node attributes and edges.
   • Graph Contrastive Learning-based.
-Graph contrastive learning (GCL)-based GNNs (Wang et al., 2022c; Perozzi et al., 2014; Lee et al., 2022; Hassani and Khasahmadi, 2020; Yuan et al., 2021; Sun et al., 2024) aim to learn representations by contrasting positive and negative examples.
+Graph contrastive learning (GCL)-based GNNs (Wang et al., 2022a; Perozzi and others, 2014; Lee et al., 2022; Hassani and Khasahmadi, 2020; Yuan et al., 2021; Sun et al., 2024) aim to learn representations by contrasting positive and negative examples.
 The basic idea is to maximize the similarity between different views or augmentations of the same node or graph (positive pairs) while minimizing the similarity between different nodes or graphs (negative pairs).
@@
     (13) $\theta^{\prime}=\arg\min_{\theta}{\frac{1}{|\mathcal{UG}|}\sum_{G_{i}\in\mathcal{UG}}\mathcal{L}_{cl}(f_{\theta},\hat{G}_{i},\tilde{G}_{i},Neg_{i})}.$
-    (15) $s.t.$ $\tilde{G}_{i},\hat{G}_{i}=\arg\min_{\tilde{G}_{i},\hat{G}_{i}}{\mathcal{L}_{positive}(G_{i},\mathcal{T})},\forall G_{i}\in\mathcal{UG},$
+    (14) $s.t.\$$\tilde{G}_{i},\hat{G}_{i}=\arg\min_{\tilde{G}_{i},\hat{G}_{i}}{\mathcal{L}_{positive}(G_{i},\mathcal{T})},\forall G_{i}\in\mathcal{UG},$
     where $\mathcal{L}_{cl}(\cdot)$ is the contrastive loss.
-Additionally, $\mathcal{L}_{positive}(\cdot)$ is the view generation loss used to generate two positive views (i.e., $\tilde{G}_{i}$ and $\hat{G}_{i}$ for each graph $G_{i}$) where $\mathcal{T}$ is a set of augmentation operations for generating the views.
-Negative views ($Neg_{i}$) are a set of graphs, which are typically randomly sampled from all graphs or from graphs with labels different from $G_{i}$.
-One typical $\mathcal{L}_{cl}(f_{\theta},\hat{G}_{i},\tilde{G}_{i},Neg_{i})$ in Equation (14) can be defined based on InfoNCE loss (Zhu et al., 2021; Yeh et al., 2022) as follows:
+Additionally, $\mathcal{L}_{positive}(\cdot)$ denotes the view generation loss used to generate two positive views (i.e., $\tilde{G}_{i}$ and $\hat{G}_{i}$ for each graph $G_{i}$) where $\mathcal{T}$ is a set of augmentation operations for generating the views.
+Negative views ($Neg_{i}$) are a set of graphs, which are typically randomly sampled from all graphs or from graphs whose labels differ from $G_{i}$.
+One typical of the contrastive loss $\mathcal{L}_{cl}(f_{\theta},\hat{G}_{i},\tilde{G}_{i},Neg_{i})$ in Equation (13) can be defined based on InfoNCE loss (Zhu and others, 2021; Yeh et al., 2022) as follows:
     (15) $\mathcal{L}_{cl}(\cdot)=-\log\frac{\text{s}(\mathbf{\hat{h}}_{i},\mathbf{\tilde{h}}_{i})}{\sum_{G^{\prime}_{i}\in\{\hat{G}_{i},\tilde{G}_{i}\}}\sum_{G_{j}\in A(G_{i})\}}s(\mathbf{\hat{h}}^{\prime}_{i},\mathbf{\tilde{h}}_{j})},$
@@
     The techniques in Equation (14) are employed to generate positive views for each node.
-Specifically, similarity-based methods (Wang et al., 2022c; Perozzi et al., 2014; Lee et al., 2022) utilize both node features and graph structures to identify nodes in the original graph, which are similar to a given target node, as the positive nodes.
+Specifically, similarity-based methods (Wang et al., 2022a; Perozzi and others, 2014; Lee et al., 2022) utilize both node features and graph structures to identify nodes in the original graph, which are similar to a given target node, as the positive nodes.
 Also, diffusion-based approaches (Hassani and Khasahmadi, 2020; Yuan et al., 2021; Sun et al., 2024) modify the graph’s structure (e.g., by deleting or adding edges) based on global topological information, such as personalized PageRank (PPR) (Haveliwala, 1999) and motif-preserving (Sun et al., 2024).
-Perturbation-based methods (Thakoor et al., 2021; Zhu et al., 2020; Zhu et al., 2021; Li et al., 2024a), on the other hand, introduce changes to both the graph’s structure and node features.
-3. A Generalizable and Expressive GNN
-We propose a simple yet effective contrastive learning-based GNN with high generalization and expressivity capabilities on graph-level tasks.
-As shown in Figure 1, our proposed GNN model consists of three components, i.e., adaptive graph augmentation, $k$-Path rooted subgraph encoder, and consistency-aware model training.
-Figure 2. Framework overview.
-3.1. Adaptive Graph Augmentation
-To enhance the generalization ability of the GNN model, we propose using contrastive learning to optimize GNNs for graph-level tasks.
-As described in Section 2, the key step is to generate positive graph pairs for each graph.
-Here, we propose a unified contrastive learning approach for graphs in diverse domains, such as geometric graphs and biology.
-The basic idea is to first measure the importance of each edge concerning the whole graph and then drop the unimportant edges based on weighted sampling to generate positive pairs for $G$.
-Specifically, given a graph $G_{i}(V_{i},\mathbf{A}_{i},\mathbf{X}_{i})$ with edges $E$, the edge importance of each edge $e_{v,u}\in E_{i}$ is calculated based on node degree of $v$ and $u$ (Li et al., 2024a; Zhu et al., 2020), i.e., $imp_{v,u}=\log\left(\frac{D_{i}(v)+D_{i}(u)}{2}+1\right)$, where edge connected with higher-degree nodes tends to be important because nodes with higher degrees in a graph are bridges and core nodes in a functional motif as follows:
-  • Geometric Graphs. High-degree nodes act as crucial bridges, and removing their edges significantly impacts the number of paths and cycles within this graph.
-  • Social Networks.
-Higher-degree nodes are crucial as they often represent key influencers or hubs in the network.
-These nodes connect different parts of the network, facilitating communication and information flow.
-Their removal can fragment the network and disrupt connectivity significantly.
-  • Chemistry Graphs.
-High-degree atoms mainly determine the properties of a motif.
-For example, phosphorus (P) in phosphate groups ($\text{PO}_{4}$) is bonded to four oxygen atoms (O).
-While oxygen atoms contribute to properties like electronegativity and polarity, phosphorus is often more critical in determining the overall functionality and behavior of the phosphate group.
-Then, the dropping probability $p_{v,u}$ of each edge $e_{v,u}$ can be normalized using min-max normalization (Zhu et al., 2020; Patro, 2015; Li et al., 2024a) as follows:
-(17) $p_{{v,u}}=\mu\cdot\frac{imp_{max}-imp_{v,u}}{imp_{max}-imp_{min}},$
-where $\mu\in[0,1]$ is a hyperparameter, and $imp_{max}=\max_{e_{i,j}\in E_{i}}imp_{i,j}$ and $imp_{min}=\min_{e_{i,j}\in E_{i}}imp_{i,j}$ are the largest and smallest edge score, respectively.
-Then, for each graph $G_{i}$, we can generate two positive graphs $\hat{G}_{i}$ and $\tilde{G}_{i}$ by randomly dropping unimportant edges based on Equation (17).
-3.2. $k$-Path Rooted Subgraph Encoder
-A more expressive GNN (Xu et al., 2019; Huang et al., 2022).
-can better encode substructures like paths, cycles, and other motifs, which are essential for solving downstream tasks.
-To enhance the expressivity ability, we propose a $k$-Path rooted subgraph encoder following (Huang et al., 2022; You et al., 2021).
-3.2.1. $k$-Path Rooted Subgraph and Feature Augmentation
-Given a graph $G(V,\mathbf{A},\mathbf{X})$, we define a $k$-tuple set of each node $v_{i}\in V$ as $\mathcal{T}_{v_{i}}^{k}=\{\vec{t}_{v_{i}}^{k}=(v_{i},v_{i+1},v_{i+2},\ldots,v_{i+k-1})\}$, where there is an edge between each consecutive pair $(v_{i},v_{i+1})$ and $v_{i+j}$ is actually a $j$-hop neighbor of node $v_{i}$.
-Particularly, for each node $v_{i}\in V$, the number of $k$-tuples starting from $v_{i}$ is $O(\overline{n}_{nei}^{k})$, where $\overline{n}_{nei}$ is the average neighbor number of nodes in $G$.
-Then, for each tuple in $\vec{t}_{v_{i}}^{k}\in\mathcal{T}_{v_{i}}^{k}$, we can extract a subgraph based on nodes in $\vec{t}_{v_{i}}^{k}$.
-Formally, given a $k$-tuple nodes $\vec{t}_{v_{i}}^{k}=(v_{i},v_{i+1},v_{i+2},\ldots,v_{i+k-1})$ and the graph $G$, we denote the rooted subgraph of $\vec{t}_{v_{i}}^{k}$ as $G^{s}_{\vec{t}_{v_{i}}^{k}}(V^{s}_{\vec{t}_{v_{i}}^{k}},\mathbf{A}^{s}_{\vec{t}_{v_{i}}^{k}},\mathbf{X}^{s}_{\vec{t}_{v_{i}}^{k}})$, which extracts $L$-hop neighbors of each node in $\vec{t}_{v_{i}}^{k}$, i.e., $V^{s}_{\vec{t}_{v_{i}}^{k}}=\cup_{v_{j}\in\vec{t}_{v_{i}}^{k}}{\cup_{l=0}^{L}\mathcal{N}^{l}(v_{j})}$ and keep the edges among nodes $V^{s}_{\vec{t}_{v_{i}}^{k}}$ in $G$.
-Then, given each rooted subgraph $G^{s}_{\vec{t}_{v_{i}}^{k}}(V^{s}_{\vec{t}_{v_{i}}^{k}},\mathbf{A}^{s}_{\vec{t}_{v_{i}}^{k}},\mathbf{X}^{s}_{\vec{t}_{v_{i}}^{k}})$ of nodes $\vec{t}_{v_{i}}^{k}$, we add identifiers for each node in $\vec{t}_{v_{i}}^{k}$ based on the consecutive order.
-Specifically, for each node $v_{i+j}\in\vec{t}_{v_{i}}^{k}$, its node feature $\mathbf{x}(v_{i+j})\in\mathbb{R}^{d_{x}}$ is augmented as $\mathbf{x}(v_{i+j})=\mathbf{x}(v_{i+j})\oplus\mathbf{e}_{j,k}\in\mathbb{R}^{(d_{x}+k)}$, where $\mathbf{e}_{j,k}\in\{0,1\}^{k}$ is a one hot vector, where the $j$-th element is 1 and others are 0.
-For nodes $v_{m}\in V^{s}_{\vec{t}_{v_{i}}^{k}}\setminus\vec{t}_{v_{i}}^{k}$, its node feature $\mathbf{x}(v_{m})\in\mathbb{R}^{d_{x}}$ of is augmented as $\mathbf{x}(v_{m})=\mathbf{x}(v_{m})\oplus\mathbf{0}_{k}\in\mathbb{R}^{(d_{x}+k)}$, where $\mathbf{0}_{k}$ is a vector of length $k$ and all its components are 0.
-Then, we denote the feature augmented subgraph $G^{s}_{\vec{t}_{v_{i}}^{k}}$ as $G^{s,a}_{\vec{t}_{v_{i}}^{k}}(V^{s}_{\vec{t}_{v_{i}}^{k}},\mathbf{A}^{s}_{\vec{t}_{v_{i}}^{k}},\mathbf{X}^{s,a}_{\vec{t}_{v_{i}}^{k}})$.
-3.2.2. Graph Representation Learning
-Given each augmented rooted subgraph $G^{s,a}_{\vec{t}_{v_{i}}^{k}}$, the node representations of nodes $V^{s}_{\vec{t}_{v_{i}}^{k}}$ can be learned by a GNN encoder as $\mathbf{H}({V^{s}_{\vec{t}_{v_{i}}^{k}}})$, such as GraphSAGE (Hamilton et al., 2017) or GatedGNN (Li et al., 2016).
-Then, the representations of each node $v_{i}$ can be learned based on all subgraphs of each $\vec{t}_{v_{i}}^{k}\in\mathcal{T}_{v_{i}}^{k}$ as follows:
-(18) $\mathbf{h}({v_{i}})=\mathsf{READOUT}_{1}(\{\mathbf{H}(V^{s}_{\vec{t}_{v_{i}}^{k}})[v_{i}]|\vec{t}_{v_{i}}^{k}\in\mathcal{T}_{v_{i}}^{k}\}),$
-Where $\mathsf{READOUT}_{1}(\cdot)$ is a readout function, such as MaxPooling. Therefore, the final graph representation of $G(V,\mathbf{A},\mathbf{X})$ can be learned as follows:
-(19) $\mathbf{h}=\mathsf{READOUT}_{2}(\{\mathbf{h}({v_{i}})|v_{i}\in V\}),$
-The graph representation $\mathbf{h}$ can be used to predict graph labels or graph properties, as mentioned in Equation (1).
-3.2.3. Theoretical Analysis
-Theorem 3.1.
-Given the graph $G(V,\mathbf{A},\mathbf{X})$, the $k$-Path rooted subgraph of the $k$-tuple node set $\mathcal{T}_{v_{i}}^{k}$ can count the number of $k+2$-Path, starting from the node $v_{i}$.
-Proof.
-We firstly give Lemma 3.2 and then prove Theorem 3.1.
-Lemma 3.2.
-Given a graph denoted as $G(V,\mathbf{A},\mathbf{X})$ and a $k$-tuple $\vec{t}_{v_{i}}^{k}=(v_{i},v_{i+1},v_{i+2},\ldots,v_{i+k-1})$ starting from node $v_{i}$, our model can count the number of $(k+2)$-Path between $v_{i}$ and any node $v_{j}\notin\vec{t}^{k}_{v_{i}}$ that pass through the path $(v_{i},v_{i+1},v_{i+2},\ldots,v_{i+k-1})$.
-Proof.
-Note that since we extract the subgraphs based on the $k$ tuple $\vec{t}_{v_{i}}^{k}$, we can know whether a node belong to this $\vec{t}_{v_{i}}^{k}$.
-Given a node $v_{j}\in V$ and $v\notin\vec{t}^{k}_{v_{i}}$, we can count the $k+2$-Path between $v_{i}$ and $v_{j}$ by aggregating information from neighbor.
-Firstly, we can identify whether a node $v_{j}$ is the one-hop neighbor of the last element of $\vec{t}_{v_{i}}^{k}$ (i.e., $v_{i+k-1}$) as follows:
-(20) $\mathbf{h}^{(1)}_{\vec{t}_{v_{i}}^{k}}(v_{j})=\mathbb{I}(v_{j}\notin\vec{t}^{k}_{v_{i}}[1:k-1])\sum_{v_{m}\in N(v_{j})}{\mathbb{I}(v_{m}=v_{i+k-1})}$
-where $\vec{t}_{v_{i}}^{k}[1:k-1]=(v_{i},v_{i+1},\cdots,v_{i+k-2})$ is the first $k-1$ elements of $\vec{t}_{v_{i}}^{k}$, the indicator function $\mathbb{I}(condition)=1$ if the $condition$ is true.
-Specifically, in Equation (20), if $v_{j}$ is the one-hop neighbor $v_{i+k-1}$, the hidden representation $\mathbf{h}^{(1)}_{\vec{t}_{v_{i}}^{k}}(v_{j})=1$ otherwise $\mathbf{h}^{(1)}_{\vec{t}_{v_{i}}^{k}}(v_{j})=0$.
-Then, in Equation (21), for each node $v_{j}\notin\vec{t}^{k}_{v_{i}}$, $\mathbf{h}^{(2)}_{\vec{t}_{v_{i}}^{k}}(v_{j})$ indicate the number of 2-Path from node $v_{i+k-1}$ to the node $v_{j}$.
-(21) $\mathbf{h}^{(2)}_{\vec{t}_{v_{i}}^{k}}(v_{j})=\mathbb{I}(v_{j}\notin\vec{t}^{k}_{v_{i}})\sum_{v_{m}\in N(v_{j})}{\mathbf{h}^{(1)}_{\vec{t}_{v_{i}}^{k}}(v_{m})}$
-Finally, in Equation (22), for each node $v_{j}\notin\vec{t}^{k}_{v_{i}}$, $\mathbf{h}^{(2)}_{\vec{t}_{v_{i}}^{k}}(v_{j})$ indicate the number of 3-Path from node $v_{i+k-1}$ to the node $v_{j}$.
-(22) $\mathbf{h}^{(3)}_{\vec{t}_{v_{i}}^{k}}(v_{j})=\mathbb{I}(v_{j}\notin\vec{t}^{k}_{v_{i}})\sum_{v_{m}\in N(v_{j})}{(\mathbf{h}^{(2)}_{\vec{t}_{v_{i}}^{k}}(v_{m})-\mathbf{h}^{(1)}_{\vec{t}_{v_{i}}^{k}}(v_{j}))},,$
-If $v_{j}$ is a one-hop neighbor of $v_{i+k-1}$, there exists a recursive 3-Path of the form $(v_{i+k-1},v_{j},v_{m},v_{j})$.
-To eliminate these recursive 3-Path for $v_{j}$, we use $\mathbf{h}^{(2)}_{\vec{t}_{v_{i}}^{k}}(v_{m})-\mathbf{h}^{(1)}_{\vec{t}_{v_{i}}^{k}}(v_{j})$.
-Since $\vec{t}^{k}_{v_{i}}$ is a $k-1$ path from node $v_{i}$ to node $v_{i+k-1}$ and $\mathbf{h}^{(3)}_{\vec{t}_{v_{i}}^{k}}(v_{j})$ is the number of 3-Path from $v_{i+k-1}$ to $v_{j}$, $\mathbf{h}^{(3)}_{\vec{t}_{v_{i}}^{k}}(v_{j})$ is the number of $k+2$-Path from $v_{i}$ to $v_{j}$ that pass through the path $\vec{t}^{k}_{v_{i}}$.
-Based on Lemma 3.2, given the graph $G(V,\mathbf{A},\mathbf{X})$, we can compute the total number of $k+2$ path in the graph $G$ that starts from node $v_{i}$ as follows: $n^{k+2}_{v_{i}}=\sum_{\vec{t}^{k}_{v_{i}}\in\mathcal{T}^{k}_{v_{i}}}\sum_{v_{j}\in V}{\mathbf{h}^{(3)}_{\vec{t}_{v_{i}}^{k}}(v_{j})}.$
-Theorem 3.3.
-Given the graph $G(V,\mathbf{A},\mathbf{X})$, the $k$-Path rooted subgraph of the $k$-tuple node set $\mathcal{T}_{v_{i}}^{k}$ can count the number of $k+3$-Cycle of the node $v_{i}$.
-Proof.
-Based on the Theorem 3.1, we can counting the number of $k+2$-Path from each node $v_{i}$ that passes through the path $\vec{t}^{k}_{v_{i}}\in\mathcal{T}^{k}_{v_{i}}$, then we can cunt the number of $k+3$-Cycle of node $v_{i}$ based on its neighbor $N(v_{i})$ as $\sum_{v_{j}\in N(v_{i})}\sum_{\vec{t}^{k}_{v_{i}}\in\mathcal{T}^{k}_{v_{i}}}\mathbf{h}^{(3)}_{\vec{t}^{k}_{v_{i}}}$.
-∎
-3.3. Consistency-aware Model Training
-We propose a contrastive loss and a task-aware consistency loss to optimize GNNs, enhancing the generalizability of learned representations and effectively capturing graph patterns.
-Given a batch of graphs $\mathcal{G}_{B}=\{(G_{i},y_{i})\}_{i=1}^{B}$ with size $B$, we generate two positive views $\hat{G}_{i}$ and $\tilde{G}_{i}$ for each $G_{i}\in\mathcal{G}_{B}$ by dropping edges from $G_{i}$ based on Equation (17).
-The contrastive loss $\mathcal{L}_{cl}(f_{\theta},G_{i})$ for each graph $G_{i}$ is defined as follows:
-$-\log\frac{\sum_{\{G_{j}|y_{j}=y_{i}\}}\sum_{\mathbf{h}^{\prime}_{i}\in\{\hat{\mathbf{h}}_{i},\tilde{\mathbf{h}}_{i}\}}\sum_{\mathbf{h}^{\prime}_{j}\in\{\hat{\mathbf{h}}_{j},\tilde{\mathbf{h}}_{j}\}}{s({\mathbf{h}}^{\prime}_{i},{\mathbf{h}}^{\prime}_{j})+s(\hat{\mathbf{h}}_{i},\mathbf{\tilde{h}}_{i})}}{\sum_{G_{i},G_{j}\in\mathcal{G}_{B}}\sum_{\mathbf{h}^{\prime}_{i}\in\{\hat{\mathbf{h}}_{i},\tilde{\mathbf{h}}_{i}\}}\sum_{\mathbf{h}^{\prime}_{i}\in\{\hat{\mathbf{h}}_{j},\tilde{\mathbf{h}}_{j}\}}{s(\hat{\mathbf{h}}_{i},\hat{\mathbf{h}}_{j})}},$
-where $\{G_{j}|y_{j}=y_{i}\}$ denotes the graphs in $\mathcal{G}_{B}$ that has the same label of $G_{i}$, and $s({\mathbf{h}}_{i},{\mathbf{h}_{j}})=\exp(cosine({\mathbf{h}}_{i},{\mathbf{h}}_{j})/\tau)$ is the similarity score between ${\mathbf{h}}_{i}$ and ${\mathbf{h}}_{j}$ based the temperature parameter $\tau$.
-Also, we define a task-aware consistency loss.
-Note that two positive pairs $\hat{G}_{i}$ and $\tilde{G}_{i}$ are created by dropping unimportant edges from $G_{i}$.
-For graph classification tasks, the graph label of $\hat{G}_{i}$ and $\tilde{G}_{i}$ should remain similar to the original graph $G_{i}$.
-For graph regression tasks, such as counting tasks, the cycle (resp.
-path ) number in $\hat{G}_{i}$ and $\tilde{G}_{i}$ should be less than that in $G_{i}$, since dropping edges decreases the cycle and path numbers in $G_{i}$.
-Based on this analysis, we define the task-aware consistency loss $\mathcal{L}_{cy}(f_{\theta},\hat{G}_{i},\tilde{G}_{i},y_{i})$ for each graph $G_{i}$ as follows:
-(23) $\mathcal{L}_{cy}(\cdot)=\begin{cases}-\sum_{y\in{y}_{i}}(\log{\hat{y}^{*}}_{i}[y]+\log{\tilde{y}^{*}}_{i}[y]),&{y}_{i}\in\{0,1\}^{|\mathcal{Y}|}\\ \mathsf{ReLU}(\hat{y}^{*}_{i}-y_{i})+\mathsf{ReLU}(\tilde{y}^{*}_{i}-y_{i}),&y_{i}\in\mathbb{R}\end{cases}$
-In summary, we pretrain the GNNs by the combination of contrastive learning loss and task-aware consistency loss.
-$\theta^{\prime}=\arg\min_{\theta}$ $\frac{1}{B}\sum_{G_{i}\in\mathcal{G}_{B}}\mathcal{L}_{cl}(f_{\theta},\hat{G}_{i},\tilde{G}_{i},Neg_{i})$
-(24) $+\mathcal{L}_{cy}(f_{\theta},\hat{G}_{i},\tilde{G}_{i},y_{i})$
-where $\mathcal{L}_{cl}(\cdot)$ is the contrastive loss in Equation (16), and $\mathcal{L}_{cy}(\cdot)$ is the task-aware consistency loss in Equation (23).
-Then, we can fine-tune the GNNs with labeled data $\mathcal{LG}$ based on the task-specific loss in Equation (5).
+Perturbation-based methods (Thakoor et al., 2021; Zhu et al., 2020; Zhu and others, 2021; Li et al., 2024a), on the other hand, introduce changes to both the graph’s structure and node features.
 4. Experiments
-We conduct experiments to analyze baselines and our model on both graph classification and regression tasks.
-Specifically, we introduce datasets and baselines in Section 4.1 and Section 4.2.
-We present the hyperparameter setting and the evaluation metric in Section 4.3 and Section 4.4, respectively.
-Effectiveness and efficiency are discussed in Section 4.5 and Section 4.6.
-Additionally, we evaluate baselines on more scenarios in Section 4.7.
+4.1. Evaluation Framework OpenGLT
+As shown in Figure 2, we design a flexible and end-to-end evaluation framework OpenGLT that consists of three key components: the data-level, model-level, and evaluation-level components.
+It ensures fair and comprehensive comparisons across different methods.
+  • Data-level Component.
+We collected 13 datasets for graph classification spanning diverse domains, including social networks, chemical graphs, and molecular graphs, as well as 13 datasets for graph regression like cycle and path counting.
+Additionally, we provide a unified data processing pipeline, which includes data transformation and standardized data splits tailored to various scenarios, such as imbalanced and few-shot settings.
+All datasets are formatted into a unified loader to facilitate fair comparisons across different methods and support future extensions.
+  • Model-level Component.
+At the model level, we provide more than 20 GNN models to support comprehensive evaluations.
+These include node-based models, pooling-based models, subgraph-based models, graph-level models (GL-based), and self-supervised learning-based models (SSL-based).
+Additionally, we incorporate various optimization strategies tailored to GNNs, such as neighbor sampling, subgraph sampling, and graph partitioning, to enhance scalability and performance.
+  • Evaluation-level Component.
+It focuses on ensuring fair and comprehensive comparisons of model performance.
+We utilize a wide range of metrics, including Macro-F1, Micro-F1, Mean Absolute Error (MAE), and $R^{2}$ scores, as well as resource-based measures such as runtime and memory consumption.
+To aid in the interpretation of results, we provide multiple visualization options, including tables, column charts, and line charts, enabling clear and insightful comparisons across different methods.
+Our framework is designed to be user-friendly and highly extensible.
+Users can easily integrate new GNN models by following a standardized interface and registering them in the model library.
+Similarly, new datasets can be incorporated by formatting them into our unified loader, which handles preprocessing and splits automatically.
+The evaluation module also supports the addition of custom metrics and visualizations.
+To assist users, we provide comprehensive documentation with examples, ensuring that the framework can be readily extended to meet diverse research needs.
+Figure 2. Evaluation framework.
 4.2. Datasets
-We use graphs from four domains, including social networks (SN), biology (BIO), chemistry (CHE), and geometric (GE) graphs.
-In particular, for dataset splitting, if a standard split is publicly available and commonly used in existing works, we adopt the standard split.
+We evaluate GNNs on four domains with 26 datasets: social networks (SN), biology (BIO), chemistry (CHE), and motif counting (MC).
+For dataset splitting, if a standard split is publicly available and commonly used in existing works, we adopt the standard split.
 If there is no standard split, we use a 10-fold cross-validation approach.
-The detailed data statistics and split ratio are in Table 2.
+Detailed statistics and description are in Table 3.
 4.2.1. Social Networks
-  • IMDB-B (Morris et al., 2020) and IMDB-M (Morris et al., 2020).
-Two movie datasets from the IMDB (imd, 2024), where nodes are actors, edges denote co-appearances, and labels indicate movie genres.
-IMDB-B and IMDB-M are binary and multiple-class classification datasets, respectively.
-  • REDDIT-B (Morris et al., 2020).
-A binary social network dataset from Reddit (red, 2024), where nodes represent users, edges indicate interactions, and graphs are labeled as QA-based or discussion-based.
-  • COLLAB-M (Morris et al., 2020). A multi-class scientific collaboration network, where nodes are researchers and edges indicate co-authorship. Each graph is labeled with different research fields.
+  • IMDB-BINARY (I-B) (Morris et al., 2020) and IMDB-MULTI (I-M) (Morris et al., 2020).
+Two movie datasets from IMDB (IMDb, 2024) where each graph represents a movie, with nodes as actors and edges indicating co-appearance.
+The task of IMDB-B is to classify movies into two genres: Action and Romance.
+IMDB-M extends this to a three-class setting: Comedy, Action, and Drama.
+  • REDDIT (RED) (Morris et al., 2020).
+We use the binary version (REDDIT-BINARY) of a dataset from Reddit (Reddit, 2024), where each graph represents a discussion thread.
+Nodes correspond to users, and edges indicate interactions.
+The task is to classify threads into two categories: Q&A-based and Discussion-based.
+  • COLLAB (COL) (Morris et al., 2020).
+A scientific collaboration dataset where each graph represents an ego-network of a researcher, with nodes as researchers and edges indicating co-authorship.
+The task is to classify the networks into one of three research fields: High Energy Physics, Condensed Matter Physics, or Astrophysics.
 4.2.2. Biology
-  • PROTEIN-B (Morris et al., 2020). A binary protein dataset where nodes represent secondary structure elements, edges denote proximity, and graphs indicate if the molecule is an enzyme.
-  • DD-B (Morris et al., 2020).
-Derived from a molecular protein dataset, where nodes represent amino acids, and edges indicate adjacency.
-Each graph is labeled to classify different types of proteins.
-  • ENZYMES (Morris et al., 2020). Derived from the BRENDA enzyme database, it represents nodes as secondary structural elements, edges as adjacency, and labels each graph with one of six EC classes.
-  • MolHIV (Hu et al., 2020b) and MolTox21 (Hu et al., 2020b).
-They are molecular datasets, where nodes represent atoms, and edges are chemical bonds.
-Labels in MolHIV indicate whether the molecule has HIV inhibition, and labels in MolTox21 indicate the toxicity type.
+  • PROTEINS (PRO) (Morris et al., 2020) and DD (Morris et al., 2020).
+Two binary classification datasets where each graph represents a protein.
+The task is to classify a protein molecule as either an enzyme or a non-enzyme.
+In PROTEINS, nodes represent secondary structure elements (SSEs), while in DD, nodes are amino acids; edges in both indicate spatial proximity.
+  • ENZYMES (ENZ) (Morris et al., 2020).
+A six-class classification dataset from the BRENDA enzyme database.
+Each graph represents a protein, in which nodes represent SSEs with features encoding atomic properties, and edges denote spatial proximity.
+The task is to classify proteins into one of six enzyme commission (EC) classes.
+  • MolHIV (HIV) (Hu et al., 2020a) and MolTox21 (TOX) (Hu et al., 2020a).
+Two molecular property prediction datasets where nodes represent atoms and edges denote chemical bonds.
+MolHIV is a single-label binary dataset and the task is to predict whether a molecule inhibits HIV replication.
+MolTox21 is a 12-label binary dataset, each label represents a specific toxicological target.
 4.2.3. Chemistry
-MUTAG, NCI1, MolBACE, and MolPCBA are molecular graph datasets, with nodes represent atoms, node labels indicating the atomic species (e.g., N, C, O), and edges represent chemical bonds between atoms.
-  • MUTAG (Morris et al., 2020). The labels are whether a molecule has mutagenic activity, referring to its potential to cause genetic mutations.
-  • NCI1 (Morris et al., 2020). The labels specify whether a molecule shows a positive response in assays for specific anticancer activity.
-  • MolBACE (Hu et al., 2020b). The labels indicate whether a molecule inhibits the BACE enzyme (Beta-secretase 1).
-  • MolPCBA (Hu et al., 2020b). The labels represent the activity of a molecule across 128 different bioassays, indicating whether the compound is active in a given experimental condition.
-Table 2. Data statistics. CV denotes Cross-Validation.
-Type Data Graphs $\overline{\mbox{Nodes}}$ $\overline{\mbox{Edges}}$ Feat Classes Split Ratio
-SN IMDB-B 1,000 19.77 96.53 N/A 2 10-fold CV
-IMDB-M 1,500 13.00 65.94 N/A 3 10-fold CV
-REDDIT-B 2,000 429.63 497.75 N/A 2 10-fold CV
-COLLAB 5,000 74.49 2,457.78 N/A 3 10-fold CV
-BIO PROTEINS 1,113 39.06 72.82 3 2 10-fold CV
-DD 1,178 284.32 715.66 89 2 10-fold CV
-ENZYMES 600 32.63 62.14 3 6 10-fold CV
-MolHIV 41,127 25.50 54.93 9 2 8/1/1
-MolTox21 7,831 18.90 39.22 9 12 8/1/1
-CHE MUTAG 188 17.93 19.79 7 2 10-fold CV
-NCI1 4,110 29.87 32.30 37 2 10-fold CV
-MolBACE 1,513 34.13 36.91 204 2 8/1/1
-MolPCBA 437,929 26.00 28.10 128 128 8/1/1
-GE $\{\mathbf{3,4,5}\}$-Cycle 5,000 18.80 31.34 1 1 3/2/5
-$\mathbf{\{6,7,8\}}$-Cycle 5,000 18.80 31.34 1 1 3/2/5
-$\mathbf{\{4,5,6\}}$-Path 5,000 18.80 31.34 1 1 3/2/5
-4-Clique 5,000 18.80 31.34 1 1 3/2/5
-Tailed Tri. 5,000 18.80 31.34 1 1 3/2/5
-Chordal Cyc. 5,000 18.80 31.34 1 1 3/2/5
-Triangle Rect. 5,000 18.80 31.34 1 1 3/2/5
-4.1.4. Geometric Graphs
-The counting task aims to predict the number of specific substructures in each graph.
-We collect 14 datasets (Zhao et al., 2022b).
-Labels include the counts of various substructures, including 3-Cycle, 4-Cycle, 5-Cycle, 6-Cycle, 7-Cycle, 8-Cycle, 4-Path, 5-Path, 6-Path, 4-Clique, Tailed-Triangle (Tailed Tri.), Chordal-Cycle, and Triangle-Rectangle.
-4.2. Baselines
-We compare and comprehensively investigate our proposed model with 14 representative and effective baselines in five types.
+MUTAG, NCI1, MolBACE, and MolPCBA are molecular graph datasets, where nodes represent atoms and edges represent chemical bonds.
+  • MUTAG (MUT) (Morris et al., 2020) and NCI1 (NCI) (Morris et al., 2020).
+In both datasets, node features are one-hot encodings of atomic species (e.g., C, N, O).
+They are binary classification tasks: MUTAG labels indicate whether a molecule is mutagenic, while NCI1 labels indicate anti-cancer activity in cell-based assays.
+  • MolBACE (BAC) (Hu et al., 2020a) and MolPCBA (PCB) (Hu et al., 2020a).
+In both datasets, node features are derived from RDKit, including properties such as chirality and aromaticity.
+MolBACE is a single-label binary task for predicting whether a molecule inhibits the BACE-1 enzyme (linked to Alzheimer’s disease).
+MolPCBA is a 128-label binary task, each representing activity in a specific biological assay.
+4.2.4. Motif-Counting Graphs
+The graphs and their splits are sourced from (Chen et al., 2020) and have been widely used in follow-up works (Chen et al., 2020; Zhao et al., 2022b; Huang et al., 2022; Yan et al., 2024).
+Each graph is labeled with the counts of specific substructures, including Cycles (3-8 nodes), Paths (4-6 nodes), 4-Cliques, Tailed Triangles, Chordal Cycles, and Triangle-Rectangles.
+Graphs are generated by sampling a $d$-regular graph with $m$ nodes and randomly deleting $m$ edges, where $(m,d)$ is uniformly chosen from $\{(10,6),(15,6),(20,5),(30,5)\}$.
+The ground-truth counts are computed using the subgraph counting algorithm proposed in (Shervashidze et al., 2009).
+Table 3.
+Data statistics.
+$\overline{\mbox{Nodes}}$ and $\overline{\mbox{Edges}}$ denote the average number of nodes and edges per graph, respectively.
+Split is the data partition strategy, where 10-fold denotes 10-fold cross-validation, others denote train/validation/test splits.
+Type Dataset Graphs $\overline{\textbf{Nodes}}$ $\overline{\textbf{Edges}}$ Classes Split
+SN IMDB-B (I-B) 1,000 19.8 96.5 2 10-fold
+IMDB-M (I-M) 1,500 13.0 65.9 3 10-fold
+REDDIT-B (RED) 2,000 429.6 497.8 2 10-fold
+COLLAB (COL) 5,000 74.5 2457.8 3 10-fold
+BIO PROTEINS (PRO) 1,113 39.1 72.8 2 10-fold
+$D$ 1,178 284.3 715.7 2 10-fold
+ENZYMES (ENZ) 600 32.6 62.1 6 10-fold
+MolHIV (HIV) 41,127 25.5 54.9 2 8/1/1
+MolTox21 (TOX) 7,831 18.9 39.2 12 8/1/1
+CHE MUTAG (MUT) 188 17.9 19.8 2 10-fold
+NCI1 (NCI) 4,110 29.9 32.3 2 10-fold
+MolBACE (BAC) 1,513 34.1 36.9 2 8/1/1
+MolPCBA (PCB) 437,929 26.0 28.1 128 8/1/1
+MC {3,4,5,6,7,8}-Cycle 5,000 18.8 31.3 1 3/2/5
+{4,5,6}-Path 5,000 18.8 31.3 1 3/2/5
+4-Clique 5,000 18.8 31.3 1 3/2/5
+Tailed Tri. 5,000 18.8 31.3 1 3/2/5
+Chor. Cyc. 5,000 18.8 31.3 1 3/2/5
+Tri. Rec. 5,000 18.8 31.3 1 3/2/5
+4.3. Evaluated GNNs
+We comprehensively evaluate 16 representative and effective GNNs across five categories as follows.
 4.3.1. Node-based GNNs
-  • GCN (Kipf and Welling, 2016). It utilizes graph convolutional layers to aggregate information from neighbors, serving as a fundamental GNN.
-  • GIN (Xu et al., 2019). The GIN baseline employs sum pooling and utilizes Graph Isomorphism Network as the message passing layer.
+  • GCN (Kipf and Welling, 2016).
+As a foundational GNN, GCN learns node features by aggregating information from their neighborhood, effectively performing a weighted average of neighbor representations in each layer.
+  • GIN (Xu et al., 2019).
+GIN utilizes a Multi-Layer Perceptron (MLP) to update node representations based on a sum of its neighbors, theoretically matching the capability of the Weisfeiler-Lehman (1-WL) graph isomorphism test.
   • GraphSAGE (SAGE) (Hamilton et al., 2017).
-It aggregates information from sampled neighbors using different functions, such as mean, LSTM, or pooling, to update node embeddings in a scalable manner.
+SAGE introduces a sampling-based neighborhood aggregation strategy.
+For each node, it uniformly samples a fixed number of neighbors to generate embeddings, making the learning process scalable and inductive.
+  • PNA (Corso et al., 2020).
+PNA combines multiple aggregators (e.g., mean, max, sum) with degree-based scalers to capture the distribution of neighbor features, allowing the model to distinguish diverse neighbor patterns better than single-aggregator methods.
 4.3.2. Pooling-based GNNs
-  • TopK (Cangea et al., 2018; Gao and Ji, 2019). It performs node drop by selecting the top-ranked nodes based on a learnable scoring function.
-  • GMT (Baek et al., 2021). It condenses nodes into important ones based on grouping-matrix and create edges via attention mechanism.
-  • EPool (Diehl, 2019). It progressively merges nodes by contracting high-scoring edges between adjacent nodes through edge clustering.
+  • TopK (Cangea et al., 2018; Gao and Ji, 2019).
+TopK uses a trainable projection vector to score nodes, with the top-k nodes passing GNN backbones selected to form the graph for the subsequent layer.
+  • GMT (Baek et al., 2021).
+GMT introduces a hierarchical pooling method where nodes are adaptively grouped into clusters using a transformer-based attention mechanism, allowing the model to learn soft cluster assignments and generate coarse-grained graph representations at multiple scales.
+  • EdgePool (EPool) (Diehl, 2019).
+EPool learns a hierarchical representation by iteratively contracting edges.
+It scores edges based on their importance for connecting nodes and merges the corresponding node pairs, effectively simplifying the graph topology layer by layer.
 4.3.3. Subgraph-based GNNs
-  • ECS (Yan et al., 2024). It integrates a base GNN with precomputed structural embeddings to efficiently count substructures using distance information and message passing.
-  • AK+ (Zhao et al., 2022c). It extends base GNNs by replacing the local aggregation mechanism with a GNN that encodes general subgraphs instead of star-shaped patterns.
-  • I2GNN (Huang et al., 2022). It enhances Subgraph MPNNs by employing multiple node identifiers, assigning distinct identifiers to the root node and its neighbors within each subgraph.
+  • ECS (Yan et al., 2024).
+ECS boosts the GNN backones by injecting pre-calculated structural information.
+It computes embeddings that encode subgraph-centric distances, allowing the model to count substructures with provable guarantees.
+  • GNNAK+ (AK+) (Zhao et al., 2022a).
+AK+ shifts the computational unit from single-node neighborhoods to k-hop subgraphs, thereby enabling the GNN backbones to capture more complex, high-order structural patterns.
+  • I2GNN (I2) (Huang et al., 2022).
+I2GNN breaks structural symmetries by assigning unique identifiers to the central root node and its neighbors within each sampled subgraph, enabling finer-grained representation learning.
 4.3.4. GL-based GNNs
-  • VIBGSL (VIB) (Sun et al., 2022). It uses the Information Bottleneck principle to mask irrelevant features and learn a task-relevant graph structure for improved GNN performance.
-  • HGPSL (HGP) (Zhang et al., 2019). It selects key nodes to form an induced subgraph and refines the graph structure using sparse attention-based structure learning.
-  • MOSGSL (MO) (Zhou et al., 2024). It reconstructs key subgraphs using a motif-driven guidance module that captures and aligns discriminative structural patterns
+  • VIBGSL (VIB) (Sun et al., 2022). VIB uses the Information Bottleneck principle to mask irrelevant features and learn a task-relevant graph structure for better modeling the key patterns.
+  • HGP-SL (HGP) (Zhang et al., 2019). HGP selects key nodes to form an induced subgraph and refines the graph structure using sparse attention-based structure learning.
+  • MOSGSL (MO) (Zhou et al., 2024). MO reconstructs key subgraphs using a motif-driven guidance module that dynamically captures and aligns discriminative structural patterns.
 4.3.5. SSL-based GNNs
-  • RGC (Sun et al., 2024). It employs a diverse-curvature GCN and motif-aware contrastive learning on a stable kernel layer to model motif regularity and learn node representations.
-  • MVGRL(MVG) (Hassani and Khasahmadi, 2020). It uses diffusion algorithms, such as Personalized PageRank, to generate two positive views of the graph and maximize their similarity.
-  • Ours.
-It enhances expressivity with a $k$-path rooted subgraph approach and improves generalization through adaptive graph contrastive learning, ensuring robust performance across diverse tasks and scenarios.
-4.3. Hyperparameter and Hardware Setting
-For classification datasets, we set the batch size as 32 for four larger datasets (REDDIT-B, COLLAB, DD, and MolPCBA) and 128 for the other datasets.
-For regression datasets, we set the batch size as 256 The baselines and our model use different hyperparameters, which are fine-tuned by grid search based on validation data.
-For example, the hidden dimension is selected from $\{64,128,256,512\}$, the learning rate from $\{1e-2,1e-3,1e-4,1e-5\}$, and the dropout rate from $\{0,0.1,0.2,0.3,0.4,0.5\}$.
-Our model sets $k=1$ for graph classification, and sets $k=3$ for regression with 3-path sampling uniformly.
-All models are trained for a maximum of 2000 epochs, with early stopping applied if no improvement is observed on the validation set within 50 epochs.
-All codes are executed on a CentOS 7 machine equipped with a 20-core Intel® Xeon® Silver 4210 CPU @ 2.20GHz, 8 NVIDIA GeForce RTX 2080 Ti GPUs, and 256GB of RAM.
+  • RGC (Sun et al., 2024). RGC employs a diverse-curvature GCN and motif-aware contrastive learning on a stable kernel layer to robustly model motif regularity and learn node representations.
+  • MVGRL (MVG) (Hassani and Khasahmadi, 2020). MVG learns graph representations by contrasting node embeddings from the original and diffusion-based views of the graph.
+  • GCA (Zhu and others, 2021). GCA learns graph representations by contrasting node embeddings across adaptively augmented views that preserve important structural and semantic information.
 4.4. Evaluation Metric
-We define the effectiveness and efficiency metrics as follows.
+We evaluate the performance of GNNs using effectiveness and efficiency metrics as follows.
 4.4.1. Effectiveness Metric
@@
   • Strict Accuracy (Acc).
+Strict accuracy is used to measure the proportion of exact matches between predicted and true labels.
 Strict accuracy is defined as $Acc=\frac{1}{|\mathcal{LG}|}\sum_{G_{i}\in\mathcal{LG}}{\mathbb{I}(y_{i}=\hat{y}_{i})}$, where $\mathbb{I}(y_{i}=\hat{y}_{i})=1$ if only $y_{i}=\hat{y}_{i}$.
   • Micro-F1 (Mi-F1).
+Micro-F1 is a performance metric that considers the overall precision and recall across all instances in the dataset.
 The Micro-precision is defined as $Mi\text{-}P=\frac{\sum_{G_{i}\in\mathcal{LG}}{|y_{i}\cap\hat{y}_{i}|}}{\sum_{G_{j}\in\mathcal{LG}}{|\hat{y}_{j}|}}$ and Micro-recall is $Mi\text{-}R=\frac{\sum_{G_{i}\in\mathcal{LG}}{|y_{i}\cap\hat{y}_{i}|}}{\sum_{G_{j}\in\mathcal{LG}}{|{y}_{j}|}}$.
@@
   • Macro-F1 (Ma-F1).
-The Macro-precision is defined as $Ma\text{-}P=\frac{1}{|\mathcal{LG}|}\sum_{G_{i}\in\mathcal{LG}}\frac{|y_{i}\cap\hat{y}_{i}|}{|\hat{y}_{i}|}$ and Macro-recall is defined as $Ma\text{-}R=\frac{1}{|\mathcal{LG}|}\sum_{G_{i}\in\mathcal{LG}}\frac{|y_{i}\cap\hat{y}_{i}|}{|y_{i}|}$, Macro-F1 $Ma\text{-}F1=\frac{2\times Ma\text{-}P\times Ma\text{-}R}{Ma\text{-}P+Ma\text{-}R}$.
+Macro-F1 evaluates the average performance of precision and recall across all instances, treating each equally regardless of size.
+The Macro-precision is defined as $Ma\text{-}P=\frac{1}{|\mathcal{LG}|}\sum_{G_{i}\in\mathcal{LG}}\frac{|y_{i}\cap\hat{y}_{i}|}{|\hat{y}_{i}|}$ and Macro-recall is defined as $Ma\text{-}R=\frac{1}{|\mathcal{LG}|}\sum_{G_{i}\in\mathcal{LG}}\frac{|y_{i}\cap\hat{y}_{i}|}{|y_{i}|}$, so the Macro-F1 is defined as $Ma\text{-}F1=\frac{2\times Ma\text{-}P\times Ma\text{-}R}{Ma\text{-}P+Ma\text{-}R}$.
 For graph regression tasks, we use the Mean Absolute Error (MAE) and R2 as follows.
-  • Mean Absolute Error (MAE). Mean Absolute Error is defined as $MAE=\frac{1}{|\mathcal{LG}|}\sum_{G_{i}\in\mathcal{LG}}{|y_{i}-\hat{y}_{i}|}$.
+  • Mean Absolute Error (MAE).
+Mean Absolute Error measures the average magnitude of errors between predicted and true values.
+MAE is defined as $MAE=\frac{1}{|\mathcal{LG}|}\sum_{G_{i}\in\mathcal{LG}}{|y_{i}-\hat{y}_{i}|}$.
+Lower $MAE$ indicates better performance.
   • R2.
+R2 evaluates the proportion of variance in the true values that is captured by the predicted values.
 The R2 is defined as $R2=1-\frac{\sum_{G_{i}\in\mathcal{LG}}(y_{i}-\hat{y}_{i})^{2}}{\sum_{G_{i}\in\mathcal{LG}}(y_{i}-\bar{y})^{2}}\in[0,1]$, where $\bar{y}=\frac{1}{|\mathcal{LG}|}\sum_{G_{i}\in\mathcal{LG}}y_{i}$ is the mean of the true values.
-Higher $R2$ indicates better prediction performance.
+Higher $R2$ indicates better performance.
 4.4.2. Efficiency Metric
-We evaluate the efficiency of models on both graph classification and regression tasks based on the training time (s), inference time (s), memory usage (MB) in training and inference phrases.
-Table 3. Evaluation on graph classification. The best and the second performance are highlighted in bold and in underline.
-Type Dataset Metric Node-based Pooling-based Subgraph-based GL-based SSL-based
-GCN GIN SAGE TopK GMT EPool ECS AK+ I2GNN VIB HGP MO RGC MVG Ours
-SN IMDB-B Acc 0.6840 0.7100 0.6860 0.6860 0.7440 0.6980 0.7150 0.7290 0.7070 0.7250 0.7030 0.7310 0.6230 0.6970 0.7550
-F1 0.6963 0.7117 0.6991 0.7105 0.7625 0.7146 0.7158 0.7391 0.7014 0.7185 0.7078 0.7422 0.6321 0.6940 0.7693
-IMDB-M Acc 0.4620 0.4820 0.4650 0.4670 0.5080 0.4750 0.4727 0.4967 OOM 0.4720 0.4650 0.5067 0.4130 0.4867 0.5233
-F1 0.4393 0.4706 0.4512 0.4435 0.4850 0.4563 0.4466 0.4839 OOM 0.4443 0.4376 0.4874 0.3867 0.4788 0.5087
-REDDIT-B Acc 0.9305 0.8965 0.9094 0.9280 0.9195 0.9260 OOM OOM OOM 0.8276 OOM 0.8625 OOM OOM 0.9325
-F1 0.9322 0.9026 0.9132 0.9284 0.9263 0.9303 OOM OOM OOM 0.8298 OOM 0.8722 OOM OOM 0.9330
-COLLAB Acc 0.7644 0.7328 0.7406 0.7556 0.8164 0.7696 OOM OOM OOM 0.7528 0.6988 0.8278 OOM 0.7688 0.7862
-F1 0.7454 0.7099 0.7169 0.7395 0.7973 0.7520 OOM OOM OOM 0.7114 0.6702 0.8046 OOM 0.7320 0.7432
-BIO PROTEINS Acc 0.7286 0.7250 0.7367 0.7277 0.7440 0.7286 0.7061 0.7475 0.7071 0.7322 0.7304 0.7232 0.7015 0.7367 0.7457
-F1 0.6596 0.6432 0.6617 0.6712 0.7058 0.6598 0.5838 0.6978 0.6026 0.6641 0.6653 0.6442 0.6386 0.6993 0.6998
-DD Acc 0.7310 0.7233 0.7542 0.7156 0.7802 0.7343 OOM 0.7776 0.7335 0.7632 0.7598 0.7632 OOM OOM 0.7632
-F1 0.6676 0.6555 0.6707 0.6384 0.7252 0.6687 OOM 0.7267 0.6402 0.6845 0.6841 0.6844 OOM OOM 0.6886
-ENZYMES Acc 0.4650 0.4833 0.5183 0.4800 0.4950 0.4783 0.4617 0.5267 0.4650 0.4417 0.4667 0.5250 0.4917 0.5333 0.5533
-F1 0.4775 0.4781 0.5098 0.4794 0.4237 0.4737 0.4592 0.5279 0.4628 0.4337 0.4695 0.5309 0.4902 0.5314 0.5671
-MolHIV Acc 0.9696 0.9701 0.9695 0.9687 0.9691 0.9696 0.9689 0.9728 0.9703 0.9686 0.9689 0.9701 0.9686 0.9705 0.9742
-F1 0.3192 0.3466 0.2991 0.2221 0.2861 0.3213 0.2537 0.3682 0.3413 0.2203 0.2518 0.2502 0.2225 0.3628 0.3693
-MolTOX21 Acc 0.5548 0.5553 0.5529 0.5523 0.5412 0.5514 0.5463 0.5561 0.5580 0.5283 0.5350 0.5384 0.5112 0.5565 0.5727
-Mi-F1 0.9120 0.9114 0.9096 0.9083 0.9109 0.9101 0.9121 0.9138 0.9137 0.8993 0.9004 0.9048 0.8986 0.9126 0.9174
-Ma-F1 0.3628 0.3601 0.3418 0.2240 0.3452 0.3765 0.3858 0.3850 0.3815 0.2020 0.2125 0.2180 0.1981 0.3847 0.3882
-CHE MUTAG Acc 0.8041 0.8570 0.8199 0.8094 0.8254 0.8091 0.8091 0.8409 0.8062 0.7643 0.7706 0.7790 0.7067 0.8607 0.9102
-F1 0.8583 0.8907 0.8644 0.8590 0.8642 0.8506 0.8558 0.8770 0.8512 0.8255 0.8296 0.8447 0.8012 0.8912 0.9310
-NCI1 Acc 0.8158 0.8154 0.8146 0.8169 0.7662 0.8160 0.7866 0.8187 0.7603 0.7816 0.7825 0.7852 0.6995 0.7560 0.8253
-F1 0.8160 0.8153 0.8149 0.8171 0.7708 0.8168 0.7914 0.8190 0.7673 0.7827 0.7831 0.7855 0.6997 0.7557 0.8249
-MoIBACE Acc 0.6711 0.6645 0.6477 0.6756 0.6513 0.6749 0.7125 0.6803 0.6078 0.6019 0.6689 0.6294 0.6089 0.7099 0.7895
-F1 0.7154 0.7351 0.7099 0.7264 0.7366 0.7436 0.7400 0.7382 0.7055 0.7001 0.6545 0.7077 0.7047 0.7366 0.8242
-MolPCBA Acc 0.5456 0.5466 0.5451 0.5461 0.5447 ¿3 days OOM 0.5523 0.5414 OOM OOM OOM OOM OOM 0.5531
-Mi-F1 0.9850 0.9853 0.6945 0.9849 0.9836 ¿3 days OOM 0.9860 0.9850 OOM OOM OOM OOM OOM 0.9858
-Ma-F1 0.1200 0.1500 0.1317 0.1333 0.0167 ¿3 days OOM 0.2140 0.2064 OOM OOM OOM OOM OOM 0.2152
-Table 4. Evaluation on graph regression. The best and the second performance are highlighted in bold and in underline.
-Type Dataset Metric Node-based Pooling-based Subgraph-based GL-based SSL-based
-GCN GIN SAGE TopK GMT EPool ECS AK+ I2GNN VIB HGP MO RGC MVG Ours
-GE 3-Cycle MAE 0.4396 0.3962 0.5119 0.4290 0.4251 0.4238 0.0192 0.0022 0.0009 0.8776 0.4369 0.5410 0.4744 0.4231 0.0006
-R2 0.6970 0.7546 0.8081 0.7044 0.7107 0.5679 0.9996 0.9998 0.9999 0.0096 0.6850 0.5003 0.6062 0.7174 0.9999
-4-Cycle MAE 0.2806 0.2536 0.5407 0.2772 0.2720 0.2696 0.0148 0.0223 0.0062 0.7851 0.2745 0.5442 0.5398 0.2734 0.0039
-R2 0.8229 0.8864 0.4011 0.8331 0.8477 0.8406 0.9997 0.9990 0.9998 0.0157 0.8350 0.4032 0.4436 0.8606 0.9999
-5-Cycle MAE 0.2782 0.1858 0.4608 0.2399 0.2338 0.2659 0.0716 0.0340 0.0115 0.9021 0.2763 0.4532 0.4570 0.2180 0.0078
-R2 0.8334 0.9364 0.4815 0.8614 0.8936 0.8553 0.9862 0.9972 0.9996 0.0019 0.8190 0.6081 0.4939 0.9060 0.9998
-6-Cycle MAE 0.3014 0.1895 0.4690 0.2302 0.1788 0.2926 0.0862 0.0584 0.0368 0.8876 0.3035 0.4557 0.5027 0.1773 0.0275
-R2 0.7926 0.9162 0.6189 0.8803 0.9279 0.8073 0.9569 0.9956 0.9969 0.0021 0.7891 0.6219 0.6077 0.9301 0.9982
-7-Cycle MAE 0.4011 0.2111 0.5898 0.2852 0.1569 0.3942 0.1560 0.0563 0.0488 0.8273 0.4220 0.5713 0.5836 0.1500 0.0414
-R2 0.5887 0.8635 0.4594 0.7921 0.9354 0.6091 0.9460 0.9937 0.9946 0.0064 0.5487 0.4883 0.4692 0.9404 0.9957
-8-Cycle MAE 0.4757 0.2632 0.5285 0.2910 0.1377 0.4701 0.1153 0.0493 0.0399 0.7430 0.4790 0.5340 0.4808 0.1286 0.0337
-R2 0.3570 0.7220 0.2212 0.7553 0.9428 0.3840 0.9703 0.9949 0.9963 0.0330 0.3520 0.2051 0.3319 0.9466 0.9974
-4-Path MAE 0.7154 0.4273 0.7340 0.5273 0.1606 0.6361 0.0243 0.0150 0.0083 0.7784 0.5019 0.7316 0.7270 0.1512 0.0066
-R2 0.1574 0.6349 0.0898 0.5270 0.9459 0.3188 0.9921 0.9996 0.9998 0.0279 0.5390 0.0909 0.1310 0.9534 0.9999
-5-Path MAE 0.6848 0.3949 0.6669 0.5189 0.1559 0.6133 0.0134 0.0153 0.0092 0.7510 0.4846 0.6933 0.7053 0.1407 0.0081
-R2 0.1050 0.6358 0.1675 0.5015 0.9473 0.3254 0.9995 0.9995 0.9998 0.0283 0.5204 0.0823 0.0650 0.9562 0.9999
-6-Path MAE 0.6164 0.3913 0.6594 0.5629 0.1394 0.6930 0.0138 0.0133 0.0089 0.7571 0.4552 0.6627 0.6589 0.1299 0.0079
-R2 0.2470 0.5958 0.1286 0.4133 0.9539 0.1408 0.9995 0.9996 0.9998 0.0250 0.5459 0.1282 0.1280 0.9607 0.9998
-4-Clique MAE 0.3432 0.3449 0.3498 0.3430 0.3427 0.3578 0.0091 0.0092 0.0005 0.3866 0.2355 0.3422 0.3424 0.1797 0.0004
-R2 0.1612 0.1348 0.1086 0.1213 0.1022 0.1057 0.9662 0.9964 0.9999 0.0890 0.8896 0.1604 0.1610 0.9002 0.9999
-Tailed Triangle MAE 0.3506 0.2886 0.3812 0.3465 0.3665 0.3405 0.0189 0.0146 0.0034 0.8934 0.3791 0.3891 0.4095 0.3277 0.0018
-R2 0.7884 0.8606 0.7303 0.7813 0.7673 0.8053 0.9989 0.9995 0.9999 0.0176 0.7336 0.7008 0.6245 0.8177 0.9999
-Chordal Cycle MAE 0.4377 0.3683 0.4390 0.4216 0.3690 0.4246 0.0295 0.0336 0.0044 0.8720 0.4441 0.4413 0.4728 0.3671 0.0028
-R2 0.6346 0.7109 0.6002 0.6414 0.6848 0.6421 0.9970 0.9951 0.9999 0.0333 0.5746 0.5841 0.5684 0.7293 0.9999
-Triangle Rectangle MAE 0.4657 0.4034 0.4794 0.4736 0.4068 0.4890 0.3619 0.3464 0.3271 0.8263 0.4286 0.4755 0.5409 0.4301 0.3123
-R2 0.5211 0.6291 0.5335 0.4921 0.6235 0.4702 0.6895 0.7063 0.7241 0.0338 0.5828 0.5037 0.3852 0.5691 0.7501
+We evaluate the efficiency of models on both graph classification and regression tasks based on the training time (s), inference time (s), memory usage (MB) in training and inference phases.
+4.5. Hyperparameter and Hardware Setting
+For classification datasets, we set the batch size as 32 for four larger datasets (REDDIT, COLLAB, DD, and MolPCBA) and 128 for the other datasets.
+For regression datasets, we set the batch size as 256.
+To efficiently tune the hyperparameters for each model, we employed the Optuna framework (Akiba et al., 2019).
+For each model, we conducted $200$ trials using the Tree-structured Parzen Estimator (TPE) sampler and a MedianPruner to terminate unpromising trials early.
+The hyperparameter search spaces were defined as follows: the hidden dimension is selected from $\{64,128,256,512\}$, the learning rate from $\{1e-2,1e-3,1e-4,1e-5\}$, GNN layers from $\{1,2,3,4\}$, and the dropout rate from $\{0,0.1,0.2,0.3,0.4,0.5\}$.
+All models are trained for a maximum of 2000 epochs, with early stopping applied if no improvement is observed on the validation set within 50 epochs.
+All experiments are executed on a CentOS 7 machine equipped with dual 10-core Intel® Xeon® Silver 4210 CPUs @ 2.20GHz, 8 NVIDIA GeForce RTX 2080 Ti GPUs (11GB each) and 256GB RAM.
 4.6. Effectiveness Evaluation
-In this subsection, we evaluate the effectiveness of GNNs on graph tasks from two aspects.
-Firstly, in Section 4.5.1, we assess whether existing GNNs can capture the core semantic patterns of graphs, where the core semantic patterns mainly determine graph labels.
-Secondly, in Section 4.5.2, we evaluate the expressivity ability of GNNs based, i.e., whether GNNs can distinguish graph geometric patterns, such as the number of cycles and paths.
-Table 5. Efficiency evaluation. We report the total training time (Train (s)), and total inference time (Infer (s)), and GPU memory during training (T-Mem) and inference (I-Mem).
-Type Dataset Metric Node-based Pooling-based Subgraph-based GL-based SSL-based
-GCN GIN SAGE TopK GMT EPool ECS AK+ I2GNN VIB HGP MO RGC MVG Ours
-SN REDDIT-B Train(s) 192.44 182.98 179.55 218.03 758.22 8365.60 OOM OOM OOM OOM OOM 371.86 OOM OOM 411.93
-Infer(s) 28.42 26.21 26.04 32.95 30.45 1953.91 OOM OOM OOM OOM OOM 79.35 OOM OOM 34.50
-T-Mem(MB) 287 288 310 287 2120 360 OOM OOM OOM OOM OOM 604 OOM OOM 4287
-I-Mem(MB) 190 122 135 190 244 194 OOM OOM OOM OOM OOM 289 OOM OOM 190
-BIO ENZYMES Train(s) 26.05 26.34 25.23 31.95 38.85 375.25 312.46 350.25 470.05 120.48 67.5 75.29 495.23 57.53 55.45
-Infer(s) 5.19 5.33 5.04 6.84 7.17 76.30 68.74 68.93 94.92 36.12 19.20 28.27 85.11 7.76 5.80
-T-Mem(MB) 83 78 77 83 228 103 582 2369 1656 323 186 506 282 324 246
-I-Mem(MB) 42 36 36 42 46 45 270 879 790 152 95 194 54 58 42
-CHE MUTAG Train(s) 10.78 11.03 10.05 13.50 15.96 60.13 52.05 88.70 67.33 44.37 35.32 39.99 80.49 25.15 18.72
-Infer(s) 3.86 3.89 3.51 5.27 4.01 17.74 18.91 21.18 25.10 15.80 11.63 15.18 29.78 7.64 4.22
-T-Mem(MB) 42 44 44 42 82 46 422 775 1306 56 48 431 134 140 102
-I-Mem(MB) 19 20 21 19 24 19 245 210 722 26 27 74 24 25 19
-GE 4-Cycle Train(s) 166.50 160.95 140.88 180.57 1224.39 482.95 1180.51 645.97 1487.66 395.71 284.34 394.60 1770.94 300.21 1503.16
-Infer(s) 256.17 241.02 235.40 279.43 2058.62 1033.68 1776.91 920.13 2219.63 549.60 441.95 583.55 2675.22 461.87 2255.94
-T-Mem(MB) 88 85 85 88 380 104 1214 1734 3080 262 269 470 238 237 4327
-I-Mem(MB) 64 62 62 64 260 68 342 608 726 104 152 270 103 96 737
+Table 4.
+Evaluation on graph classification.
+The best and second performance are highlighted in bold and underlined, respectively.
+“Met.” means Metrics, “OOM” means out-of-memory, and “>3>3days” means training cannot be completed within 3 days.
+Type Data Met. Node-based Pooling-based Subgraph-based GL-based SSL-based
+GCN GIN SAGE PNA TopK GMT EPool ECS AK+ I2 VIB HGP MO RGC MVG GCA
+SN I-B Acc 0.6840 0.7100 0.6860 0.7280 0.6860 0.7440 0.6980 0.7150 0.7290 0.7070 0.7250 0.7030 0.7310 0.6230 0.6970 0.7120
+F1 0.6963 0.7117 0.6991 0.7391 0.7105 0.7625 0.7146 0.7158 0.7391 0.7014 0.7185 0.7078 0.7422 0.6321 0.6940 0.7147
+I-M Acc 0.4620 0.4820 0.4650 0.4980 0.4670 0.5080 0.4750 0.4727 0.4967 OOM 0.4720 0.4650 0.5067 0.4130 0.4867 0.4833
+F1 0.4393 0.4706 0.4512 0.4842 0.4435 0.4850 0.4563 0.4466 0.4839 OOM 0.4443 0.4376 0.4874 0.3867 0.4788 0.4726
+RED Acc 0.9029 0.8965 0.9094 OOM 0.9280 0.9195 0.9260 OOM OOM OOM 0.8276 OOM 0.8625 OOM OOM OOM
+F1 0.9123 0.9026 0.9132 OOM 0.9322 0.9263 0.9303 OOM OOM OOM 0.8298 OOM 0.8722 OOM OOM OOM
+COL Acc 0.7644 0.7328 0.7406 OOM 0.7556 0.8164 0.7696 OOM OOM OOM 0.7528 0.6988 0.8278 OOM 0.7688 OOM
+F1 0.7454 0.7099 0.7169 OOM 0.7395 0.7973 0.7520 OOM OOM OOM 0.7114 0.6702 0.8046 OOM 0.7320 OOM
+BIO PRO Acc 0.7286 0.7250 0.7367 0.7372 0.7277 0.7440 0.7286 0.7061 0.7475 0.7071 0.7322 0.7304 0.7232 0.7015 0.7367 0.7322
+F1 0.6596 0.6432 0.6617 0.6645 0.6712 0.7058 0.6598 0.5838 0.6978 0.6026 0.6641 0.6653 0.6442 0.6386 0.6993 0.6647
+$D$ Acc 0.7310 0.7233 0.7542 0.7488 0.7156 0.7802 0.7343 OOM 0.7776 0.7335 0.7632 0.7598 0.7632 OOM OOM OOM
+F1 0.6676 0.6555 0.6707 0.7245 0.6384 0.7252 0.6687 OOM 0.7267 0.6402 0.6845 0.6841 0.6844 OOM OOM OOM
+ENZ Acc 0.4650 0.4833 0.5183 0.5233 0.4800 0.4950 0.4783 0.4617 0.5267 0.4650 0.4417 0.4667 0.5250 0.4917 0.5317 0.4817
+F1 0.4775 0.4781 0.5098 0.5204 0.4794 0.4237 0.4737 0.4592 0.5279 0.4628 0.4337 0.4695 0.5309 0.4902 0.5314 0.4772
+HIV Acc 0.9696 0.9701 0.9695 0.9703 0.9687 0.9691 0.9696 0.9689 0.9728 0.9703 0.9686 0.9689 0.9701 0.9686 0.9705 0.9701
+F1 0.3192 0.3466 0.2991 0.3622 0.2221 0.2861 0.3213 0.2537 0.3682 0.3413 0.2203 0.2518 0.2502 0.2225 0.3628 0.3448
+TOX Acc 0.5548 0.5553 0.5529 0.5561 0.5523 0.5412 0.5514 0.5463 0.5561 0.5580 0.5283 0.5350 0.5384 0.5112 0.5565 0.5558
+F1-i 0.9120 0.9114 0.9096 0.9135 0.9083 0.9109 0.9101 0.9121 0.9138 0.9137 0.8993 0.9004 0.9048 0.8986 0.9126 0.9123
+F1-a 0.3628 0.3601 0.3418 0.3839 0.2240 0.3452 0.3765 0.3858 0.3850 0.3815 0.2020 0.2125 0.2180 0.1981 0.3847 0.3695
+CHE MUT Acc 0.8041 0.8570 0.8199 0.8588 0.8094 0.8254 0.8091 0.8091 0.8409 0.8062 0.7643 0.7706 0.7790 0.7067 0.8607 0.8842
+F1 0.8583 0.8907 0.8644 0.8909 0.8590 0.8642 0.8506 0.8558 0.8770 0.8512 0.8255 0.8296 0.8447 0.8012 0.8912 0.9070
+NCI Acc 0.8158 0.8154 0.8146 0.8183 0.8169 0.7662 0.8160 0.7866 0.8187 0.7603 0.7816 0.7825 0.7852 0.6995 0.7560 0.8122
+F1 0.8160 0.8153 0.8149 0.8184 0.8171 0.7708 0.8168 0.7914 0.8190 0.7673 0.7827 0.7831 0.7855 0.6997 0.7557 0.8119
+BAC Acc 0.6711 0.6645 0.6477 0.7074 0.6756 0.6513 0.6749 0.7125 0.6803 0.6078 0.6019 0.6689 0.6294 0.6089 0.7099 0.6729
+F1 0.7154 0.7351 0.7099 0.7342 0.7264 0.7366 0.7436 0.7400 0.7382 0.7055 0.7001 0.6545 0.7077 0.7047 0.7366 0.7186
+PCB Acc 0.5456 0.5466 0.5451 0.5470 0.5461 0.5447 >3>3days OOM 0.5523 0.5414 OOM OOM OOM OOM OOM OOM
+Mi-F1 0.9850 0.9853 0.9839 0.9853 0.9849 0.9836 >3>3days OOM 0.9860 0.9850 OOM OOM OOM OOM OOM OOM
+Ma-F1 0.1200 0.1500 0.1317 0.1517 0.1333 0.0167 >3>3days OOM 0.2140 0.2064 OOM OOM OOM OOM OOM OOM
+Table 5. Evaluation on graph regression. The best and second performance are highlighted in bold and underlined, respectively.
+Type Data Met. Node-based Pooling-based Subgraph-based GL-based SSL-based
+GCN GIN SAGE PNA TopK GMT EPool ECS AK+ I2 VIB HGP MO RGC MVG GCA
+MC 3-Cyc MAE 0.4396 0.3962 0.5119 0.4060 0.4290 0.4251 0.4238 0.0192 0.0022 0.0009 0.8776 0.4369 0.5410 0.4744 0.4231 0.3873
+R2 0.6970 0.7546 0.8081 0.7493 0.7044 0.7107 0.5679 0.9996 0.9998 0.9999 0.1025 0.6850 0.5003 0.6062 0.7174 0.7595
+4-Cyc MAE 0.2806 0.2536 0.5407 0.2509 0.2772 0.2720 0.2696 0.0148 0.0223 0.0062 0.6451 0.2745 0.5442 0.5398 0.2734 0.2204
+R2 0.8229 0.8864 0.4011 0.8924 0.8331 0.8477 0.8406 0.9997 0.9990 0.9998 0.1424 0.8350 0.4032 0.4436 0.8606 0.8987
+5-Cyc MAE 0.2782 0.1858 0.4608 0.2580 0.2399 0.2338 0.2659 0.0716 0.0340 0.0115 0.9021 0.2763 0.4532 0.4570 0.2180 0.1756
+R2 0.8334 0.9364 0.4815 0.8866 0.8614 0.8936 0.8553 0.9862 0.9972 0.9996 0.0019 0.8190 0.6081 0.4939 0.9060 0.9408
+6-Cyc MAE 0.3014 0.1895 0.4690 0.2841 0.2302 0.1788 0.2926 0.0862 0.0584 0.0368 0.8876 0.3035 0.4557 0.5027 0.1773 0.1784
+R2 0.7926 0.9162 0.6189 0.8252 0.8803 0.9279 0.8073 0.9569 0.9956 0.9969 0.0021 0.7891 0.6219 0.6077 0.9301 0.9223
+7-Cyc MAE 0.4011 0.2111 0.5898 0.2237 0.2852 0.1569 0.3942 0.1560 0.0563 0.0488 0.8273 0.4220 0.5713 0.5836 0.1500 0.2045
+R2 0.5887 0.8635 0.4594 0.8469 0.7921 0.9354 0.6091 0.9460 0.9937 0.9946 0.0064 0.5487 0.4883 0.4692 0.9404 0.8770
+8-Cyc MAE 0.4757 0.2632 0.5285 0.2841 0.2910 0.1377 0.4701 0.1153 0.0493 0.0399 0.7430 0.4790 0.5340 0.4808 0.1286 0.2592
+R2 0.3570 0.7220 0.2212 0.7152 0.7553 0.9428 0.3840 0.9703 0.9949 0.9963 0.0330 0.3520 0.2051 0.3319 0.9466 0.7265
+4-Path MAE 0.7154 0.4273 0.7340 0.2940 0.5273 0.1606 0.6361 0.0243 0.0150 0.0083 0.7784 0.5019 0.7316 0.7270 0.1512 0.4093
+R2 0.1574 0.6349 0.0898 0.8619 0.5270 0.9459 0.3188 0.9921 0.9996 0.9998 0.0279 0.5390 0.0909 0.1310 0.9534 0.6579
+5-Path MAE 0.6848 0.3949 0.6669 0.3628 0.5189 0.1559 0.6133 0.0134 0.0153 0.0092 0.7510 0.4846 0.6933 0.7053 0.1407 0.3805
+R2 0.1050 0.6358 0.1675 0.6973 0.5015 0.9473 0.3254 0.9995 0.9995 0.9998 0.0283 0.5204 0.0823 0.0650 0.9562 0.6562
+6-Path MAE 0.6164 0.3913 0.6594 0.3397 0.5629 0.1394 0.6930 0.0138 0.0133 0.0089 0.7571 0.4552 0.6627 0.6589 0.1299 0.3784
+R2 0.2470 0.5958 0.1286 0.7429 0.4133 0.9539 0.1408 0.9995 0.9996 0.9998 0.0250 0.5459 0.1282 0.1280 0.9607 0.6076
+4-Cliq MAE 0.3432 0.3449 0.3498 0.2504 0.3430 0.3427 0.3578 0.0090 0.0092 0.0005 0.3866 0.2355 0.3422 0.3424 0.1797 0.3401
+R2 0.1612 0.1348 0.1086 0.8230 0.1213 0.1022 0.1057 0.9665 0.9964 0.9999 0.0890 0.8896 0.1604 0.1610 0.9002 0.1644
+Tailed Tri MAE 0.3506 0.2886 0.3812 0.3470 0.3465 0.3665 0.3405 0.0189 0.0146 0.0034 0.8934 0.3791 0.3891 0.4095 0.3277 0.2747
+R2 0.7884 0.8606 0.7303 0.7792 0.7813 0.7673 0.8053 0.9989 0.9995 0.9999 0.0176 0.7336 0.7008 0.6245 0.8177 0.8662
+Chor. Cyc MAE 0.4377 0.3683 0.4390 0.3500 0.4216 0.3690 0.4246 0.0295 0.0336 0.0044 0.8720 0.4441 0.4413 0.4728 0.3671 0.3608
+R2 0.6346 0.7109 0.6002 0.8121 0.6414 0.6848 0.6421 0.9970 0.9951 0.9999 0.0333 0.5746 0.5841 0.5684 0.7293 0.7149
+Tri. Rec. MAE 0.4657 0.4034 0.4794 0.4410 0.4736 0.4068 0.4890 0.3619 0.3464 0.3271 0.8263 0.4286 0.4755 0.5409 0.4301 0.3943
+R2 0.5211 0.6291 0.5335 0.5598 0.4921 0.6235 0.4702 0.6895 0.7063 0.7241 0.0338 0.5828 0.5037 0.3852 0.5691 0.6425
+Table 6.
+Efficiency evaluation.
+We report total training (Train (s)) and inference (Infer (s)) times with two decimal places, and GPU memory usage for training (T-Mem (MB)) and inference (I-Mem (MB)) in integers.
+“OOM” indicates out-of-memory.
+Type Data Met. Node-based Pooling-based Subgraph-based GL-based SSL-based
+GCN GIN SAGE PNA TopK GMT EPool ECS AK+ I2 VIB HGP MO RGC MVG GCA
+SN RED Train(s) 192.44 182.98 179.55 OOM 218.03 758.22 8365.60 OOM OOM OOM OOM OOM 371.86 OOM OOM OOM
+Infer(s) 28.42 26.21 26.04 OOM 32.95 30.45 1953.91 OOM OOM OOM OOM OOM 79.35 OOM OOM OOM
+T-Mem(MB) 287 288 310 OOM 287 2120 360 OOM OOM OOM OOM OOM 604 OOM OOM OOM
+I-Mem(MB) 190 122 135 OOM 190 244 194 OOM OOM OOM OOM OOM 289 OOM OOM OOM
+BIO ENZ Train(s) 26.05 26.34 25.23 72.33 31.95 38.85 375.25 312.46 350.25 470.05 120.48 67.50 75.29 495.23 57.53 125.45
+Infer(s) 5.19 5.33 5.04 17.42 6.84 7.17 76.30 68.74 68.93 94.92 36.12 19.20 28.27 85.11 7.76 5.07
+T-Mem(MB) 83 78 77 289 83 228 103 582 2369 1656 323 186 506 282 324 272
+I-Mem(MB) 42 36 36 96 42 46 45 270 879 790 152 95 194 54 58 42
+CHE MUT Train(s) 10.78 11.03 10.05 24.53 13.50 15.96 60.13 52.05 88.70 67.33 44.37 35.32 39.99 80.49 25.15 20.72
+Infer(s) 3.86 3.89 3.51 7.94 5.27 4.01 17.74 18.91 21.18 25.10 15.80 11.63 15.18 29.78 7.64 3.94
+T-Mem(MB) 42 44 44 152 42 82 46 422 775 1306 56 48 431 134 140 104
+I-Mem(MB) 19 20 21 82 19 24 19 245 210 722 26 27 74 24 25 19
+MC 4-Cyc Train(s) 166.50 160.95 140.88 458.70 180.57 1224.39 482.95 1180.51 645.97 1487.66 395.71 284.34 394.60 1770.94 300.21 304.95
+Infer(s) 256.17 241.02 235.40 590.20 279.43 2058.62 1033.68 1776.91 920.13 2219.63 549.60 441.95 583.55 2675.22 461.87 262.40
+T-Mem(MB) 88 85 85 228 88 380 104 1214 1734 3080 262 269 470 238 237 165
+I-Mem(MB) 64 62 62 209 64 260 68 342 608 726 104 152 270 103 96 64
 4.6.1. Graph Classification Tasks
-We evaluate baselines and our model on the graph classification tasks in Table 3.
-Firstly, node-based GNNs, such as GCN and GIN, cannot achieve satisfactory performance.
-These models learn node representations and use global pooling, which overlooks local graph structures, resulting in suboptimal outcomes compared to other GNN types.
-Secondly, pooling-based approaches propose to progressively reduce the graph size hierarchically and progressively learn the graph representations, which cannot achieve the best performance.
-This is because they simply keep the top-k most important nodes and concentrate similar nodes into a community to learn graph representation, ignoring local graph structure information.
-Thirdly, subgraph-based approaches effectively capture intricate local structures within the graphs, making them well-suited for bioinformatics applications where understanding detailed interactions is essential.
-However, their computational intensity and potential overfitting to subgraph patterns present notable challenges.
-Fourthly, graph learning-based approaches reconstruct graph structure and node features to remove noisy information, thereby enhancing performance on noisy datasets, such as social networks.
-Lastly, SSL-based approaches propose to learn generalizable representations based on pretraining on unlabeled graphs and fine-tuning on labeled graphs.
-Particularly, our proposed model achieves the best or comparable performance on all datasets.
-This is because we propose adaptive contrastive learning and sample subgraphs to learn local structure information, thereby learning a generalizable and informative graph representation.
+As shown in Table 4, node-based GNNs, such as GCN and SAGE, cannot achieve satisfactory performance.
+These models learn node representations and use global pooling.
+Global pooling overlooks local structural information, which is critical for distinguishing graphs in bioinformatics (e.g., ENZYMES) and chemistry (e.g., MUTAG).
+PNA achieves comparatively good performance as it captures richer neighborhood distributions by leveraging multiple aggregators and degree-scalers.
+Secondly, pooling-based approaches, such as GMT and TopK, achieve competitive performance, particularly on social networks like COLLAB and REDDIT.
+These methods progressively reduce the graph size by grouping nodes or retaining only the most important ones, thereby preserving multi-level structural information.
+However, they are less effective on datasets where fine-grained local structures (e.g., motifs) play a critical role, such as ENZYMES and NCI1.
+Subgraph-based methods, such as ECS, AK+, and I2, demonstrate highly competitive performance on bioinformatics and chemistry datasets because they break graphs into meaningful substructures, enabling them to capture important patterns that other methods often miss.
+However, they are out-of-memory (OOM) on large graphs or high node counts, such as REDDIT and COLLAB.
+Fourthly, graph learning-based approaches, such as MO and HGP, perform well on noisy social datasets, such as IMDB-B and IMDB-M.
+By dynamically reconstructing graph structures and removing irrelevant edges or nodes, these approaches enhance robustness and improve generalization.
+However, they are less effective on datasets where the original graph structure is already well-formed molecular graphs, such as MUTAG.
+Lastly, SSL-based methods like MVG and GCA perform robustly across multiple datasets, leveraging pretraining on unlabeled graphs to learn generalizable representations.
+However, SSL-based approaches are computationally expensive due to graph augmentation and can lead to OOM issues.
 4.6.2. Graph Regression Tasks
-For graph regression, we mainly evaluate the ability of GNNs to capture the semantics and core patterns of graphs, based on cycle counting and path counting for each graph.
-Lower MAE and higher R2 values indicate better performance.
-As shown in Table 4, node-based (except GIN), pooling-based, GL-based, and SSL-based approaches (except our model) do not achieve satisfactory performance.
-This is because these approaches are not designed to enhance the expressivity of GNNs, i.e., they cannot distinguish whether two graphs are isomorphic or have the same cycles.
-In contrast, GIN, subgraph-based GNNs, and our model aim to theoretically increase the expressivity of GNNs, thereby achieving satisfying performance.
-Particularly, our model performs the best on all datasets.
-This is because our model is guaranteed to have better expressivity, such as counting 8-Cycle, while other approaches have low expressivity.
-Additionally, we propose a unified contrastive learning approach to augment graphs and achieve more generalizable graph representations.
-4.6. Efficiency Evaluation
-We evaluate the efficiency of GNNs in terms of time and GPU peak memory consumption during both the training and inference phases.
-As shown in Table 5, node-based GNNs are the most efficient because they directly learn node representations from the original graphs without any additional processing.
-Secondly, the pooling-based approach, TopK also achieves good efficiency and low memory usage by retaining only the most important nodes However, GMT and EPool are slightly less efficient as it computes use nodes features to compute the cluster representation and pairwise similarities between connected nodes to form clusters, respectively.
-Thirdly, subgraph-based and GL-based approaches are less efficient due to their reliance on additional graph processing.
-Subgraph-based GNNs sample subgraphs for each root node or generate multiple subgraphs and learn representations for each one, which increases time and memory costs.
-Similarly, GL-based approaches iteratively optimize the graph structure and node features while training the GNNs, further adding to resource consumption.
-Lastly, SSL-based approaches are computationally intensive in training because they augment graphs into positive pairs and optimize GNNs by maximizing the similarity between these pairs, which increases time and memory usage.
-However, during inference, they do not require graph augmentation and operate directly on the original graphs, resulting in efficiency comparable to node-based GNNs.
+For graph regression tasks, the focus is primarily on evaluating how well GNNs can capture the semantics and key structural patterns of graphs, such as cycle and path counts for each graph.
+Lower MAE and higher $R^{2}$ scores reflect better performance.
+As shown in Table 5, node-based methods (excluding GIN and PNA), pooling-based models, GL-based techniques, and SSL-based approaches generally fail to deliver satisfactory results.
+This is because these methods are not specifically designed to enhance the expressiveness of GNNs, meaning they cannot effectively differentiate between isomorphic graphs or graphs with identical cycles.
+In contrast, GIN, PNA, and subgraph-based GNNs explicitly aim to improve the theoretical expressiveness of GNNs, resulting in superior performance.
+GIN and PNA aim to improve the theoretical expressiveness to approximate the Weisfeiler-Lehman (1-WL) isomorphism test.
+Subgraph-based models like ECS, AK+, and I2 consistently outperform other approaches on almost all regression datasets.
+By breaking graphs down into overlapping or rooted subgraphs, these methods capture rich structural details that enhance their theoretical expressivity.
+As a result, they are better able to distinguish between graph isomorphism classes and accurately identify important motifs, which leads to improved performance.
+Lastly, as the complexity of the target motif increases (e.g., from 3-Cycle to 8-Cycle), most GNNs (except subgraph-based ones) experience a drastic performance drop.
+This highlights the limitations of many GNN architectures in capturing higher-order dependencies and complex structural relationships.
 (a) IMDB-M
@@
 (d) MUTAG
-Figure 3. Robustness evaluation.
+Figure 3. Robustness evaluation on noisy graphs.
 (a) IMDB-M
@@
 Figure 5. Few-shot evaluation.
+4.7. Efficiency Evaluation
+We evaluate the efficiency of GNNs in terms of total training time, inference time, GPU peak memory usage during training and inference.
+As shown in Table 6, node-based GNNs exhibit the highest efficiency across all datasets.
+Their simplicity in aggregating neighbors’ information without additional graph processing ensures minimal time and memory usage.
+However, this efficiency comes at the cost of limited expressiveness as discussed in Section 4.6.
+Secondly, pooling-based methods strike a balance between efficiency and performance.
+TopK is efficient due to its node-pruning strategy, which reduces graph size while preserving key features.
+However, more advanced pooling approaches, such as GMT, are computationally expensive due to their clustering operations based on node similarity.
+Thirdly, subgraph-based GNNs (e.g., ECS, AK+, I2) are computationally intensive because they rely on generating multiple subgraphs for each graph.
+Despite their high resource requirements, these methods excel at capturing graph motifs and complex structural patterns, as shown in Section 4.6.
+Their computational cost makes them less efficient for real-time or large-scale applications.
+GL-based methods (e.g., VIB, HGP, and MO) dynamically reconstruct graph structures during training, which adds significant computational overhead.
+While they improve robustness to noise and graph quality, their iterative optimization process limits scalability to larger datasets.
+SSL-based methods (e.g., RGC, MVG, GCA) are computationally expensive during training because they require graph augmentations and additional contrastive learning stage.
+In summary, node-based GNNs are the most efficient but lack expressiveness, while pooling-based models strike a balance between efficiency and performance.
+Subgraph-based and GL-based approaches offer superior expressiveness but suffer from high computational costs.
+SSL-based methods, though computationally expensive during training, are efficient during inference, making them suitable for pretraining scenarios.
 4.8. More Scenarios
-We evaluate GNNs in more realistic scenarios, including noisy graphs, imbalanced graphs, and few-shot graphs.
+We further evaluate GNNs in three realistic and challenging scenarios, including noisy graphs, imbalanced graphs, and few-shot graphs.
 To ensure clarity, we select one representative graph dataset from each domain for evaluation, i.e., IMDB-M, ENZYMES, NCI1, and 4-Cycle.
-Additionally, we compare the most effective GNN model from each GNN category, including GCN, GMT, with AK+ and MO for classification tasks, I2GNN and HGPSL for regression tasks, with our model.
-4.7.1. Robustness Evaluation
-We evaluate the robustness of existing GNNs on noisy and incomplete graphs.
-Specifically, for each graph $G(V,\mathbf{A},\mathbf{X})$, we randomly delete $\alpha||\mathbf{A}||_{0}$ existing edges by varying a ratio $\alpha\in\{0.1,0.2,0.3,0.4,0.5\}$.
-As shown in Figure 3, as $\alpha$ increases, the performance of GNNs decreases and our proposed model outperform other baselines under the same ratio $\alpha$.
-Specifically, node-based GCN shows the sharply decline due to their reliance on the original graph structure and global pooling, which fail to capture local patterns effectively.
-Also, Pooling-based approach GMT also performs unsatisfying as it separates nodes into different clusters based on node similarity and connectivity, while edge deletions significantly affect the cluster results.
-Subgraph-based approach AK+ performs better by capturing intricate local patterns, but their reliance on subgraph sampling makes them sensitive to edge deletions.
-GL-based method MO shows resilience by dynamically learning graph structures.
-Particularly, our proposed model outperforms all baselines, maintaining strong performance on all $\alpha$ values.
-By leveraging adaptive contrastive learning and subgraph sampling, it effectively captures local structures and mitigates the impact of noise, resulting in superior robustness.
+Additionally, we compare some of the most representative models from each GNN category, including GCN (node-based), GMT (pooling-based), and GCA (SSL-based).
+For top perform, we select AK+ (subgraph-based), MO (GL-based) for classification tasks, while I2 (subgraph-based) and HGP (GL-based) for regression tasks.
+4.8.1. Robustness Evaluation on Noisy Graphs
+[…]
 4.8.2. Imbalance Data Evaluation
-In the graph classification task, given a dataset $\mathcal{G}={(G_{i},y_{i})}$, we simulate class imbalance by setting the proportions of training samples per class as $\{1,\frac{1}{2^{\beta}},\frac{1}{3^{\beta}},\dots,\frac{1}{|\mathcal{Y}|^{\beta}}\}$, where $\beta\in\{0,0.5,1,1.5,2\}$ controls the imbalance ratio.
-The number of samples in the first class is fixed under all $\beta$ values.
-For graph regression tasks, where $y_{i}\in[y_{\text{min}},y_{\text{max}}]$, we partition the label range into three equal buckets and vary the proportions of the groups as $\{1,\frac{1}{2^{\beta}},\frac{1}{3^{\beta}}\}$ to simulate imbalance.
-A higher $\beta$ indicates a greater level of imbalance.
-As shown in Figure 4, GCN and GMT struggle with both classification and regression tasks due to their reliance on global graph aggregation, which dilutes the features of minority class nodes and edges in imbalanced datasets.
-Similarly, AK+ (and I2GNN) and MO (and HGP), despite incorporating subgraph-level information or reconstructing graph structures, fail to explicitly address class imbalance, limiting their ability to preserve minority class features under skewed distributions.
-In contrast, our model demonstrates superior robustness across all tasks.
-By effectively capturing task-relevant subgraphs and leveraging adaptive contrastive learning, it mitigates the dominance of majority class features and balances feature representations, preserving the features of minority classes even under severe imbalance.
+For the graph classification task, given a dataset $\mathcal{G}={(G_{i},y_{i})}$, we simulate class imbalance by adjusting the proportion of training samples in each class to follow the sequence $\{1,\frac{1}{2^{\beta}},\frac{1}{3^{\beta}},\dots,\frac{1}{|\mathcal{Y}|^{\beta}}\}$, where $\beta\in\{0,0.5,1,1.5,2\}$ determines the degree of imbalance.
+The sample count for the first class remains constant across all $\beta$ settings.
+For graph regression tasks, where $y_{i}$ spans the interval $y_{\text{min}},y_{\text{max}}]$, we divide the label range into three equal-sized buckets and set the group proportions as $\{1,\frac{1}{2^{\beta}},\frac{1}{3^{\beta}}\}$ to create imbalance, with larger $\beta$ corresponding to higher level of imbalance.
+As shown in Figure 4, accuracy drops across various GNN architectures as imbalance increases, particularly in datasets with minority classes.
+Node-based GNN (e.g., GCN) and pooling-based GNN (GMT) struggle because their global aggregation mechanisms tend to average out minority signals, making it difficult to distinguish rare patterns.
+Subgraph-based (AK+) and graph learning-based (MO) GNNs, while capturing richer structural information, do not explicitly address class imbalance, leading to performance degradation under extreme imbalance, especially for less classes.
+SSL-based model (GCA), although capable of leveraging abundant unlabeled data, likewise fails to correct imbalance on its own and converges to performance similar to other baselines unless augmented with imbalance-aware strategies (Liu et al., 2023).
 4.8.3. Few-shot Evaluation
-Specifically, For graph classification tasks, given a training graph dataset $\mathcal{G}=\{(G_{i},y_{i})\}$, we set the number of training graphs per class as $\gamma\in\{5,10,15,20,25\}$.
-For graph regression tasks, we partition the range $[y_{\text{min}},y_{\text{max}}]$ into five equal buckets to create discrete groups and sample $\gamma$ training instances for each bucket.
-As shown in Figure 5, GCN and GMT struggle in few-shot settings due to their reliance on global graph aggregation, which oversmooths features and fails to preserve task-relevant substructures.
-Similarly, AK+, I2GNN, MO, and HGP, while incorporating subgraph-level information, lack mechanisms to prioritize task-relevant features or adapt to low-data conditions, leading to suboptimal performance.
-In contrast, our model demonstrates superior performance across all tasks by explicitly capturing task-relevant subgraphs and leveraging adaptive contrastive learning, achieving robust generalization even with very limited training data.
-5. Conclusion
-In this paper, we systematically revisit existing GNNs for graph-level tasks, categorizing them into five types: node-based, hierarchical pooling-based, subgraph-based, graph learning-based, and self-supervised learning-based approaches.
-To address limitations in current evaluation practices, we proposed a unified evaluation framework that ensures comprehensive and fair comparisons by incorporating diverse datasets, multiple tasks, and challenging scenarios such as noisy, imbalanced, and few-shot graphs.
-Furthermore, we propose a novel GNN model that enhances expressivity through a $k$-path rooted subgraph approach and improves generalization via adaptive graph contrastive learning.
-Extensive experiments demonstrated the superior performance of our model across a wide range of graph-level tasks and twenty-seven datasets.
+We simulate data scarcity by limiting the number of training samples.
+For classification tasks, we construct training sets where the number of labeled graphs per class is in $\gamma\in\{5,10,15,20,25\}$.
+For regression tasks, we partition the label into five equal-width buckets and uniformly sample $\gamma$ training instances per bucket.
+The results, as summarized in Figure 5, reveal that most GNN architectures exhibit significant performance degradation as $\gamma$ decreases.
+Node-based (GCN) and pooling-based (GMT) models are particularly affected, as their global aggregation mechanisms rely heavily on abundant labeled data to learn meaningful representations.
+Interestingly, while subgraph-based models such as AK+ and I2, as well as graph learning-based models like MO and HGP, are theoretically capable of leveraging local structural patterns, they do not demonstrate substantial resilience to data scarcity in practice.
+This is primarily because current implementations lack explicit mechanisms to identify and prioritize the most informative subgraphs or adaptively focus on critical features when labeled data is limited.
+Their performance improvements over standard baselines are thus marginal in few-shot scenarios, indicating that richer local modeling alone does not guarantee data efficiency (Wang et al., 2023c).
+4.9. Correlation Analysis
+4.9.1. Graph Topological Features
+To investigate the intrinsic relationship between graph topology and GNN model performance, we characterize each graph using a comprehensive set of topological features.
+These features capture properties ranging from macroscopic global characteristics to fine-grained local patterns.
+We categorize them into three groups: (i) Global Structural Features, (ii) Local Structural Features, and (iii) Node Distributions.
+Global Structural Features. These metrics provide a high-level summary of the entire graph’s connectivity and organization. They typically return a single scalar value for each graph.
+  • Average Degree (Deg.) and Density (Den.).
+For a graph $G=(V,E)$, let $|V|$ and $|E|$ be the number of nodes and edges, respectively.
+$\text{Avg.
+Deg}=\frac{2|E|}{|V|}$.
+Density is the ratio of actual edges to possible edges:
+    (16) $\text{Density}=\frac{2|E|}{|V|(|V|-1)}.$
+    They measure the overall sparsity.
+  • Average Shortest Path Length (SPL) and Diameter (Dia.).
+Let $d(u,v)$ denote the shortest path distance between nodes $u$ and $v$.
+$\text{SPL}=\frac{1}{|V|(|V|-1)}\sum_{u\neq v}d(u,v)$, and $\text{Diameter}=\max_{u\neq v}d(u,v)$.
+For disconnected graphs, these are computed on the largest connected component.
+  • Degree Assortativity (Assr.).
+Calculated as the Pearson correlation coefficient between the degrees of connected nodes.
+Positive values indicate that high-degree nodes tend to connect with other high-degree nodes.
+  • Modularity (Mod.). Quantifies the strength of community structure. A positive value indicates that connections within communities are denser than expected by random chance.
+Local Structural Features. These metrics focus on small, recurring substructures. For node-level metrics, we aggregate them into graph-level features.
+  • Average Clustering Coefficient (CC).
+For a node $v$, the local clustering coefficient $C_{v}$ is the density of the subgraph induced by its neighbors.
+We report the average over all nodes: $\text{Avg.
+CC}=\frac{1}{|V|}\sum_{v\in V}C_{v}$.
+  • Triangle Counts (Tri.).
+Let $T_{v}$ be the number of triangles node $v$ participates in.
+Unlike global metrics, to capture both the prevalence and distribution of triangles, we report both the mean ($\mu$) and standard deviation ($\sigma$) of $T_{v}$ across all nodes.
+High $\mu$ implies a dense graph, while high $\sigma$ indicates a significant degree of heterogeneity.
+Node Distributions.
+These features analyze the distribution of node importance.
+Since these are originally defined at the node level, we characterize the entire graph by reporting both the Mean ($\mu$) and Standard Deviation ($\sigma$) for each metric.
+  • Betweenness Centrality (BC.).
+$C_{B}(v)$ represents the fraction of all-pairs shortest paths passing through node $v$.
+We report $\mu_{\text{BC}}$ and $\sigma_{\text{BC}}$.
+A high $\sigma_{\text{BC}}$ indicates the presence of critical ”bridge” nodes controlling information flow.
+  • PageRank (PR.).
+Measures node influence based on transitive importance.
+We report $\mu_{\text{PR}}$ and $\sigma_{\text{PR}}$.
+A high $\sigma_{\text{PR}}$ reveals a hierarchical structure with distinct authority nodes.
+  • K-Core Number (KC.).
+The core number $k_{v}$ is the largest $k$ such that $v$ belongs to a $k$-core.
+We report $\mu_{\text{KC}}$ and $\sigma_{\text{KC}}$.
+A high $\mu_{\text{KC}}$ indicates global robustness, while high $\sigma_{\text{KC}}$ reflects a core-periphery structure.
+4.9.2. Evaluation
+Given GNNs $\mathcal{M}=\{M_{i}\}_{i=1}^{n}$ and graph dataset $\mathcal{D}=\{G_{j}\}_{j=1}^{m}$, we construct a performance matrix $\mathbf{P}\in\mathbb{R}^{n\times m}$ where $\mathbf{P}[i][j]$ is the result of $M_{i}$ on $G_{j}$ (e.g., accuracy or MAE).
+We denote each $k$-th topological feature of $\mathcal{D}$ as $\mathbf{f}_{k}$, where $\mathbf{f}_{k}[j]$ is the value of the $k$-th feature for graph $G_{j}$.
+We assess the relationship between model performance and $k$-th feature with Spearman’s correlation $\rho(\mathbf{P}[i],\mathbf{f}_{k})$, with a significance threshold of $\alpha=0.05$.
+4.9.3. Insights
+Figure 6. Correlation between graph features and GNNs. Each cell shows the Spearman’s $\rho$ with * for $p<0.05$ and $\dagger$ for $0.05\leq p<0.1$.
+As illustrated in Figure 6, we observe:
+  • Structurally simpler models (e.g., Node-based) show weak correlations with most features, reflecting their reliance on local aggregation rather than global topology.
+  • Graph density negatively correlates with most models, suggesting high connectivity exacerbates over-smoothing and noise.
+  • High local clustering and assortativity negatively impact Subgraph-based and SSL-based models, likely due to the “rich-club” effect distracting from peripheral structures.
+  • Conversely, graph sparsity and high Betweenness Centrality positively correlate with Hierarchical (e.g., HGP) and SSL-based models, demonstrating their ability to leverage clear structures for long-range information flow.
+Most importantly, the absence of universally significant correlations confirms that (i) model selection cannot rely on a single structural indicator, and (ii) no single architecture consistently dominates.
+This underscores the critical importance of comprehensive benchmarks across diverse domains, as conducted in OpenGLT, to provide an empirical basis for scenario-specific model selection.
+5. Conclusion and Future Direction
+This paper presents a comprehensive experimental study on graph neural networks for graph-level tasks, categorizing existing models, introducing a unified evaluation framework (OpenGLT), and benchmarking 16 GNNs across 26 diverse datasets under challenging scenarios such as noise, imbalance, and limited data.
+Our results reveal that no single model excels universally, highlighting important trade-offs between expressiveness and efficiency, and emphasizing the need for robust evaluation in realistic conditions.
+Based on these findings, promising future directions include the development of scenario-adaptive or hybrid GNN architectures that dynamically leverage different model strengths, research into lightweight and scalable algorithms for practical deployment, and the incorporation of transfer and foundation model techniques to improve generalization and data efficiency, particularly when labeled data is scarce.
+Our study provides valuable benchmarks and guidance for advancing GNN research on graph-level tasks.